### v1

Abstract
We present simple to implement Wald-type statistics that deliver a general nonparametric inference theory for linear restrictions on varying coefficients in a range of spatial autoregressive models. Our theory covers error dependence of a general form, allows for a degree of misspecification robustness via nonparametric spatial weights and permits inference on both varying regression and spatial coefficients. One application of our method finds evidence for constant returns to scale in the production function of the Chinese nonmetal mineral industry, while another finds a nonlinear impact of the distance to the employment center on housing prices in Boston. A simulation study confirms that our tests perform well in finite-samples.
JEL codes: C21, C31
Keywords: Spatial autoregression, varying coefficients, inference
1.   Introduction
This paper develops a general framework for inference on varying coefficient spatial autoregressive (SAR) models that is as easy to implement in practice as familiar Wald tests. SAR models are fast becoming a workhorse for researchers working in settings where economic agents interact via economic or social links. Such links can be geographic proximity, or proximity in some more general sense such as a social network. The SAR model can accommodate these types of general links easily, and this goes some way to explain its appeal. No surprise then that the baseline SAR model, due to Cliff and Ord (1981), has become much studied by econometricians. This model has long been popular with regional scientists, see Anselin (1988) for an early book length treatment; important econometric contributions that established the foundation for rigorous econometric analysis of such models include Kelejian and Prucha (1998) and Lee (2004).
In keeping with the growing importance of SAR models, various strands of the literature have explored how to widen their scope. One such is the focus on varying coefficient SAR models, wherein model parameters are allowed to be unknown functions of some observed economic variable(s). This permits a flexible effect of the regressors and spatial dependence on the outcome variable, and indeed varying coefficient models are abundant in statistics and econometrics (see e.g. Fan and Zhang, 1999, 2008). In the SAR context, varying coefficient models have been studied by Sun et al. (2014) and a sequence of important papers by Sun (2016), Malikov and Sun (2017), Sun and Malikov (2018) and Sun (2024), for example. More generally, semiparametric SAR models have attracted much interest, see e.g. Su and Jin (2010), Su (2012), Robinson and Thawornkaiwong (2012) and Zhang (2013), but their focus tends to be on inference on the parametric component.
In this paper, we show how inference on varying coefficients in a range of SAR models can be conducted just like familiar parametric Wald tests for linear restrictions. The key idea is that if series approximations are used for the nonparametric varying coefficients then the series coefficients can be employed for very easy practical inference. This is motivated by the scientific aim of simplification of nonparametric/semiparametric inference. Indeed, a growing literature stresses such ideas, see e.g Ackerberg et al. (2012), Gupta (2018) and Korolev (2019), amongst others.
[…]
Using this series expansion approach, we theoretically justify Wald-type test statistics for inference on varying coefficients. We cover a baseline high-order SAR model with varying regression coefficients but constant spatial lag parameters and extend the model to allow for spatial error dependence. The baseline model is similar to the one estimated by Sun et al. (2014), but also features a growing number of spatial lags and covariates. We then show how to allow for a degree of misspecification robustness by incorporating nonparametric spatial weights à la Pinkse et al. (2002). Our approach is also shown to work for models with varying coefficients on the spatial lags, in the spirit of Malikov and Sun (2017).
We then apply our methods on two different studies where varying coefficients have been shown to play an important role. In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry. This can be seen as a follow-up application to that of Li et al. (2002), who sugges that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e. output elasticities of capital and labor sum to unity) for most observed managerial expense levels. However, they do not have a test for CRS and do not model spatial dependence. Our analysis confirms their observations as we cannot reject the CRS hypothesis using their model and in much richer settings. This suggests that CRS technology may be a salient feature of this industry.
Our second study is an illustration on the famous Boston housing data. We estimate a model similar to Sun et al. (2014), who used a model selection procedure to recommend that location should have a varying effect on house prices with respect to some other features of the property. We formally test for the varying location effects of these features using our statistics and indeed find them to be statistically significant.
In a set of Monte Carlo simulation studies, we show that our tests have excellent finite-sample performance. In addition to tests that are based on the asymptotic normal approximation, we also propose a bootstrap version of our tests; the bootstrap tests perform particularly well for small sample sizes. We experiment with a range of SAR models and spatial dependence structures to demonstrate the full range of applicability and generality of our approach.
The next section introduces our method in the context of a baseline higher-order SAR model. Section 3 extends the model to allow for spatial error dependence of Kelejian and Prucha (2007) type, while Section 4 allows for nonparametric spatial weights (Pinkse et al., 2002). In Section 5 we allow the coefficient on the spatial lag to also vary (Malikov and Sun, 2017) and show how our approach can be adapted to this setting. Section 6 presents our empirical applications to the Chinese nonmetal mineral industry and Boston house prices, while Section 7 contains the simulation study. Proofs are in the appendix, and a supplementary online appendix contains a number of additional simulation results.
2.   Baseline spatial autoregression
Consider a vector of unknown functions $\delta\left(\cdot\right)=\left(\delta_{1}\left(\cdot\right),\ldots,\delta_{d_{\delta}}\left(\cdot\right)\right)^{\prime}$, and the model
$y_{in}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{in,j}^{\prime}y_{n}+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (2.1)
where $y_{n}$ is the $n\times 1$ vector with typical element $y_{in}$, $w_{in,j}^{\prime}$ is the $i$th row of the spatial weight matrix $W_{jn}$, $j=1,\ldots,d_{\lambda}$, $d_{\lambda}\rightarrow\infty$ as $n\rightarrow\infty$, and $\epsilon_{in}$ is $i.i.d.$ with mean 0 and unit variance. Also, $\left(x_{in},p_{in},z_{in}\right)\in\mathcal{X}\times\mathcal{P}\times\mathcal{Z}\subseteq\mathbb{R}^{d_{\beta}\times d_{\delta}\times d_{z}}$, and $d_{\beta}\rightarrow 0$ as $n\rightarrow\infty$ but $d_{z}$ and $d_{\delta}$ are fixed. For all $i=1,\ldots,n$, $z_{in}$ is throughout the paper uncorrelated with $\epsilon_{in}$ while $x_{in}$ is allowed to be correlated with $\epsilon_{in}$.
Permitting $d_{\lambda}$ and $d_{\beta}$ to diverge with sample size allows a flexible modelling approach, with the model becoming richer as more observations become available to the practitioner. This approach has been studied in statistics since at least the work of Huber (1973), but in spirit the sequence of experiments considered by Le Cam (1960) provides an even earlier reference. Econometricians have also studied such models in least squares and GMM settings, see e.g. Andrews (1985), Koenker and Machado (1999) and Cattaneo et al. (2018), to name just three references. In a SAR setting, Gupta and Robinson (2015), Gupta (2018) and Gupta and Qu (2024) emphasize the advantages of this approach, especially when clustered data naturally imply asymptotic regimes in which $d_{\lambda}$ diverges with $n$. It is worth noting that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed but we state our theorems for the more complex case where they diverge.
We now mostly drop $n$ subscripting, but occasionally remind the reader of the dependence on sample size of various quantities. The paper takes the spatial weight matrices $W_{j}$ to be non-stochastic everywhere except in Section 4. In all sections these weight matrices are also uniformly bounded in row and column sums, a commonly used restriction to control spatial dependence (Kelejian and Prucha, 1998; Lee, 2004). This property is assumed to hold almost surely in Section 4, where the spatial weights are stochastic.
Our aim is to do inference on $\delta(\cdot)$, specifically we wish to test a fixed number $m$ restrictions of the form $S\delta(z)=s,z\in\mathcal{Z}$, where $S$ is a known, constant $m\times d_{\delta}$ matrix and $s$ is a known, constant $m\times 1$ vector. Because we can convert any linear restriction to an exclusion restriction, we will focus on tests of the null hypothesis
$H_{0}^{true}:\delta(z)=0,z\in\mathcal{Z}.$ (2.2)
Our test statistics will approximate $\delta(\cdot)$ by a series expansion and test if the coefficients in the expansion are jointly zero, thus converting the restriction in $H_{0}^{true}$ to an increasing-dimensional one, up to some suitably negligible approximation error.
2.1   Test statistic
We approximate $\delta_{k}(z)$ by $\psi_{k}^{h_{k}}(z)^{\prime}\alpha_{k}^{h_{k}}$, where $\psi_{k}^{h_{k}}(z)=\left(\psi_{k1}(z),\ldots,\psi_{kh_{k}}(z)\right)^{\prime}$ and $\alpha_{k}^{h_{k}}=\left(\alpha_{k1},\ldots,\alpha_{kh_{k}}\right)^{\prime}$, for some basis functions $\psi_{k\ell}(\cdot)$, $\ell=1,\ldots,h_{k}$, $k=1,\ldots,d_{\delta}$. Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$. Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$, with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$. Then we can write (2.1) as
$y_{i}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{i,j}^{\prime}y+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (2.3)
with $u_{i}=r_{i}+\epsilon_{i}$, where $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error. Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$. Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can be written in matrix notation as
$y=Q\lambda+X\beta+\Psi\alpha+u=L\xi+u,$ (2.4)
where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$.
The null hypothesis to approximate $H_{0}^{true}$ is
$H_{0}:\alpha=0,$ (2.5)
and our test will be implemented using the 2-stage least squares (2SLS) estimator
$\hat{\xi}\equiv\left(\hat{\lambda}^{\prime},\hat{\beta}^{\prime},\hat{\alpha}^{\prime}\right)^{\prime}=\left(L^{\prime}{P_{K_{1}}}L\right)^{-1}L^{\prime}{P_{K_{1}}}y,$ (2.6)
where $P_{K_{1}}=K_{1}\left(K_{1}^{\prime}K_{1}\right)^{-1}K_{1}^{\prime}$, with $K_{1}$ being an $n\times J_{1}$ instrument matrix with $J_{1}\geq d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$ but having the same asymptotic order i.e. ${J_{1}}/{d_{\xi}}$ tends to a constant that is at least unity, as $n\rightarrow\infty$. Let $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\lambda}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix. . Then, the test statistic is
$\mathbb{W}_{1}=\frac{n\hat{\alpha}^{\prime}\mathscr{D}_{1}^{-1}\hat{\alpha}-d_{\alpha}}{\sqrt{2d_{\alpha}}}.$ (2.7)
2.2   Asymptotic properties
We begin with restrictions on the moments of various objects, noting that throughout the paper $C$ denotes a generic positive constant, arbitrarily large but independent of $n$. Let $k_{ri}$ be the elements of the instrumental variable matrix, which corresponds to $K_{1},K_{2}$ and $K_{3}$ as defined in this section, 4 and 5, respectively, and $l_{ri}$ denote the elements of $L$.
Assumption 1.
$\sup_{i\geq 1}\mathrm{E}\left(r_{i}^{2}\right)=o(n^{-1})$.
Assumption 2.
$\mathrm{E}\left|\epsilon_{i}\right|^{q}<C$, for some $q>4$.
Assumption 3.
$\mathrm{E}\left(l_{ri}^{2}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
Assumption 1 is a general assumption that controls the approximation error. Sufficient conditions for various cases can be found in Chen (2007). We avoid using specific rates for the approximation error because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions. Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition. Assumption 3 imposes moment conditions on regressors and instruments.
To ease notation, we introduce the following definition which captures boundedness and non-multicollinearity properties of random or fixed matrices of growing dimension.
Definition 1.
For any square, symmetric and positive semi-definite matrix $A$, let $\overline{\alpha}(A)$ and $\underline{\alpha}(A)$ denote its largest and smallest eigenvalues, respectively. If $A$ is random, we say that $A$ has Property G if
$\overline{\alpha}(A)=O_{p}(1)\text{ and }\left\{\underline{\alpha}(A)\right\}^{-1}=O_{p}(1).$
If $A$ is non-random, we say that $A$ has Property G if
$\limsup_{n\rightarrow\infty}\overline{\alpha}(A)<\infty\text{ and }\liminf_{n\rightarrow\infty}\underline{\alpha}(A)>0.$
Assumption 4.
$n^{-1}K_{1}^{\prime}K_{1}$, $n^{-1}L^{\prime}K_{1}K_{1}^{\prime}L$ and $n^{-1}\Psi^{\prime}\Psi$ have 1.
We first show that the statistic $\mathbb{W}_{1}$ can be approximated by a quadratic form in $\epsilon$. Define $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$, with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$. Then, we have the following theorem.
Theorem 2.1.
Under $H_{0}$, Assumptions 1-4 and
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}}{n}\rightarrow 0,$ (2.8)
as $n\rightarrow\infty$,
$\mathbb{W}_{1}-\frac{\epsilon^{\prime}\mathcal{M}_{1}\epsilon-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
We take this opportunity to remind the reader that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
Assumption 5.
$\left(I_{n}-\sum_{j=1}^{d_{\lambda}}\lambda_{j}W_{j}\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$.
Define $H_{\ell 1}\equiv H_{\ell 1,n}:\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$, with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proof of the next theorem. This sequence of local alternatives features a $d_{\alpha}^{1/4}$ damping factor that accounts for the cost of our nonparametric approach, and has been found in similar problems by Hong and White (1995) and Gupta (2018), amongst others. We can now state the main theorem of the section.
Theorem 2.2.
Under Assumptions 1-5 and
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}\rightarrow 0,$ (2.9)
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{1}$ provides a consistent test.
[…]
3.   SAR with error spatial dependence
Having demonstrated our methods with a baseline model, we now consider (2.1) but relax the $i.i.d$ assumption on $\epsilon_{i}$ to
$\epsilon_{i}\equiv\epsilon_{in}=\sum_{j=1}^{n}b_{ij}v_{j},$ (3.1)
where the $v_{j}$ are $i.i.d$ with mean 0 and variance 1, and the $b_{ij}\equiv b_{ijn}$ can depend on $n$. This is the Kelejian and Prucha (2007) approach to error spatial dependence, and indeed we will employ the spatial heteroskedasticity and autocorrelation consistent (SHAC) method pioneered in that paper. Our imposition that $z_{i}$ be uncorrelated with $\epsilon_{i}$ now, and subsequently, requires that $z_{i}$ is uncorrelated with all the $v_{j}$, $j=1,\ldots,n$.
3.1   Test statistic
Let $\mathscr{D}_{2}=\frac{1}{n}R\mathcal{U}_{1}R^{\prime}$ with $\mathcal{U}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\Xi_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$ and $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$ with $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, where $\hat{\Xi}_{1}$ is the Kelejian and Prucha (2007) SHAC estimate of $\Xi_{1}=\frac{1}{n}K_{1}^{\prime}\Sigma K_{1}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$ being the error variance matrix. We still use the 2SLS estimator $\hat{\xi}$ and the test statistic is now
$\mathbb{W}_{2}=\frac{n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{2}^{-1}\hat{\alpha}-d_{\alpha}}{\sqrt{2d_{\alpha}}}.$ (3.2)
We introduce the following assumptions:
Assumption 6.
$\Sigma$ has 1.
Assumption 7.
$\sup_{i\geq 1}\sum_{j=1}^{n}\left|b_{ij}\right|+\sup_{j\geq 1}\sum_{i=1}^{n}\left|b_{ij}\right|<\infty$.
Assumption 8.
$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel function that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g. Kelejian and Prucha (2007), Delgado and Robinson (2015) and Gupta and Qu (2024) for similar assumptions. Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g. Kelejian and Prucha (2007).
Now we introduce distance measures $d_{ij,m}=d_{ji,m},m=1,\ldots,M$. As in Kelejian and Prucha (2007), we allow for measurement errors and so, in the following, let $d^{*}_{ij,m}=d^{*}_{ji,m}\geq 0$ be the actual distance measures used in practice. Corresponding to each measure, assume that the researcher can select a distance $d_{m}>0$ satisfying $d_{m}\uparrow\infty$ as $n\rightarrow\infty$. For each unit $i=1,\ldots,n$, let $\ell_{i}=\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)$ and set $\ell=\max_{i}\ell_{i}$. Observe that $\ell_{i}$ is the number of units $j$ for which $d^{*}_{ij,m}\leq d_{m}$ for at least one $m=1,\ldots,M$. Also let $\sigma_{ij}$ be a typical element of $\Sigma$.
Assumption 9.
(a) $\mathrm{E}(\ell^{2})=o\left(n^{2\eta}\right)$ where $\eta<\frac{1}{2}(q-2)/(q-1)$ with $q>4$ in Assumption 2. (b) $\sum_{j=1}^{n}\left|\sigma_{ij}\right|d^{\chi}_{ij,1}<C$ for some $\chi\geq 1$. (c) $d^{*}_{ij,m}=d_{ij,m}+\nu_{ij,m}\geq 0$, with $\left|\nu_{ij,m}\right|<C$ and $\nu_{ij,m}$ independent of $v_{i}$ for all $m=1,\ldots,M$.
Then, the $(r,s)$-th element of $\hat{\Xi}_{1}$ is
$\hat{\Xi}_{1,rs}=n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{ri}k_{sj}\hat{u}_{j}\hat{u}_{j}\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),r,s=1,\ldots,J_{1}.$ (3.3)
3.2   Asymptotic properties
In all subsequent lemmas and theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9. In Lemma 3.1 and Theorems 3.1-3.2, which follow, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$. Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e. the spectral norm of $A$.
Lemma 3.1.
Let Assumptions 1, 3, 4, 6-9 hold, and
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\xi}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (3.4)
Then,
$\big{\|}\hat{\Xi}_{1}-\Xi_{1}\big{\|}=o_{p}(1).$
The condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $d_{\xi}/n\rightarrow 0$ by Assumption 9(a). Now, define $\mathcal{M}_{2}=\frac{1}{n}B^{\prime}K_{1}\mathcal{V}_{2}K_{1}^{\prime}B$, with $\mathcal{V}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$. Then we have the following theorem.
Theorem 3.1.
Let (3.4) hold. Then, under $H_{0}$, Assumptions 1, 3, 4, 6-9, as $n\rightarrow\infty$,
$\mathbb{W}_{2}-\frac{v^{\prime}\mathcal{M}_{2}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
To establish the asymptotic distribution of $\mathbb{W}_{2}$, the following assumption is imposed to check a Lyapunov condition, as also e.g. by Delgado and Robinson (2015) and Gupta and Qu (2024).
Assumption 10.
$\mathrm{E}\left|v_{i}\right|^{s}<C$, for some $s\geq 8$.
Define $H_{\ell 2}\equiv H_{\ell 2,n}:\alpha=\alpha^{*}\equiv\nu_{2n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{2n}^{\prime}\Gamma_{2n}\nu_{2n})^{\frac{1}{2}}$, with $\nu_{2n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{2n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proof of the next theorem.
Theorem 3.2.
Under Assumptions 1, 3-10, with
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}+n^{\eta-1}d_{\xi}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,$ (3.5)
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{2}$ provides a consistent test.
[…]
Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $\frac{d_{\xi}^{3}}{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
4.   Misspecification robustness
We now consider a degree of misspecification robustness in (2.1), which allows $\epsilon_{i}$ to have the structure in Section 3. Focusing on the case $d_{\lambda}=1$ for notational simplicity, and writing $W_{1}=W$, (2.1) imposed the parametric form $\lambda Wy$ and took $W$ as known in the “spatial lag” term. We now allow the spatial lag in outcomes to take the nonparametric form $Gy$, where $G$ has elements $g\left(d_{rs}\right)$, $r,s=1,\ldots,n$, for some vector of exogenous (independent of $v_{j}$, $j=1,\ldots,n$) economic distance measures $d_{rs}$. This idea was introduced by Pinkse et al. (2002), and has been much used since (see e.g. Sun, 2016 and Gupta and Qu, 2024).
4.1   Test statistic
We approximate the elements $g\left(d_{ij}\right)$ of $G$ with a series of basis functions. Following Pinkse et al. (2002), we suppose
$g(d_{ij})=\sum_{l=1}^{\infty}\tau_{l}e_{l}(d_{ij}),$ (4.1)
where $\tau_{l}$ are unknown coefficients, and the $e_{l}$ form a basis of the function space to which $g$ belongs. Let $c_{i}^{\prime}$ be the $i$-th row of the $n\times d_{\tau}$ matrix $\mathfrak{C}$ with typical $(i,l)$-th element $\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $\tau=\left(\tau_{1},\ldots,\tau_{d_{\tau}}\right)^{\prime}$, $d_{\tau}\rightarrow\infty$ as $n\rightarrow\infty$. Then we generalize (2.1) to
$y_{in}=g_{in}^{\prime}y+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (4.2)
and (2.3) to
$y_{i}=c_{i}^{\prime}\tau+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (4.3)
with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$. In the matrix notation, we now have
$y=\mathfrak{C}\tau+X\beta+\Psi\alpha+u=F\theta+u,$ (4.4)
where $F=[\mathfrak{C},X,\Psi]$ and $\theta=\left(\tau^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. Our test is implemented using the 2SLS estimator
$\hat{\theta}=\left(F^{\prime}{P_{K_{2}}}F\right)^{-1}F^{\prime}{P_{K_{2}}}y,$ (4.5)
where $P_{K_{2}}=K_{2}\left(K_{2}^{\prime}K_{2}\right)^{-1}K_{2}^{\prime}$, $K_{2}$ being an $n\times J_{2}$ instrument matrix with $J_{2}\geq d_{\theta}=d_{\tau}+d_{\beta}+d_{\alpha}$, but with $J_{2}$ and $d_{\theta}$ having the same asymptotic order.
Let $\mathscr{D}_{3}=\frac{1}{n}R\mathcal{U}_{2}R^{\prime}$ with $\mathcal{U}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\Xi_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$ and $T_{2}=\frac{1}{n}F^{\prime}{P_{K_{2}}}F$, and let $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$ with $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\tau}+d_{\beta}\right)},I_{d_{\alpha}}\right]$ and $\hat{\Xi}_{2}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{2}=\frac{1}{n}K_{2}^{\prime}\Sigma K_{2}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. Then, the test statistic is
$\mathbb{W}_{3}=\frac{n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{3}^{-1}\hat{\alpha}-d_{\alpha}}{\sqrt{2d_{\alpha}}}.$ (4.6)
Denote the elements of $F$ as $f_{ri}$. We have following assumptions.
Assumption 11.
$\mathrm{E}\left(f^{2}_{ri}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
Assumption 12.
$\sup_{i\geq 1}\mathrm{E}\left(r_{iG}^{2}\right)=o(n^{-1})$ and $\sup_{i\geq 1}\mathrm{E}\left(r_{i\delta}^{2}\right)=o(n^{-1})$.
Assumption 13.
$n^{-1}K_{2}^{\prime}K_{2}$ and $n^{-1}F^{\prime}K_{2}K_{2}^{\prime}F$ have 1.
Lemma 4.1.
Let Assumptions 6-9, 11-13 hold, and
$\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\theta}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,\text{as }n\rightarrow\infty.$ (4.7)
Then,
$\bigg{\|}\hat{\Xi}_{2}-\Xi_{2}\bigg{\|}=o_{p}(1).$
4.2   Asymptotic properties
Define $\mathcal{M}_{3}=\frac{1}{n}B^{\prime}K_{2}\mathcal{V}_{3}K_{2}^{\prime}B$, with $\mathcal{V}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\mathscr{D}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$. Then, we have the following theorem.
Theorem 4.1.
Let (4.7) hold. Then, under $H_{0}$, Assumptions 6-9, 11-13, as $n\rightarrow\infty$,
$\mathbb{W}_{3}-\frac{v^{\prime}\mathcal{M}_{3}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. The following assumption generalizes Assumption 5 to the setting of this section.
Assumption 14.
$\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $d_{rs},r,s=1,\ldots,n$.
Theorem 4.2.
Under Assumptions 6-14, with
$\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\theta}^{3}}{n}+n^{\eta-1}d_{\theta}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,$ (4.8)
as $n\rightarrow\infty$, the following hold:
[…]
(2) $\mathbb{W}_{3}$ provides a consistent test.
[…]
5.   Varying spatial coefficients
In this section we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but now allow the spatial lag parameters $\lambda_{j}$ to vary. The model we consider is closely related to Malikov and Sun (2017) and Sun and Malikov (2018), with the difference that we do not study panel data settings. Instead, we allow $d_{\beta}\rightarrow\infty$ and general spatial dependence in the disturbances. Our model is also related to the one studied by Sun (2024), wherein the model includes a nonparametric nuisance regression function but does not feature dependent disturbances. Thus, our model complements the extant literature.
5.1   Test statistic
Specifically, we have
$y_{i}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}\left(z_{i}\right)w_{i,j}^{\prime}y+x_{i}^{\prime}\beta+p_{i}^{\prime}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$ (5.1)
with $\lambda_{j}(\cdot)$ unknown scalar-valued functions and $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j}$, i.e. error with spatial dependence. Introduce the approximation $\lambda_{j}\left(z_{i}\right)=\sum_{k=1}^{l_{j}}\mu_{kj}\phi_{kj}\left(z_{i}\right)+r_{ij\lambda},j=1,\ldots,d_{\lambda}$. Set $d_{\mu}=\sum_{j=1}^{d_{\lambda}}l_{j}$, $\phi_{j}=\left(\phi_{1j},\ldots,\phi_{l_{j}j}\right)^{\prime}$, and let $H(z)$ be the $n\times d_{\mu}$ matrix with $i$-th row $\left(y^{\prime}w_{i,1}\phi_{1}\left(z_{i}\right),\ldots,y^{\prime}w_{i,d_{\lambda}}\phi_{d_{\lambda}}\left(z_{i}\right)\right)$.
Then,
$y=G(z)\gamma+u,$ (5.2)
where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$. Defining $\Lambda_{j}(z)=diag\left(\lambda_{j}\left(z_{1}\right),\ldots,\lambda_{j}\left(z_{n}\right)\right)$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
$y=S(z)^{-1}\left(X\beta+t(z)+\epsilon\right),$ (5.3)
with $P$ having $i$-th row $p_{i}^{\prime}$ and $t(z)$ serving to stack $p_{i}^{\prime}\delta\left(z_{i}\right)$ into an $n\times 1$ vector with conformable subscript to $y_{i}$ and $\epsilon_{i}$.
The null hypothesis in this setting is
$H_{0}:\vartheta=0,$ (5.4)
where $\vartheta$ is either $\mu$ or $\alpha$ and our test will be implemented using the 2SLS estimator
$\hat{\gamma}=\left(G^{\prime}{P_{K_{3}}}G\right)^{-1}G^{\prime}{P_{K_{3}}}y,$ (5.5)
where $P_{K_{3}}=K_{3}\left(K_{3}^{\prime}K_{3}\right)^{-1}K_{3}^{\prime}$ with $K_{3}$ being an $n\times J_{3}$ instrument matrix with $J_{3}\geq d_{\gamma}=d_{\mu}+d_{\beta}+d_{\alpha}$, but with $J_{3}$ and $d_{\gamma}$ having the same asymptotic order. Note that we have omitted the $z$ argument for brevity.
Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. Then, the test statistic is
$\mathbb{W}_{4}=\frac{n\hat{\vartheta}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}-d_{\vartheta}}{\sqrt{2d_{\vartheta}}},$ (5.6)
where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$.
Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
Assumption 15.
$\sup_{z\in\mathcal{Z}}\mathrm{E}\left(g^{2}_{ri}(z)\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
Assumption 16.
$\sup_{i\geq 1}\mathrm{E}\left(r_{i\lambda}^{2}\right)=o(n^{-1})$ and $\sup_{i\geq 1}\mathrm{E}\left(r_{i\delta}^{2}\right)=o(n^{-1})$.
Assumption 17.
$n^{-1}K_{3}^{\prime}K_{3}$ and $\sup_{z}n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1.
Lemma 5.1.
Let Assumptions 6-9, and 15 -17 hold, and
$\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (5.7)
Then, we have
$\bigg{\|}\hat{\Xi}_{3}-\Xi_{3}\bigg{\|}=o_{p}(1).$
5.2   Asymptotic properties
Define $\mathcal{M}_{4}=\frac{1}{n}B^{\prime}K_{3}\mathcal{V}_{4}K_{3}^{\prime}B$, with $\mathcal{V}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\mathscr{D}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$. Then, we have the following theorem.
Theorem 5.1.
Let (5.7) hold. Then, under $H_{0}$, Assumptions 6-9, 15-17, as $n\rightarrow\infty$,
$\mathbb{W}_{4}-\frac{v^{\prime}\mathcal{M}_{4}v-d_{\vartheta}}{\sqrt{2d_{\vartheta}}}=o_{p}(1).$
The following assumption is the appropriate version of Assumptions 5 and 14 for the model considered in this section.
Assumption 18.
$\left(I_{n}-S(z)\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $z$.
Define $H_{\ell 4}\equiv H_{\ell 4,n}:\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{\frac{1}{2}}$, with $\nu_{4n}$ a $d_{\vartheta}\times 1$ non-zero vector and $\Gamma_{4n}$ a $d_{\vartheta}\times d_{\vartheta}$ matrix defined in the proof of the next theorem.
Theorem 5.2.
Under Assumptions 6-10 and 15-18, suppose that
$\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\gamma}^{3}}{n}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,$ (5.8)
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{4}$ provides a consistent test.
[…]
6.   Empirical applications
6.1   Testing for the CRS in a production function
In this subsection, we are interested in studying whether the production function in China’s nonmetal mineral manufacturing industry has a constant returns to scale (CRS) technology. CRS refers to a scenario where a simultaneous proportional increase in all inputs results in an identical proportional increase in output. Returns to scale is a classical concept in economics to assess the efficiency of a production function, and it has accordingly received much attention in the empirical literature (for recent works, see Ackerberg et al., 2015; Basu and Bundick, 2017; Attanasio et al., 2020; Combes et al., 2021). Existing works, however, mostly do not account for cross-sectional dependence between firms. The role of cross-firms correlation has recently been highlighted empirically by Iyoha (2023), who shows that this can capture spillover effects and would lead to model misspecification if left unaccounted for.
We consider models that generalize Li et al. (2002)’s production function by incorporating spatial dependence in the analysis, thereby allowing for cross-sectional dependence. Their analysis centers on a Cobb-Douglas style production function that allows the elasticities of capital and labor to be nonparametrically varying with management expenses. Management expenses are costs that are indirectly associated with output production, such as research and development (R&D), equipment upgrades and employee training. In this framework, CRS corresponds to the elasticities of capital and labor summing to one.
Using data from the Third Industrial Census of China (conducted by the National Statistical Bureau in 1995) on the non-metal mineral manufacturing industry, Li et al. (2002) find that ignoring the varying nature of these coefficients leads to an underestimation of the returns to scale. Moreover, they suggest their estimate of the returns to scale of the technology is close to being constant for most values of managerial expense, but they do not have a test for the CRS hypothesis.
Our data comes from the Chinese Industrial Enterprise Database for the year of 2014 as this is the most recent information available in the database. We use data on all firms in the nonmetal mineral manufacturing industry with the survey code 31, resulting in a sample of 7,355 observations. As advocated by Li et al. (2002), the appeal for focusing on this industry is that production functions of firms are expected to be homogeneous since it has a very small proportion of foreign entity ownership. This is important because evidence suggests that production performance may vary across ownership types (Murakami et al., 1994). Table 1 introduces all variables after taking logarithms.
Our variables are output ($y_{i}$), capital ($p_{i1}$), labor ($p_{i2}$), and management expenses ($z_{i}$). As in Li et al. (2002), we include a parametric benchmark where the Cobb-Douglas production function has a quadratic form of management expenses without the varying coefficient entering additively, namely:
$y_{i}=\delta_{0}+\delta_{1}p_{i1}+\delta_{2}p_{i2}+\beta_{1}z_{i}+\beta_{2}z_{i}^{2}+\epsilon_{i}.$ (6.1)
The elasticities of capital and labor are $\delta_{1}$ and $\delta_{2}$, respectively. The semiparametric production function where the elasticities of of capital and labor vary with management expenses is:
$y_{i}=\delta_{0}(z_{i})+\delta_{1}(z_{i})p_{i1}+\delta_{2}(z_{i})p_{i2}+\epsilon_{i}.$
Based on the CRS hypothesis, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
$y_{i}^{*}=\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i},$
where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate
$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.2)
where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for $k=1,...,h/2$. We approximate the unknown functions by polynomial and trigonometric basis functions of different orders. We define the series approximation based Wald statistic derived from this model with no spatial autoregressive component as $\mathbb{W}_{0}$.
Next, we incorporate the SAR structure and consider:
$y_{i}^{*}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\beta_{0}+\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i}.$
We estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.3)
with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$. We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
$\begin{split}&w_{ij}=[W_{n}^{p}]_{ij}=\begin{cases}1\quad\text{iffirms}\ i\ \text{and}\ j\ \text{are in the same prefecture-level city},\\ 0\quad\text{otherwise},\end{cases}\\ &w_{ij}=[W_{n}^{d}]_{ij}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise},\end{cases}\end{split}$
where $d_{ij,m}^{\ast}$ measures the geographical distance between firms $i$ and $j$, and $d_{m}$ represents the 10th percentile of all geographical distances.
For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$. Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for $k=1,...,h$, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function $k=1,...,h/2$ for $h=2,4$. To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function to estimate $\mathscr{K}\left({d_{ij,m}^{\ast}}/{d_{m}}\right)$. For $\mathbb{W}_{3}$, we denote $\widehat{w}_{ij,l}=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$ if $i\neq j$ for $l=1,...,h$. Then, $c_{il}=\sum_{j=1}^{n}\widehat{w}_{ij,l}y_{j}$ for $l=1,...,h$. For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$. Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$. For the IVs, we use $K_{1}=[W\Psi,X,\Psi]$ with respect to $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$, and $K_{2}=[X,\Psi]$ and $K_{3}=[H\Psi,X,\Psi]$ with respect to $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$.
The results from estimating the parametric model can be found in Table 2, and the results for other models with different basis functions are in Tables 3-6. In the latter group of tables, we define $\mathscr{T}_{n}$ as the asymptotic p-values corresponding to $\mathbb{W}_{i}$, $i=0,...,4$. It is well known that nonparametric tests, which rely on their asymptotic normal null distributions, may not perform well with finite samples. As an alternative, we provide bootstrap p-values that are denoted by $\mathscr{T}_{n}^{\ast}$ basing on 150 bootstrap replications. The details of the bootstrap procedure are given in Section 7.1.
Our results show that, in all models, the hypothesis of CRS cannot be rejected. These findings corroborate those in Li et al. (2002). A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
We end this application with some remarks that, in addition to inference on returns to scale, our analysis also reveals other similarities with prior works. Notably, from Table 7, we observe that the parametric method underestimates the returns to scale compared to the semiparametric method. [^1] The returns to scale in a semiparametric model can be computed using $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$. The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses. To illustrate this, take the semiparametric estimates with spatial weights $W_{n}^{d}$ and produce Figure 1 and Figure 2. The former shows that the output elasticity of capital is increasing in management expenses, and the latter shows the output elasticity of labor is decreasing in management expenses.
Intuitively, by the nature of management expenses in production, the elasticities may be expected to increase with both inputs. However, the decreasing elasticities of labor is a known empirical fact in this industry due to the practice of “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest. We also note that the returns to scale is relatively flat in the semiparametric model and is within the 95% bounds for most $z_{i}$, as illustrated in Figure 3. Further discussions on these findings can be found in Li et al. (2002). The results we have described here are not unique to the estimates from $W_{n}^{d}$. Other choices of spatial weights give the same conclusion; we simply selected one to illustrate our findings.
6.2   Testing the impact of the distance to the employment center on housing price
In this subsection, we use the Boston house price dataset, which has been used in several spatial studies, to illustrate another application of our proposed test. The dataset [^2] The related shp file can be downloaded at https://github.com/simonbrewer/geog6000/blob/main/datafiles/boston.tr.zip. Data descriptions are in Table 8. consists of the median value of owner-occupied homes in 506 census tracts in the Boston Standard Metropolitan Statistical Area in 1970 (Harrison Jr and Rubinfeld, 1978). Specifically, we explore how some factors such as the per capita crime rate by town (denoted by CRIM), average number of rooms per dwelling (RM), full-value property-tax rate per \$10,000 dollar (TAX), percentage of lower socioeconomic status population (LSTAT) and index of accessibility to radial highways (RAD) affect the median value of owner-occupied homes in $1000$s (MEDV), and whether the effects of these factors vary over location or not. DIS, which represents the weighted distances between a property and the five Boston employment centers that are associated with higher housing prices due to closer proximity to the city center, will drive the varying coefficient.
The motivation for considering spatial autocorrelation in house prices is straightforward, as housing prices in a specific location or region are commonly influenced by the prices of nearby houses due to factors such as environmental and geographic considerations. [^3] There are also some studies that use the SAR semiparametric model with the Boston house price data, for example, Malikov and Sun (2017) and Li et al. (2019). However, the former considers the covariate $z_{i}$ as NO2, which is more about the impact of environmental pollution on housing prices. The latter concentrates on the geographical weighted regression (GWR) method. Sun et al. (2014) proposed a semiparametric coefficient varying spatial model and derived the distribution theory of their estimator. While they do not have inference results on functions of the random coefficients, they have a model selection procedure to determine the parametric and nonparametric components in their model. We thus take the model they selected as our starting point.
Let $y_{i},x_{i1},x_{i2},p_{i1},p_{i2},p_{i3},z_{i}$ denote the log of MEDV, RAD, LSTAT, CRIM, RM, TAX and DIS respectively. Guided by model selection results, Sun et al. (2014) argue that DIS doesn’t affect LSTAT and RAD. Intuitively, for LSTAT, the geographical distribution of employment opportunities does not necessarily correlate with a community’s socio-economic status; a community might be far from employment centers but could have good public transportation links or a high number of remote work opportunities, mitigating the impact on the community’s lower socioeconomic status. For RAD, the accessibility to highways might depend more on urban planning and infrastructure investment rather than the proximity to employment centers.
The baseline model, based on Sun et al. (2014), is:
$y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
We focus on a more general version of their model, by allowing the spatial lag coefficient to be varying with DIS as well:
$y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
Therefore, we estimate:
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2}\\ &y_{i}=\sum\limits_{j=1}^{n}g(d_{ij})y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4},\end{split}$
and the estimated models are
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4}.\end{split}$ (6.4)
We consider two row-normalized spatial weight matrices. The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017). The second, denoted by $W_{n}^{t}$, indicates whether two locations are in the same tract. For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
From Tables 9-12, we find that $log(DIS)$ indeed affects $p_{i1}$, $p_{i2}$ and $p_{i3}$, and has a significantly nonlinear impact on neighboring house prices. All test statistics show a rejection of the null $H_{0}:\alpha=0$. The bootstrap p-value also confirms this conclusion. In addition, testing with $\mathbb{W}_{4}$ also leads to a rejection of the null $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices. Furthermore, we estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$. From Table 13, we thus find that the parametric method underestimates the spatial coefficient. Our results, which are based on more general models, thus complement those in Sun et al. (2014).
7.   Monte Carlo Simulation
We present a basic set of Monte Carlo results in the main body of the paper. Because our theory covers such a wide range of models, which include several important special cases, we have performed a wide range of simulations for them too. However, the sheer number of tables we have makes it impractical to include them all in the main body of the paper, and we provide references to an online supplementary appendix.
7.1   Basic setting
Taking $n=100,300,500,$ we choose two specifications to generate $y$ with $d_{\delta}=1$:
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$ (7.1)
and
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},i=1,\ldots,n,$ (7.2)
[…]
Kelejian and Prucha (2007) consider the kernel function with a distance measure $d_{ij,m}^{*}$ and a threshold $d_{m}$. In estimation, we use the Epanechnikov kernel function, we choose the Euclidean distance $d_{ij}$ mentioned in $W_{1}$ to represent $d_{ij,m}^{*}$, and $d_{m}$ is considered as the quantile value of 10% of all distances. If $d_{ij,m}^{*}>d_{m}$, $\mathscr{K}\left(\frac{d_{ij,m}^{*}}{d_{m}}\right)=0$. We also consider $d_{\lambda}=2$ with $W_{1}$, $W_{2}$, and $\lambda=(0.5,0.3)^{\prime}$. The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[W_{1}X,W_{2}X,X,\Psi]$. We calculate the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%. We consider 1000 Monte Carlo replications for all experiments and $B=150$ bootstrap replications. We define $\mathscr{S}_{n}$ as the test statistic corresponding to $\mathbb{W}_{i}$ for $i=1,\ldots,4$ and $\mathscr{S}^{\ast}_{n}$ as the bootstrap version of $\mathscr{S}_{n}$.
The bootstrap version of our test, e.g., focusing on the setting described in Section 3, can be carried out by taking these steps:
  • Compute $\widehat{\epsilon}=y-L\hat{\xi}$ under null or alternative and obtain $\widetilde{\epsilon}=(I_{n}-\frac{1}{n}l_{n}l_{n}^{\prime})\widehat{\epsilon}$, where $l_{n}$ is the $n\times 1$ column vector of ones.
  • Sample randomly with replacement $n$ times from elements of $\widetilde{\epsilon}$ to obtain $\tilde{\epsilon}^{*}$.
  • Generate the bootstrap sample $\hat{y}=\left(I_{n}-\sum\limits_{j=1}^{d_{\lambda}}\hat{\lambda}_{j}W_{j}\right)^{+}\left([X,\Psi](\hat{\beta},\hat{\alpha})^{\prime}+\tilde{\epsilon}^{*}\right)$, where $(\cdot)^{+}$denotes the Moore-Penrose generalized inverse.
  • For the bootstrap, compute $\hat{\xi}^{\ast}$ and obtain the bootstrap residual $\hat{\epsilon}^{\ast}=\hat{y}-\hat{L}\hat{\xi}^{\ast}$ to construct $\mathscr{S}_{n}^{\ast}$, where $\hat{L}=[W_{1}\hat{y},\ldots,W_{d_{\lambda}}\hat{y},X,\Psi]$.
  • Repeat $B$ times to obtain the sequence of bootstrap test statistics $\left\{\mathscr{S}_{nj}^{\ast}\right\}_{j=1}^{B}$. Calculate bootstrap p-values $\mathscr{T}_{n}^{\ast}=\frac{1}{B}\sum\limits_{j=1}^{B}\mathbf{1}\left(\mathscr{S}_{n}<\mathscr{S}_{nj}^{\ast}\right)$ and reject the null hypothesis if $\mathscr{T}_{n}^{\ast}$ is smaller than the prescribed nominal level of significance.
As before, we denote by $\mathscr{T}_{n}$ the p-value based on the asymptotic test. From Tables 14-15, we find that the bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases. Specifically, we observe that the pairs $(n,h)=(500,8)$ deliver acceptable sizes under three disturbances. Moreover, we observe that the asymptotic test is powerful against local alternatives as $n$ increases from Tables S1-S2 in the supplement.
We also consider the case of $d_{\lambda}=2$ in the supplement. The size performances [^5] Power performances are close to or equal to 1 in most cases so we omit the power tables but they are available on request. are reported in Tables S3-S4 in the supplement and the finite sample properties are similar to the case of $d_{\lambda}=1$.
7.2   Nonparametric spatial matrix
For the nonparametric spatial matrix, we generate $w_{1ij}=\Phi\left(-b_{ij}\right)\mathbf{1}\left(c_{ij}<0.1\right)$ if $i\neq j$, and $w_{1ii}=0$, where $\Phi(\cdot)$ is the standard normal cdf, $b_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-3,3]$, and $c_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. From this construction, we ensure that $W_{1}$ is sparse with no more than $10\%$ elements being nonzero. Then, $W=W_{1}/1.2\bar{\alpha}(W_{1})$, ensuring the existence of $(I-W)^{-1}$. For $\mathbb{W}_{3}$, we approximate elements in $W$ by $\widehat{w}_{ij}=\sum_{l=0}^{r}\tau_{l}[d_{ij,m}^{*}]^{l}\mathbf{1}(d_{ij,m}^{*}<d_{m})$ if $i\neq j$ with $r=2,4$, where $d_{ij,m}^{*}$ and $d_{m}$ are discussed in Section 7.1, and the IV matrix for $\mathbb{W}_{3}$ is $K_{2}=[X,\Psi]$. We define this DGP setting as (D3).
Table 16 reports the simulation results at the 1%, 5% and 10% asymptotic level with $r=2,4$ and $h=2,4,8$. We find that there is a clear pattern of rejection rates approaching the theoretical level as sample size increases, and sizes are acceptable for $n=500$ and $h=8$. When $r$ increases, the results are similar. From Table S5, the power performances are close to or equal to 1 as $n$ increases. We also consider $W=W_{1}+W_{2}$ which corresponds to $d_{\lambda}=2$. Specifically, we set $w_{ij2}=d_{m}\exp(-d_{ij,m}^{*}/d_{m})$ if $d_{ij,m}^{*}\leq d_{m}$ and $w_{ij2}=0$ if $d_{ij,m}^{*}>d_{m}$. From Table S6 in the supplement, a similar pattern for size is observed.
7.3   Varying spatial coefficient
We now consider varying $\lambda$ such that $\lambda=\lambda(z_{i})$. When $d_{\lambda}=1$, we set $\lambda(z_{i})=\sin(0.75\pi\times z_{i})$, when $d_{\lambda}=2$, we set $\lambda_{1}(z_{i})=\sin(0.38\pi\times z_{i})$ and $\lambda_{2}(z_{i})=\cos(0.37\pi\times z_{i})$. Others are same as those in Section 7.1. For estimation, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$. Then, $h_{1il}=y_{i}\sum_{j=1}^{n}w_{1ij}\phi_{jl}(z)$ and $h_{2il}=y_{i}\sum_{j=1}^{n}w_{2ij}\phi_{jl}(z)$. The IV matrix is $K_{3}=[H_{1}X,X,\Psi]$ for $d_{\lambda}=1$ and $K_{3}=[H_{1}X,H_{2}X,X,\Psi]$ for $d_{\lambda}=2$. We define this DGP setting as (D4).
Tables 17 and 18 report the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%. Sizes are acceptable as $n$ and $h$ increase, especially, the trigonometric approximation performs slightly better than polynomial approximations for Chi-square distribution when testing $\lambda=0$. Bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases. When $d_{\lambda}=2$, the results are similar, as shown in Tables S7-S8 in the supplement.
From Tables S9 and S10 in the supplement, power improves as $h$ increases for all designs, and the test for $\lambda$ and $\delta$ have similar results. As for $d_{\lambda}=2$, we find that our proposed test statistics still perform well as both $n$ and $h$ increase.
7.4   Additional simulation results in Online Supplement
In a range of experiments reported in the online supplement, we study how large $n$ with $h$ increasing to see how size changes. So we set $n=1000$ with $h=12,16$, $d_{\lambda}=1,2$ and $d_{\delta}=1$. From Table S11 of the supplement, we find that as $h$ increases, the performance of size improves.
We also consider tests of equality of varying coefficients and CRS hypothesis, where $d_{\delta}=2$ such that
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+\epsilon_{i},$ (7.3)
and
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},$ (7.4)
for $i=1,\ldots,n$. Our null of interest when testing equality is $H_{0}^{true}:\delta_{1}(z)=\delta_{2}(z),z\in\mathcal{Z},$ while for CRS it is $H_{0}^{true}:\delta_{1}(z)+\delta_{2}(z)=1,z\in\mathcal{Z}.$ In the first case, we reparameterize with $\theta(z)=\delta_{1}(z)-\delta_{2}(z)$, and the model is $y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i1}+p_{i2})\delta_{2}(z_{i})+\epsilon_{i}.$ In the second case, we reparameterize with $\theta(z)=\delta_{1}(z)+\delta_{2}(z)-1$ and the model is $y_{i}-p_{i1}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i2}-p_{i1})\delta_{2}(z_{i})+\epsilon_{i}.$
The results are included in the supplement. Taking CRS as an example, from Tables S12-S13, bootstrap tests are closer to the nominal levels than the asymptotic tests although the sizes of both types of tests improve generally as the sample size increases. For the nonparametric spatial matrix, from Table S14, there is a clear pattern of rejection rates approaching the theoretical level as sample size increases. The results of testing equality are similar, as shown in Tables S15-S17.
Appendix
Appendix A General Central Limit Theorem
Introduce an $n\times n$ matrix $\mathscr{A}$, which is symmetric, idempotent and has rank $J\rightarrow\infty$.
Assumption A.
Assume that
$\mathscr{A}=B^{\prime}\mathscr{S}B,$ (A.1)
where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7.
(b) The entries $s_{ij}$ of the symmetric matrix $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$.
(c) $\mathscr{S}$ satisfies 1.
Theorem A.1.
Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 10. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
[…]
Proof.
The proof is identical to that of Theorem 4.3 in Gupta and Qu (2024) and is omitted. ∎
Appendix B Proofs of Theorems
Proof of Theorem 2.1.
Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$. Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Then we are going to investigate
$\begin{split}&n(r_{\delta}+\epsilon)^{\prime}P_{K_{1}}L(L^{\prime}P_{K_{1}}L)^{-1}R^{\prime}\mathscr{D}_{1}^{-1}R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)\\ =&\epsilon^{\prime}\mathcal{M}_{1}\epsilon+(A_{1}+A_{2})\end{split}$ (B.1)
where $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$ with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, and $A_{1}=r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$ and $A_{2}=2r_{\delta}^{\prime}\mathcal{M}_{1}\epsilon$. Now, $|A_{1}|\leq\|r_{\delta}\|^{2}\|\mathcal{M}_{1}\|=o_{p}(1)$, since $E\|r_{\delta}\|^{2}=o(1)$, and $\|\mathcal{M}_{1}\|=O_{p}(1)$ by Assumptions 1 and 4. A similar argument holds for $A_{2}$ because it has (conditionally) zero mean and variance $4r_{\delta}^{\prime}\mathcal{M}_{1}E(\epsilon\epsilon^{\prime})\mathcal{M}_{1}r_{\delta}=4r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$, by idempotence of $\mathcal{M}_{1}$. Thus $\frac{A_{1}}{\sqrt{d_{\alpha}}}$ and $\frac{A_{2}}{\sqrt{d_{\alpha}}}$ are both $o_{p}(1)$ and establishing the theorem. ∎
Proof of Theorem 2.2.
This is a special case of Theorem 3.2 and therefore omitted. ∎
Proof of Lemma 3.1.
This is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
Proof of Theorem 3.1.
$\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(L^{\prime}{P_{K_{1}}}L)^{-1}L^{\prime}{P_{K_{1}}}(r_{\delta}+Bv)$ and under $H_{0}:\alpha=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{2}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{2}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.2)
where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, ^$\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and ^$\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
Proof of Theorem 3.2.
The proof follows by setting $J=J_{1}$, $\mathscr{A}=\mathcal{M}_{2}$ and $\mathscr{S}=\frac{1}{n}K_{1}\mathcal{V}_{2}K_{1}^{\prime}$ in Theorem A.1. Checking Assumption A follows exactly as in the proof of Theorem 5.2 below and is omitted. ∎
Proof of Lemma 4.1.
This proof is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
Proof of Theorem 4.1.
Define $r_{G}=(r_{1G},\cdots,r_{nG})^{\prime}$ and $r_{G\delta}=r_{G}+r_{\delta}$. Then, we have $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(F^{\prime}{P_{K_{2}}}F)^{-1}F^{\prime}{P_{K_{2}}}(r_{G\delta}+Bv)$ and $\big{\|}r_{G\delta}\big{\|}^{2}=o_{p}(1)$ by Assumption 12. Under $H_{0}:\alpha=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{3}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{3}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.3)
where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, ^$\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$ with $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and ^$\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
Proof of Theorem 4.2.
The proof is similar to that of Theorem 5.2 and omitted. ∎
Proof of Lemma 5.1.
Define $r_{\lambda}=(r_{1\lambda},...,r_{n\lambda})^{\prime}$. First, note that
$\hat{\gamma}-\gamma=(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}u=(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta}+\epsilon),$ (B.4)
and
$\begin{split}\mathrm{E}\left\|n^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta})\right\|^{2}&\leq n^{-2}\left\|K_{3}\right\|^{2}\mathrm{E}\left\|r_{\lambda}+r_{\delta}\right\|^{2}\\ &\leq Cn^{-1}\overline{\alpha}(n^{-1}K_{3}^{\prime}K_{3})\left(\mathrm{E}\left\|r_{\lambda}\right\|^{2}+\mathrm{E}\left\|r_{\delta}\right\|^{2}\right)=o(1),\end{split}$ (B.5)
because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17.
Next,
$\mathrm{E}\left\|n^{-1}K_{3}^{\prime}\epsilon\right\|^{2}$$\leq$ $n^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}\mathrm{E}\epsilon_{i}^{2}+n^{-2}\sum_{i\neq j}\left\|k_{i}\right\|\left\|k_{j}\right\|\left|\mathrm{E}\epsilon_{i}\epsilon_{j}\right|$ (B.6)
$=$ $n^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}\sum_{r}b_{ir}^{2}+2n^{-2}\sum_{j}\left\|k_{j}\right\|\sum_{i<j}\left\|k_{i}\right\|\sup_{r}\left|b_{jr}\right|\left(\sum_{r}\left|b_{ir}\right|\right)$
$\leq$ $Cn^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}+Cn^{-2}\sup_{j}\left\|k_{j}\right\|\left(\sup_{r}\sum_{j}\left|b_{jr}\right|\right)\sum_{i<j}\left\|k_{i}\right\|$
$=$ $O(d_{\gamma}/n),$
by Assumptions 7 and 15.
Combining Assumption 17, (B.4), (B.5) and (B.6), we obtain
$\left\|\hat{\gamma}-\gamma\right\|=O_{p}(\sqrt{d_{\gamma}/n}).$ (B.7)
Now, observe that
$\hat{\Xi}_{r,s}-\Xi_{r,s}=a_{1,rs}+a_{2,rs}+a_{3,rs},$ (B.8)
where
$a_{1,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}(\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j})\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),$ (B.9)
$a_{2,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}(\epsilon_{i}\epsilon_{j}-\sigma_{ij})\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),$ (B.10)
$a_{3,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}\sigma_{ij}\left(\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right)-1\right),$ (B.11)
where recall that $\sigma_{ij}$ is a typical element of $\Sigma$.
We will establish $a_{l,rs}=o_{p}(1),l=1,2,3$. To begin, write $r_{i}=r_{i\lambda}+r_{i\delta}$, $g_{i}^{\prime}$ for the $i$-th row of $G$ and note that
$\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j}=g_{i}^{\prime}(\hat{\gamma}-\gamma)g_{j}^{\prime}(\hat{\gamma}-\gamma)+g_{i}^{\prime}(\hat{\gamma}-\gamma)r_{j}+r_{i}r_{j}+g_{j}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{i}+g_{i}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{j}+\epsilon_{i}r_{j}+\epsilon_{j}r_{i}.$ (B.12)
Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$.
Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, also using (B.7),
$a_{11,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|^{2}\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}^{2}}{n}\right),$
$a_{12,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}}{n}\right),$
$a_{13,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|r_{i}\right|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{1}{n}\right),$
$a_{14,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
$a_{15,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
$a_{16,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{i}\right|\left|r_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
$a_{17,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{j}\right|\left|r_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007). Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$. By Assumption 9(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$. We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$.
$a_{2,rs}=o_{p}(1)$ and $a_{3,rs}=o_{p}(1)$ are established exactly as the negligibility of $b_{rs,n}$ and $c_{rs,n}$ is established in the proof of Theorem 1 of Kelejian and Prucha (2007). This concludes the proof. ∎
Proof of Theorem 5.1.
Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$. Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big{\|}r_{\lambda\delta}\big{\|}^{2}=o(1)$ by Assumption 16. Under $H_{0}:\vartheta=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{4}v+A_{11}+A_{12}+2A_{21}+2A_{22})\end{split}$ (B.13)
where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, ^$\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$ with $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and ^$\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$. Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17. Thus, $\big{\|}\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big{\|}\leq\big{\|}\widehat{\mathcal{U}}_{3}^{-1}\big{\|}\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\big{\|}\mathcal{U}_{3}^{-1}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$. Similarly, we have ^$\big{\|}\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big{\|}=o_{p}(1)$, and ^$\big{\|}\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big{\|}=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
Thus, we have
$\frac{1}{\sqrt{d_{\vartheta}}}\big{|}v^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v\big{|}\leq Cr_{n}\cdot O_{p}\left(d_{\vartheta}^{-1/2}\right)=o_{p}(1)$ (B.14)
since $\mathrm{E}(v_{i}^{2})=1$.
Similar to (B.14) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\alpha}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\alpha}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\alpha}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\alpha}}}|A_{22}|=o_{p}(1)$. Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible.
Proof of Theorem 5.2.
(i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1. We now proceed as in the proof Theorem 4.3 of Gupta and Qu (2024) and note that the entries of $\mathscr{S}$ are,
$s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ (B.15)
where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so
$\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.16)
uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above.
Similarly, we also observe that
$\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.17)
uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 17. The conclusion now follows by Theorem A.1.
(ii) This proof is similar to that of Theorem 4.4 in Gupta and Qu (2024).
(iii) We prove all local alternatives here. We define $H_{\ell i}\equiv H_{\ell i,n}$, with the corresponding statistic being $\mathbb{W}_{i}^{*}$, for $i=1,2,3,4$. For $\mathbb{W}_{1}^{*}$, we consider $\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$ with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. Under $H_{\ell 1}$, we have
$\mathbb{W}_{1}^{*}=\mathbb{W}_{1}+\frac{1}{\sqrt{2d_{\alpha}}}d_{\alpha}^{\frac{1}{2}}\mathcal{A}_{1}(\frac{1}{n}\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{-1}+o_{p}(1),$
where $\mathcal{A}_{1}=\frac{1}{n}\nu_{1n}^{\prime}\mathscr{D}_{1}^{-1}\nu_{1n}$. We choose $\Gamma_{1n}=\mathscr{D}_{1}^{-1}$ and then finish the proof. For $\mathbb{W}_{4}^{*}$, we consider $\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{\frac{1}{2}}$. Under $H_{\ell 4}$, we have
$\mathbb{W}_{4}^{*}=\mathbb{W}_{4}+\frac{1}{\sqrt{2d_{\vartheta}}}d_{\vartheta}^{\frac{1}{2}}\mathcal{A}_{4}(\frac{1}{n}\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{-1}+o_{p}(1),$
where $\mathcal{A}_{4}=\frac{1}{n}\nu_{4n}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\nu_{4n}$, and $\big{\|}\frac{1}{n}\nu_{4n}^{\prime}(\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1})\nu_{4n}\big{\|}=o_{p}(1)$ by (5.7). We choose $\Gamma_{4n}=\widehat{\mathscr{D}}_{4}^{-1}$ and then finish the proof. The proofs of $\mathbb{W}_{2}^{*}$ and $\mathbb{W}_{3}^{*}$ are similar. ∎
Appendix C Empirical results
Table 1: Data descriptions in Section 6.1, logarithm form.
Variables Descriptions Mean Min Max Sd
$y$ Value added 11.539 3.611 19.216 1.847
$p_{1}$ Capital assets 10.280 1.609 19.160 1.980
$p_{2}$ Average number of employ 5.647 2.021 11.865 1.130
$z$ Management expense 7.707 2.303 16.139 1.853
Table 2: Results of baseline model according to (6.1) in Section 6.1.
                 $\delta_{0}$                  3.313***
                 $\delta_{1}$                  0.554***
                 $\delta_{2}$                  0.402***
                 $\beta_{1}$                  0.224
                 $\beta_{2}$                  0.018
                 $H_{0}:\delta_{1}+\delta_{2}-1=0$
                 t-test                  -0.238
                 p-value                  0.812
                 $n$                  7355
                 The standard errors of the estimators
                 are reported in parentheses.
                  $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 3: Estimation and testing results in Section 6.1 using polynomial series with $h=2$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.189* 0.130** 0.224*** 0.256*** 0.451*** 0.457*** 0.499***
(0.112) (0.062) (0.014) (0.081) (0.157) (0.042) (0.028)
$\tau_{2}/\mu_{2}$0.095 0.405*** 0.484***
(0.126) (0.047) (0.030)
$\beta_{0}$ 1.073** 0.167* 0.153** 0.167*** 0.153*** -0.485*** -0.093*** 0.077***
(0.514) (0.100) (0.076) (0.016) (0.020) (0.013) (0.004) (0.003)
$\alpha_{11}$ 1.254** -0.036*** -0.032*** -0.036 -0.032 -0.023 0.005 0.004
(0.551) (0.007) (0.004) (0.027) (0.035) (0.975) (0.033) (0.026)
$\alpha_{12}$ 0.375 -0.015 0.017 -0.015 0.017 -0.032 0.015*** 0.027***
(3.048) (0.035) (0.019) (0.010) (0.036) (0.050) (0.002) (0.002)
$\alpha_{21}$0.194 -0.004 -0.004 -0.004 -0.004 -0.002 0.000 0.003
(0.786) (0.020) (0.033) (0.005) (0.012) (0.111) (0.001) (0.002)
$\alpha_{22}$0.023 0.092*** 0.088*** 0.092*** 0.088*** 0.032*** -0.017*** 0.012***
(0.077) (0.013) (0.013) (0.004) (0.029) (0.008) (0.000) (0.000)
$\alpha_{31}$0.046 0.031 0.019 0.031 0.019* 0.011 0.001 0.003
(0.891) (0.017) (0.010) (0.022) (0.011) (0.136) (0.018) (0.014)
$\alpha_{32}$0.062 0.004 0.005 0.004 0.005 0.059 -0.033 0.026
(0.092) (0.023) (0.023) (0.001) (0.004) (0.060) (0.031) (0.051)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 1.593 3.311 3.280 1.419 4.020 0.226 1.716 2.589
$\mathscr{T}_{n}$ 0.451 0.191 0.194 0.492 0.134 0.893 0.424 0.274
$\mathscr{T}_{n}^{*}$ 0.767 0.213 0.167 0.413 0.167 0.867 0.433 0.213
$H_{0}:\mu=0$ 68.365 92.106
$\mathscr{T}_{n}$ $<$0.001 $<$0.001
$\mathscr{T}_{n}^{*}$ 0.007 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 4: Estimation and testing results in Section 6.1 using trigonometric series with $h=2$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.065 0.162* 0.065** 0.162*** 0.582*** 0.607*** 0.297***
(0.052) (0.087) (0.033) (0.014) (0.098) (0.166) (0.088)
$\tau_{2}/\mu_{2}$0.276*** 0.147** 0.494***
(0.041) (0.074) (0.088)
$\beta_{0}$ 1.111*** -0.381*** -0.382** -0.381*** -0.382*** -0.187*** 0.747*** 1.076***
(0.020) (0.136) (0.162) (0.042) (0.021) (0.012) (0.245) (0.380)
$\alpha_{11}$ 0.949*** 0.167*** 0.160*** 0.167*** 0.160*** 0.125*** -0.253*** -0.024***
(0.150) (0.023) (0.025) (0.014) (0.012) (0.016) (0.025) (0.004)
$\alpha_{12}$ 0.861*** -0.049 -0.028** -0.049 -0.028 -0.124*** 0.146*** 0.034***
(0.151) (0.033) (0.012) (0.035) (0.032) (0.008) (0.040) (0.007)
$\alpha_{21}$ 0.001 0.0096 0.094 0.096*** 0.094*** 0.014*** -0.091*** 0.021***
(0.007) (0.112) (0.124) (0.003) (0.003) (0.002) (0.019) (0.005)
$\alpha_{22}$ 0.002 0.026 0.024 0.026*** 0.024*** -0.018*** 0.020*** 0.009***
(0.007) (0.069) (0.043) (0.004) (0.003) (0.002) (0.005) (0.003)
$\alpha_{31}$ 0.017 0.085 0.068 0.085 0.068 0.001 -0.020 0.019
(0.018) (0.246) (0.191) (0.105) (0.064) (0.001) (0.017) (0.012)
$\alpha_{32}$ 0.009 0.033 0.025 0.033 0.025 -0.003 0.047 0.021
(0.022) (0.090) (0.084) (0.002) (0.002) (0.003) (0.007) (0.006)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 0.777 1.178 0.877 1.086 0.943 1.575 1.707 2.169
$\mathscr{T}_{n}$ 0.678 0.555 0.645 0.581 0.624 0.455 0.426 0.338
$\mathscr{T}_{n}^{*}$ 0.660 0.533 0.680 0.553 0.600 0.467 0.267 0.333
$H_{0}:\mu=0$ 321.305 406.511
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.027
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 5: Estimation and testing results in Section 6.1 using polynomial series with $h=4$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.330*** 0.265** 0.330*** 0.065*** 0.444*** 0.236*** 0.313***
(0.095) (0.134) (0.019) (0.024) (0.135) (0.043) (0.0969)
$\tau_{2}/\mu_{2}$0.013 0.219*** 0.330***
(0.045) (0.052) (0.064)
$\tau_{3}/\mu_{3}$ 0.025** 0.203** 0.258**
(0.012) (0.082) (0.126)
$\tau_{4}/\mu_{4}$0.088 0.189* 0.096*
(0.068) (0.110) (0.051)
$\beta_{0}$ 1.866 0.040*** 0.005*** 0.040*** 0.005 -0.036 0.075 -0.278***
(4.659) (0.011) (0.001) (0.011) (0.023) (0.159) (0.221) (0.062)
$\alpha_{11}$ 1.471 0.002*** 0.000*** 0.002 0.000 0.000 0.003 -0.006
(4.108) (0.000) (0.000) (1.759) (2.686) (10.471) (26.990) (3.119)
$\alpha_{12}$1.067 0.005*** 0.002*** 0.005 0.002 0.010 0.009 -0.005
(5.407) (0.001) (0.000) (0.329) (0.588) (2.398) (5.235) (0.903)
$\alpha_{13}$ 1.186 0.013*** 0.007*** 0.013 0.007 0.035 0.014 0.010
(8.484) (0.001) (0.002) (0.028) (0.053) (0.232) (0.466) (0.100)
$\alpha_{14}$0.020 0.121*** -0.004 0.121*** -0.004*** -0.174*** 0.180*** -0.046***
(0.288) (0.021) (0.042) (0.001) (0.002) (0.008) (0.016) (0.004)
$\alpha_{21}$0.032 0.015*** 0.003*** 0.015 0.003 0.000 0.006 -0.004
(0.676) (0.004) (0.001) (0.180) (0.462) (0.861) (3.899) (0.134)
$\alpha_{22}$0.029 0.041*** 0.020*** 0.041 0.020 0.004 0.013 0.022
(2.100) (0.008) (0.006) (0.031) (0.100) (0.190) (0.642) (0.071)
$\alpha_{23}$ 0.003 -0.010 0.092*** -0.010*** 0.092*** -0.011 -0.018 0.019*
(0.213) (0.019) (0.026) (0.002) (0.009) (0.017) (0.046) (0.010)
$\alpha_{24}$ $0$0.005*** -0.003** -0.005*** -0.003*** 0.005*** -0.001 -0.001***
(0.007) (0.000) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000)
$\alpha_{31}$ 0.005 -0.005 -0.002*** -0.005 -0.002 -0.001 -0.002 0.006
(0.012) (0.061) (0.001) (0.344) (0.556) (1.561) (3.200) (0.827)
$\alpha_{32}$0.009 -0.008 -0.009 -0.008 -0.009 -0.009 -0.002 0.004
(0.042) (0.042) (0.053) (0.068) (0.124) (0.428) (0.543) (0.297)
$\alpha_{33}$ 0.007 0.059 -0.041 0.059 -0.041 -0.032 0.029 0.031
(0.032) (0.096) (0.113) (0.046) (0.033) (0.041) (0.040) (0.031)
$\alpha_{34}$0.002 0.001 0.016*** 0.001 0.016 0.008 -0.004 -0.003
(0.026) (0.073) (0.157) (0.006) (0.062) (0.011) (0.083) (0.044)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 0.075 1.707 1.464 2.401 1.808 0.030 0.098 0.385
$\mathscr{T}_{n}$ 0.963 0.426 0.481 0.301 0.405 0.985 0.952 0.825
$\mathscr{T}_{n}^{*}$ 0.947 0.440 0.500 0.287 0.467 0.933 0.887 0.793
$H_{0}:\mu=0$ 101.231 122.398
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 6: Estimation and testing results in Section 6.1 using trigonometric series with $h=4$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.224*** 0.256*** 0.224*** 0.256*** 0.196** 0.123* 0.136***
(0.040) (0.083) (0.014) (0.081) (0.098) (0.067) (0.033)
$\tau_{2}/\mu_{2}$ 0.261* 0.298** 0.269**
(0.136) (0.143) (0.110)
$\tau_{3}/\mu_{3}$ 0.029* 0.293** 0.299**
(0.018) (0.145) (0.140)
$\tau_{4}/\mu_{4}$0.011 0.107* 0.138*
(0.124) 0.026** 0.035***
$\beta_{0}$ 1.108*** 0.060** 0.004** 0.060*** 0.004*** -0.067*** (0.008) (0.009)
(0.021) (0.031) (0.002) (0.009) (0.001) (0.025) (0.055) (0.077)
$\alpha_{11}$ 0.995*** 0.167*** 0.153*** 0.167*** 0.153*** -0.415*** 0.143*** 0.150***
(0.162) (0.010) (0.011) (0.016) (0.020) (0.014) (0.016) (0.014)
$\alpha_{12}$ 0.904*** -0.036*** -0.032*** -0.036 -0.032 0.735*** -0.032 -0.022
(0.164) (0.005) (0.006) (0.027) (0.035) (0.019) (0.023) (0.021)
$\alpha_{13}$ 0.084 -0.015** 0.017*** -0.015 0.017 -0.198*** -0.003 -0.026***
(0.158) (0.007) (0.005) (0.010) (0.036) (0.006) (0.008) (0.006)
$\alpha_{14}$0.251 -0.004*** -0.004 -0.004 -0.004 -0.160*** -0.002 -0.019***
(0.160) (0.001) (0.003) (0.005) (0.012) (0.007) (0.005) (0.004)
$\alpha_{21}$ 0.001 0.092*** 0.088*** 0.092*** 0.088*** -0.006*** 0.037*** 0.037***
(0.007) (0.015) (0.016) (0.004) (0.029) (0.001) (0.003) (0.003)
$\alpha_{22}$ 0.001 0.031*** 0.019*** 0.031*** 0.019*** -0.037*** 0.013*** 0.014***
(0.008) (0.012) (0.003) (0.002) (0.003) (0.003) (0.002) (0.002)
$\alpha_{23}$ 0.000 0.004 0.035 0.004*** 0.035*** 0.014*** 0.007*** -0.003***
(0.007) (0.021) (0.022) (0.001) (0.004) (0.001) (0.001) (0.001)
$\alpha_{24}$ 0.005 0.003 0.006 0.003*** 0.006*** 0.003*** 0.002*** 0.002***
(0.008) (0.004) (0.008) (0.000) (0.001) (0.001) (0.000) (0.000)
$\alpha_{31}$ 0.018 0.087 0.068 0.087 0.068 -0.027 0.041 0.041
(1.320) (1.065) (1.012) (0.127) (0.126) (0.863) (0.079) (0.166)
$\alpha_{32}$ 0.020 0.029 0.025 0.029 0.025 -0.081 0.013 0.006
(0.685) (2.392) (2.003) (0.268) (0.212) (0.062) (0.091) (0.099)
$\alpha_{33}$ 0.003 0.028 0.005 0.028 0.005 0.019 0.010 0.009
(0.331) (5.691) (4.001) (0.316) (0.305) (0.891) (1.063) (0.085)
$\alpha_{34}$ 0.005 -0.022 -0.020 -0.022 -0.020 0.019 -0.011 -0.013
(0.369) (0.355) (0.093) (0.287) (0.365) (1.605) (0.062) (0.076)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 0.290 1.848 1.497 0.038 1.443 0.034 0.375 1.022
$\mathscr{T}_{n}$ 0.865 0.397 0.473 0.981 0.486 0.983 0.829 0.600
$\mathscr{T}_{n}^{*}$ 0.873 0.393 0.453 0.953 0.467 0.927 0.833 0.580
$H_{0}:\mu=0$ 86.301 90.233
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 7: Comparison of parametric method and semiparametric method in Section 6.1.
No $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
parametric $\hat{\delta}_{1}$ 0.5536 0.5438 0.5330 0.5989 0.5996 0.5960
$\hat{\delta}_{2}$ 0.4022 0.3794 0.3725 0.3145 0.3226 0.3099
$\hat{\delta}_{1}$+$\hat{\delta}_{2}$ 0.9558 0.9232 0.9055 0.9134 0.9222 0.9060
polynomial, $h=2$ $\hat{\delta}_{1}$ 0.6228 0.6488 0.6179 0.6263 0.6051 0.6108
$\hat{\delta}_{2}$ 0.3664 0.3212 0.3691 0.2930 0.3502 0.3220
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9892 0.9699 0.9869 0.9194 0.9553 0.9328
trigonometric,$h=2$ $\hat{\delta}_{1}$ 0.6233 0.6257 0.6149 0.6284 0.6298 0.6806
$\hat{\delta}_{2}$ 0.3517 0.3099 0.3381 0.3040 0.3377 0.2816
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9750 0.9356 0.9530 0.9324 0.9675 0.9621
polynomial, $h=4$ $\hat{\delta}_{1}$ 0.6840 0.6390 0.5432 0.6077 0.6213 0.6342
$\hat{\delta}_{2}$ 0.3089 0.3395 0.4345 0.3252 0.3471 0.3025
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9929 0.9785 0.9777 0.9330 0.9683 0.9368
trigonometric,$h=4$ $\hat{\delta}_{1}$ 0.6792 0.6896 0.6814 0.6152 0.6289 0.6569
$\hat{\delta}_{2}$ 0.3121 0.2702 0.3002 0.3043 0.3649 0.3199
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9913 0.9598 0.9816 0.9195 0.9938 0.9768
$n$ 7355
The result in the first column does not consider any spatial weight matrix.
Figure 1: Comparison of parametric method and semiparametric method in Section 6.1, capital.
Figure 2: Comparison of parametric method and semiparametric method in Section 6.1, labor.
Figure 3: Comparison of parametric method and semiparametric method in Section 6.1, CRS.
Table 8: Data descriptions in Section 6.2.
Variables Descriptions Mean Min Mean SD
$y$ log(MEDV) 3.035 1.609 3.912 0.408
$x_{1}$ log(LSTAT) 2.371 0.548 3.637 0.601
$x_{2}$ log(RAD) 1.868 0 3.178 0.875
$p_{1}$ log(CRIM) -0.780 -5.064 4.488 2.162
$p_{2}$ log(RM) 1.832 1.270 2.172 0.112
$p_{3}$ log(TAX) 5.931 5.231 6.567 0.396
$z$ log(DIS) 1.188 0.122 2.495 0.540
Table 9: Estimation and testing results in Section 6.2 using polynomial series with $h=2$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.238*** 0.219*** 0.238*** 0.219*** 0.214** 0.512*** 0.468***
(0.049) (0.044) (0.028) (0.022) (0.084) (0.073) (0.089)
$\tau_{2}/\mu_{2}$ 0.330** -0.356*** -0.311***
(0.129) (0.049) (0.057)
$\beta_{0}$ 4.081*** 3.987*** 4.081*** 3.987*** -0.003 0.132*** 0.022*
(0.267) (0.272) (0.147) (0.156) (0.006) (0.018) (0.012)
$\beta_{1}$0.432*** -0.432*** -0.432*** -0.432*** -0.143* -0.053*** -0.162
(0.025) (0.024) (0.027) (0.025) (0.075) (0.004) (0.175)
$\beta_{2}$ 0.166*** 0.154*** 0.166*** 0.154*** 0.086* 0.025*** 0.044**
(0.024) (0.023) (0.018) (0.017) (0.049) (0.001) (0.020)
$\alpha_{11}$0.044*** -0.041*** -0.044*** -0.041* -0.056* -0.033** -0.017***
(0.006) (0.006) (0.006) (0.024) (0.030) (0.001) (0.004)
$\alpha_{12}$ 0.028*** 0.025*** 0.028*** 0.025*** 0.043* 0.022*** 0.014***
(0.004) (0.004) (0.005) (0.005) (0.022) (0.001) (0.003)
$\alpha_{21}$ 0.042** 0.046** 0.042** 0.046*** 0.115** -0.001 -0.068***
(0.019) (0.018) (0.019) (0.006) (0.051) (0.002) (0.011)
$\alpha_{22}$0.012 -0.015 -0.012 -0.015 0.136*** 0.008*** 0.035***
(0.012) (0.011) (0.009) (0.022) (0.049) (0.001) (0.004)
$\alpha_{31}$$0**$027*** -$0**$027*** $0*$005*** -0.019***
(0.007) (0.007) (0.006) (0.002) (0.030) (0.002) (0.003)
$\alpha_{32}$ 0.076* 0.070* 0.076*** $0**$084*** 0.082*** 0.101***
(0.040) (0.040) (0.003) (0.004) (0.015) (0.021) (0.011)
Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 183.122 91.459 123.439
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
$H_{0}:\mu=0$ 32.336 35.141
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.013
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 10: Estimation and testing results in Section 6.2 using trigonometric series with $h=2$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.254*** 0.225*** 0.254*** 0.225*** 0.259*** 0.177*** 0.121**
(0.049) (0.044) (0.028) (0.023) (0.069) (0.027) (0.057)
$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
(0.149) (0.040) (0.083)
$\beta_{0}$ 3.527*** 3.547*** 3.527*** 3.547*** 0.001 0.567*** 0.675***
(0.236) (0.229) (0.175) (0.139) (0.009) (0.091) (0.028)
$\beta_{1}$0.410*** -0.415*** -0.410*** -0.415*** -0.235*** -0.396*** -0.456***
(0.024) (0.023) (0.023) (0.024) (0.064) (0.021) (0.032)
$\beta_{2}$ 0.156*** 0.146*** 0.156*** 0.146*** 0.170*** 0.154*** 0.170***
(0.024) (0.023) (0.017) (0.017) (0.031) (0.018) (0.020)
$\alpha_{11}$0.056* -0.035* -0.056*** -0.035*** -0.024* -0.006** -0.007*
(0.040) (0.021) (0.003) (0.003) (0.012) (0.003) (0.004)
$\alpha_{12}$0.023*** -0.022*** -0.023*** -0.022*** -0.163*** -$0**$024***
(0.005) (0.005) (0.006) (0.005) (0.027) (0.005) (0.005)
$\alpha_{21}$ 0.054*** 0.049*** 0.054*** 0.049*** 0.241** 0.037*** 0.048***
(0.010) (0.010) (0.008) (0.007) (0.097) (0.009) (0.011)
$\alpha_{22}$0.016 -0.007 -0.016*** -0.007*** 0.212*** 0.106** 0.132**
(0.018) (0.018) (0.003) (0.002) (0.026) (0.052) (0.065)
$\alpha_{31}$0.027*** -0.023*** -0.027*** -0.023*** -0.017 -0.025*** -0.030***
(0.004) (0.004) (0.002) (0.002) (0.036) (0.002) (0.002)
$\alpha_{32}$ 0.010* 0.009* 0.010*** 0.009** 0.622*** 0.053*** 0.065***
(0.006) (0.005) (0.004) (0.004) (0.015) (0.004) (0.005)
Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 38.897 34.423 229.115 325.327 266.301 175.332 144.982
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.020 0.007 0.007 0.020 0.007 0.013 0.007
$H_{0}:\mu=0$ 28.883 21.326
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 11: Estimation and testing results in Section 6.2 using polynomial series with $h=4$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.239*** 0.269*** 0.239** 0.269*** 0.162** 0.169** 0.169***
(0.050) (0.043) (0.095) (0.027) (0.069) (0.080) (0.019)
$\tau_{2}/\mu_{2}$ 0.175*** -0.363** -0.363***
(0.042) (0.177) (0.080)
$\tau_{3}/\mu_{3}$ 0.241*** 0.345** 0.345***
(0.062) (0.147) (0.113)
$\tau_{4}/\mu_{4}$ 0.082** -0.122** -0.123**
(0.035) (0.060) (0.053)
$\beta_{0}$ 3.285*** 2.925*** 3.285 2.925*** -$0**$092* -0.440**
(0.464) (0.444) (7.140) (0.347) (0.003) (0.047) (0.211)
$\beta_{1}$0.399*** -0.386*** -0.399*** -0.386*** -0.516*** 0.021* -0.105***
(0.026) (0.025) (0.039) (0.025) (0.092) (0.012) (0.013)
$\beta_{2}$ 0.158*** 0.137*** 0.158*** 0.137*** 0.313*** 0.019* 0.052***
(0.025) (0.025) (0.023) (0.018) (0.057) (0.011) (0.013)
$\alpha_{11}$ 0.041 0.031 0.041 0.031 0.048* 0.034** 0.006**
(0.053) (0.050) (0.067) (0.039) (0.027) (0.015) (0.003)
$\alpha_{12}$0.224* -0.185* -0.224* -0.185** -0.120*** -0.072** -0.012*
(0.118) (0.112) (0.135) (0.086) (0.024) (0.034) (0.006)
$\alpha_{13}$ 0.228*** 0.193** 0.228*** 0.193*** -0.055* 0.050*** 0.080*
(0.087) (0.083) (0.087) (0.062) (0.028) (0.016) (0.046)
$\alpha_{14}$0.064*** -0.055*** -0.064*** -0.055*** 0.058*** -0.011** -0.018*
(0.021) (0.020) (0.018) (0.015) (0.018) (0.005) (0.010)
$\alpha_{21}$0.706*** -$0**$706*** -$0**$073 0.019 0.003**
(0.240) (0.226) (0.226) (0.114) (0.073) (0.011) (0.001)
$\alpha_{22}$ 1.574*** 1.596*** 1.574*** 1.596*** -0.014 -0.026* -0.006**
(0.484) (0.453) (0.460) (0.229) (0.045) (0.015) (0.002)
$\alpha_{23}$1.040*** -1.078*** -1.040*** -1.078*** -0.020 0.056*** 0.030***
(0.312) (0.291) (0.295) (0.147) (0.016) (0.014) (0.009)
$\alpha_{24}$ 0.217*** 0.230*** 0.217*** 0.230*** -0.057* $0**$045**
(0.065) (0.060) (0.060) (0.030) (0.030) (0.003) (0.018)
$\alpha_{31}$ 0.262*** 0.264*** 0.262*** 0.264*** 0.068*** -$0**$017***
(0.084) (0.079) (0.080) (0.041) (0.015) (0.003) (0.006)
$\alpha_{32}$0.572*** -0.577*** -0.572*** -0.577*** 0.032 0.019*** 0.035**
(0.166) (0.155) (0.164) (0.079) (0.027) (0.007) (0.014)
$\alpha_{33}$ 0.371*** 0.381*** 0.371*** 0.381*** -0.026 -0.089* -0.021***
(0.106) (0.099) (0.105) (0.050) (0.035) (0.046) (0.004)
$\alpha_{34}$0.076*** -$0**$076*** -0.080*** 0.029* 0.011* 0.038***
(0.022) (0.020) (0.021) (0.010) (0.015) (0.006) (0.007)
$\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 33.876 20.486 392.020 928.892 601.967 394.776 352.386
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.013 0.013 0.007 0.020 0.013 0.013
$H_{0}:\mu=0$ 56.124 66.981
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 12: Estimation and testing results in Section 6.2 using trigonometric series with $h=4$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.230*** 0.275*** 0.230*** 0.275*** 0.211*** 0.039 0.043***
(0.050) (0.043) (0.026) (0.031) (0.065) (0.284) (0.005)
$\tau_{2}/\mu_{2}$ 0.258*** -0.397** -0.437***
(0.039) (0.186) (0.048)
$\tau_{3}/\mu_{3}$ 0.255*** 0.229** 0.253***
(0.055) (0.107) (0.028)
$\tau_{4}/\mu_{4}$ 0.092*** 0.333* 0.367***
(0.031) (0.190) (0.041)
$\beta_{0}$ 3.207*** 2.762*** 3.207*** 2.762*** -0.005 0.001** 0.001***
(0.454) (0.432) (0.871) (0.290) (0.004) (0.001) (0.000)
$\beta_{1}$0.392*** -0.376*** -0.392*** -0.376*** -0.296*** -0.257** -0.200***
(0.026) (0.024) (0.031) (0.024) (0.061) (0.111) (0.025)
$\beta_{2}$ 0.146*** 0.124*** 0.146*** 0.124*** -0.046 0.103 0.069***
(0.025) (0.024) (0.019) (0.016) (0.036) (0.299) (0.010)
$\alpha_{11}$0.089*** -0.082*** -0.089*** -0.082*** 0.234*** -0.074** -0.060***
(0.027) (0.025) (0.019) (0.021) (0.078) (0.032) (0.017)
$\alpha_{12}$ 0.201*** 0.175*** 0.201*** 0.175*** 0.355*** 0.162** 0.068*
(0.068) (0.064) (0.046) (0.045) (0.093) (0.071) (0.038)
$\alpha_{13}$$0**$084*** -$0**$084*** -0.315*** -0.075** -0.020
(0.031) (0.029) (0.022) (0.019) (0.095) (0.033) (0.017)
$\alpha_{14}$$0**$081*** -$0**$081*** 0.236*** -0.073*** -0.054***
(0.026) (0.025) (0.019) (0.021) (0.066) (0.022) (0.017)
$\alpha_{21}$ 0.086*** 0.104*** 0.086*** 0.104*** -0.283*** 0.064*** 0.107***
(0.031) (0.029) (0.021) (0.016) (0.085) (0.019) (0.011)
$\alpha_{22}$0.611*** -0.688*** -0.611*** -0.688*** -0.005 -0.394 -0.506***
(0.179) (0.165) (0.170) (0.146) (0.062) (1.827) (0.138)
$\alpha_{23}$ 0.354*** 0.396*** 0.354*** 0.396*** 0.305*** 0.229*** 0.278***
(0.103) (0.095) (0.100) (0.086) (0.108) (0.071) (0.080)
$\alpha_{24}$ 0.025 0.050 0.025 $0**$042 0.036 0.091***
(0.036) (0.034) (0.019) (0.014) (0.033) (0.076) (0.011)
$\alpha_{31}$0.027** -0.027** -0.027*** -0.027*** 0.128*** -0.025*** -0.035***
(0.013) (0.012) (0.009) (0.007) (0.032) (0.006) (0.005)
$\alpha_{32}$ 0.216*** 0.238*** 0.216*** 0.238*** -0.138 0.150* 0.217***
(0.061) (0.056) (0.057) (0.048) (0.086) (0.088) (0.043)
$\alpha_{33}$0.123*** -0.136*** -0.123*** -0.136*** -0.047 -0.086** -0.121***
(0.035) (0.033) (0.033) (0.029) (0.046) (0.038) (0.025)
$\alpha_{34}$0.004 -0.012 -0.004 -0.012** -0.014 -0.009 -0.030***
(0.013) (0.012) (0.008) (0.005) (0.026) (0.018) (0.005)
Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 258.704 91.459 123.439
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
$H_{0}:\mu=0$ 43.155 46.397
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 13: Comparison of parametric method and semiparametric method in Section 6.2.
polynomial,$h=2$ polynomial, $h=4$ trigonometric,$h=2$ trigonometric,$h=4$
$W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$
0.3136 0.3371 0.3252 0.3406 0.3074 0.3221 0.2996 0.3125
Appendix D Simulation results
Table 14: Empirical sizes for $\mathbb{W}_{1}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.072 0.102 0.038 0.072 0.094 0.048 0.076 0.092
$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.088 0.029 0.069 0.118 0.019 0.078 0.127
$h=4$ $\mathscr{T}_{n}$ 0.044 0.090 0.118 0.032 0.080 0.110 0.036 0.082 0.110
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.088 0.078 0.127 0.137 0.019 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.058 0.132 0.184 0.032 0.106 0.146 0.038 0.064 0.094
$\mathscr{T}_{n}^{*}$ 0.049 0.088 0.137 0.028 0.127 0.137 0.019 0.098 0.108
7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.100 0.132 0.026 0.076 0.110 0.048 0.082 0.110
$\mathscr{T}_{n}^{*}$ 0.029 0.137 0.167 0.020 0.088 0.098 0.020 0.059 0.118
$h=4$ $\mathscr{T}_{n}$ 0.048 0.100 0.142 0.024 0.066 0.100 0.026 0.066 0.100
$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.186 0.019 0.078 0.118 0.019 0.078 0.137
$h=8$ $\mathscr{T}_{n}$ 0.060 0.120 0.178 0.042 0.074 0.120 0.018 0.054 0.104
$\mathscr{T}_{n}^{*}$ 0.038 0.176 0.196 0.039 0.078 0.147 0.020 0.078 0.137
7.1 $h=2$ $\mathscr{T}_{n}$ 0.062 0.108 0.140 0.028 0.058 0.092 0.026 0.056 0.098
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.008 0.020 0.039 0.019 0.088 0.118
$h=4$ $\mathscr{T}_{n}$ 0.072 0.116 0.170 0.020 0.066 0.116 0.036 0.078 0.108
$\mathscr{T}_{n}^{*}$ 0.049 0.078 0.127 0.029 0.069 0.088 0.015 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.094 0.168 0.216 0.026 0.062 0.112 0.034 0.078 0.136
$\mathscr{T}_{n}^{*}$ 0.058 0.127 0.186 0.029 0.108 0.167 0.014 0.088 0.137
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.064 0.098 0.044 0.070 0.104 0.046 0.070 0.092
$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.078 0.029 0.039 0.088 0.019 0.048 0.118
$h=4$ $\mathscr{T}_{n}$ 0.052 0.092 0.122 0.030 0.076 0.140 0.032 0.074 0.116
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.069 0.039 0.108 0.118 0.019 0.058 0.137
$h=8$ $\mathscr{T}_{n}$ 0.068 0.138 0.180 0.038 0.110 0.134 0.034 0.066 0.094
$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.137 0.029 0.118 0.157 0.019 0.038 0.108
7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.084 0.112 0.038 0.084 0.118 0.048 0.084 0.114
$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.176 0.029 0.059 0.108 0.010 0.039 0.088
$h=4$ $\mathscr{T}_{n}$ 0.050 0.108 0.152 0.022 0.066 0.104 0.032 0.070 0.114
$\mathscr{T}_{n}^{*}$ 0.069 0.118 0.186 0.049 0.098 0.127 0.019 0.058 0.127
$h=8$ $\mathscr{T}_{n}$ 0.064 0.108 0.162 0.032 0.086 0.124 0.022 0.052 0.080
$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.020 0.108 0.157 0.019 0.058 0.157
7.1 $h=2$ $\mathscr{T}_{n}$ 0.060 0.090 0.120 0.034 0.070 0.112 0.022 0.062 0.098
$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.098 0.020 0.039 0.078 0.018 0.068 0.137
$h=4$ $\mathscr{T}_{n}$ 0.064 0.094 0.164 0.022 0.076 0.114 0.032 0.070 0.106
$\mathscr{T}_{n}^{*}$ 0.078 0.098 0.147 0.039 0.088 0.098 0.017 0.068 0.118
$h=8$ $\mathscr{T}_{n}$ 0.098 0.158 0.206 0.024 0.060 0.102 0.032 0.078 0.120
$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.039 0.127 0.167 0.016 0.068 0.127
Table 15: Empirical sizes for $\mathbb{W}_{2}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.088 0.200 0.248 0.094 0.156 0.212 0.090 0.190 0.266
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.108 0.049 0.088 0.127 0.020 0.039 0.069
$h=4$ $\mathscr{T}_{n}$ 0.060 0.150 0.206 0.084 0.170 0.218 0.086 0.172 0.204
$\mathscr{T}_{n}^{*}$ 0.010 0.039 0.039 0.029 0.049 0.069 0.049 0.127 0.157
$h=8$ $\mathscr{T}_{n}$ 0.062 0.134 0.194 0.062 0.136 0.174 0.068 0.144 0.202
$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.108 0.049 0.078 0.098 0.020 0.127 0.137
7.1 $h=2$ $\mathscr{T}_{n}$ 0.102 0.142 0.182 0.054 0.096 0.152 0.052 0.106 0.136
$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.118 0.059 0.108 0.127 0.020 0.078 0.176
$h=4$ $\mathscr{T}_{n}$ 0.098 0.158 0.212 0.042 0.090 0.136 0.036 0.086 0.114
$\mathscr{T}_{n}^{*}$ 0.020 0.069 0.118 0.069 0.118 0.137 0.029 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.120 0.210 0.282 0.066 0.122 0.162 0.034 0.088 0.144
$\mathscr{T}_{n}^{*}$ 0.010 0.059 0.088 0.049 0.127 0.157 0.029 0.069 0.167
7.1 $h=2$ $\mathscr{T}_{n}$ 0.068 0.118 0.142 0.024 0.054 0.084 0.028 0.078 0.092
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.000 0.020 0.088 0.069 0.069 0.108
$h=4$ $\mathscr{T}_{n}$ 0.054 0.106 0.164 0.024 0.062 0.098 0.048 0.076 0.108
$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.069 0.020 0.059 0.088 0.039 0.098 0.127
$h=8$ $\mathscr{T}_{n}$ 0.062 0.138 0.174 0.028 0.058 0.092 0.032 0.084 0.132
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.029 0.069 0.098 0.039 0.078 0.127
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.082 0.160 0.212 0.106 0.198 0.246 0.084 0.176 0.232
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.078 0.108 0.127 0.020 0.088 0.118
$h=4$ $\mathscr{T}_{n}$ 0.072 0.146 0.208 0.076 0.160 0.204 0.086 0.150 0.208
$\mathscr{T}_{n}^{*}$ 0.029 0.118 0.118 0.029 0.039 0.078 0.029 0.069 0.098
$h=8$ $\mathscr{T}_{n}$ 0.086 0.144 0.194 0.076 0.154 0.204 0.084 0.152 0.184
$\mathscr{T}_{n}^{*}$ 0.039 0.108 0.137 0.029 0.078 0.098 0.049 0.108 0.118
7.1 $h=2$ $\mathscr{T}_{n}$ 0.086 0.134 0.180 0.056 0.112 0.140 0.056 0.108 0.144
$\mathscr{T}_{n}^{*}$ 0.039 0.078 0.118 0.039 0.108 0.137 0.029 0.049 0.147
$h=4$ $\mathscr{T}_{n}$ 0.102 0.188 0.240 0.040 0.100 0.140 0.052 0.092 0.138
$\mathscr{T}_{n}^{*}$ 0.020 0.059 0.069 0.078 0.137 0.167 0.049 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.112 0.192 0.270 0.052 0.106 0.160 0.032 0.076 0.132
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.049 0.029 0.167 0.176 0.049 0.118 0.167
7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.096 0.120 0.026 0.064 0.098 0.026 0.072 0.102
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.069 0.020 0.059 0.078 0.059 0.098 0.127
$h=4$ $\mathscr{T}_{n}$ 0.040 0.104 0.150 0.022 0.066 0.104 0.034 0.068 0.118
$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.118 0.049 0.088 0.088 0.059 0.088 0.127
$h=8$ $\mathscr{T}_{n}$ 0.056 0.132 0.158 0.022 0.064 0.096 0.024 0.076 0.130
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.078 0.029 0.078 0.127 0.049 0.088 0.108
Table 16: Empirical sizes for $\mathbb{W}_{3}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
$r=2$ 7.1 $h=2$ 0.032 0.096 0.164 0.040 0.096 0.116 0.014 0.056 0.096
$h=4$ 0.028 0.056 0.084 0.024 0.052 0.084 0.036 0.060 0.080
$h=8$ 0.016 0.036 0.052 0.024 0.036 0.048 0.024 0.040 0.092
7.1 $h=2$ 0.040 0.088 0.148 0.044 0.096 0.120 0.020 0.060 0.112
$h=4$ 0.040 0.088 0.104 0.032 0.100 0.120 0.012 0.044 0.066
$h=8$ 0.016 0.048 0.064 0.008 0.016 0.028 0.028 0.056 0.092
7.1 $h=2$ 0.040 0.104 0.116 0.024 0.076 0.104 0.048 0.092 0.124
$h=4$ 0.028 0.052 0.088 0.032 0.048 0.068 0.024 0.072 0.104
$h=8$ 0.020 0.048 0.052 0.024 0.056 0.072 0.004 0.024 0.052
$r=4$ 7.1 $h=2$ 0.032 0.120 0.156 0.040 0.084 0.108 0.016 0.052 0.106
$h=4$ 0.036 0.064 0.092 0.032 0.076 0.088 0.020 0.040 0.076
$h=8$ 0.012 0.032 0.040 0.012 0.032 0.048 0.004 0.032 0.044
7.1 $h=2$ 0.032 0.100 0.136 0.032 0.076 0.092 0.048 0.100 0.144
$h=4$ 0.044 0.096 0.100 0.052 0.092 0.116 0.004 0.028 0.056
$h=8$ 0.008 0.044 0.056 0.008 0.036 0.052 0.024 0.052 0.080
7.1 $h=2$ 0.056 0.092 0.128 0.048 0.080 0.132 0.060 0.100 0.132
$h=4$ 0.028 0.060 0.084 0.008 0.044 0.056 0.020 0.068 0.120
$h=8$ 0.032 0.044 0.060 0.024 0.060 0.084 0.020 0.048 0.064
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
$r=2$ 7.1 $h=2$ 0.028 0.092 0.160 0.030 0.096 0.116 0.012 0.060 0.106
$h=4$ 0.020 0.060 0.084 0.024 0.052 0.084 0.028 0.056 0.072
$h=8$ 0.008 0.032 0.048 0.020 0.032 0.048 0.024 0.040 0.072
7.1 $h=2$ 0.044 0.096 0.152 0.044 0.096 0.120 0.014 0.058 0.116
$h=4$ 0.040 0.080 0.108 0.032 0.104 0.116 0.012 0.040 0.088
$h=8$ 0.024 0.044 0.064 0.008 0.016 0.028 0.032 0.060 0.088
7.1 $h=2$ 0.036 0.100 0.112 0.024 0.076 0.104 0.048 0.092 0.132
$h=4$ 0.016 0.052 0.092 0.028 0.048 0.068 0.028 0.068 0.104
$h=8$ 0.012 0.036 0.048 0.032 0.056 0.068 0.004 0.028 0.052
$r=4$ 7.1 $h=2$ 0.032 0.116 0.156 0.040 0.084 0.108 0.018 0.050 0.102
$h=4$ 0.032 0.060 0.088 0.032 0.076 0.088 0.016 0.028 0.080
$h=8$ 0.012 0.032 0.048 0.012 0.032 0.048 0.004 0.032 0.098
7.1 $h=2$ 0.032 0.104 0.136 0.032 0.076 0.092 0.014 0.062 0.104
$h=4$ 0.040 0.100 0.108 0.052 0.092 0.116 0.008 0.028 0.076
$h=8$ 0.024 0.056 0.064 0.008 0.040 0.056 0.024 0.056 0.110
7.1 $h=2$ 0.060 0.096 0.136 0.028 0.080 0.132 0.014 0.048 0.102
$h=4$ 0.036 0.064 0.088 0.008 0.044 0.056 0.024 0.068 0.112
$h=8$ 0.020 0.044 0.060 0.020 0.056 0.080 0.016 0.048 0.094
Table 17: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\delta(z)$.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.088 0.128 0.064 0.104 0.156 0.044 0.112 0.152
$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.064 0.004 0.022 0.074 0.012 0.036 0.082
$h=4$ $\mathscr{T}_{n}$ 0.044 0.080 0.116 0.016 0.048 0.080 0.040 0.076 0.096
$\mathscr{T}_{n}^{*}$ 0.000 0.024 0.076 0.000 0.040 0.072 0.012 0.036 0.084
$h=8$ $\mathscr{T}_{n}$ 0.008 0.024 0.040 0.004 0.008 0.012 0.004 0.016 0.032
$\mathscr{T}_{n}^{*}$ 0.020 0.064 0.072 0.008 0.020 0.076 0.024 0.052 0.090
7.1 $h=2$ $\mathscr{T}_{n}$ 0.048 0.124 0.160 0.032 0.060 0.108 0.060 0.136 0.172
$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.024 0.004 0.012 0.016 0.016 0.028 0.060
$h=4$ $\mathscr{T}_{n}$ 0.060 0.112 0.132 0.028 0.068 0.104 0.044 0.096 0.128
$\mathscr{T}_{n}^{*}$ 0.000 0.012 0.036 0.000 0.000 0.020 0.008 0.012 0.024
$h=8$ $\mathscr{T}_{n}$ 0.016 0.044 0.056 0.008 0.032 0.056 0.012 0.040 0.068
$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.076 0.004 0.038 0.072 0.006 0.044 0.072
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.104 0.140 0.052 0.116 0.168 0.028 0.104 0.156
$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.068 0.000 0.060 0.076 0.000 0.068 0.094
$h=4$ $\mathscr{T}_{n}$ 0.004 0.016 0.024 0.000 0.004 0.012 0.044 0.068 0.108
$\mathscr{T}_{n}^{*}$ 0.024 0.072 0.100 0.016 0.044 0.080 0.012 0.040 0.088
$h=8$ $\mathscr{T}_{n}$ 0.000 0.028 0.048 0.012 0.032 0.068 0.016 0.036 0.068
$\mathscr{T}_{n}^{*}$ 0.006 0.032 0.084 0.000 0.034 0.078 0.004 0.040 0.076
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.104 0.144 0.044 0.092 0.116 0.048 0.116 0.160
$\mathscr{T}_{n}^{*}$ 0.004 0.022 0.068 0.016 0.020 0.076 0.004 0.034 0.088
$h=4$ $\mathscr{T}_{n}$ 0.012 0.044 0.076 0.032 0.076 0.116 0.052 0.116 0.132
$\mathscr{T}_{n}^{*}$ 0.004 0.024 0.074 0.004 0.030 0.080 0.004 0.038 0.078
$h=8$ $\mathscr{T}_{n}$ 0.000 0.004 0.004 0.000 0.004 0.008 0.008 0.020 0.032
$\mathscr{T}_{n}^{*}$ 0.016 0.036 0.078 0.008 0.020 0.080 0.016 0.028 0.082
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.108 0.132 0.032 0.084 0.108 0.040 0.112 0.148
$\mathscr{T}_{n}^{*}$ 0.000 0.008 0.024 0.000 0.000 0.012 0.008 0.024 0.056
$h=4$ $\mathscr{T}_{n}$ 0.048 0.076 0.116 0.044 0.072 0.104 0.036 0.088 0.128
$\mathscr{T}_{n}^{*}$ 0.008 0.012 0.028 0.000 0.004 0.012 0.000 0.016 0.044
$h=8$ $\mathscr{T}_{n}$ 0.012 0.060 0.072 0.012 0.024 0.040 0.016 0.040 0.116
$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.072 0.004 0.038 0.068 0.008 0.040 0.092
7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.128 0.168 0.044 0.084 0.128 0.040 0.108 0.152
$\mathscr{T}_{n}^{*}$ 0.004 0.030 0.064 0.000 0.044 0.088 0.000 0.046 0.094
$h=4$ $\mathscr{T}_{n}$ 0.016 0.024 0.040 0.008 0.056 0.080 0.044 0.104 0.136
$\mathscr{T}_{n}^{*}$ 0.044 0.088 0.100 0.012 0.030 0.068 0.008 0.040 0.094
$h=8$ $\mathscr{T}_{n}$ 0.032 0.044 0.060 0.008 0.036 0.052 0.020 0.040 0.068
$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.082 0.004 0.028 0.062 0.006 0.032 0.068
Table 18: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\lambda(z)$.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.136 0.168 0.048 0.096 0.140 0.060 0.124 0.156
$\mathscr{T}_{n}^{*}$ 0.024 0.048 0.088 0.020 0.048 0.084 0.004 0.036 0.082
$h=4$ $\mathscr{T}_{n}$ 0.040 0.068 0.096 0.040 0.080 0.100 0.032 0.088 0.132
$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.094 0.012 0.032 0.076 0.004 0.052 0.092
$h=8$ $\mathscr{T}_{n}$ 0.020 0.056 0.100 0.024 0.068 0.084 0.040 0.096 0.116
$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.088 0.012 0.036 0.068 0.020 0.036 0.100
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.076 0.108 0.036 0.104 0.136 0.072 0.124 0.168
$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.052 0.012 0.032 0.076 0.020 0.052 0.104
$h=4$ $\mathscr{T}_{n}$ 0.032 0.108 0.152 0.028 0.076 0.120 0.060 0.108 0.136
$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.068 0.008 0.020 0.068 0.020 0.040 0.092
$h=8$ $\mathscr{T}_{n}$ 0.044 0.088 0.100 0.020 0.040 0.106 0.012 0.040 0.096
$\mathscr{T}_{n}^{*}$ 0.020 0.060 0.080 0.012 0.028 0.036 0.012 0.028 0.040
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.136 0.192 0.048 0.116 0.148 0.072 0.184 0.216
$\mathscr{T}_{n}^{*}$ 0.008 0.048 0.080 0.012 0.040 0.068 0.020 0.064 0.108
$h=4$ $\mathscr{T}_{n}$ 0.028 0.064 0.108 0.032 0.076 0.092 0.072 0.144 0.204
$\mathscr{T}_{n}^{*}$ 0.008 0.032 0.072 0.008 0.028 0.056 0.024 0.056 0.128
$h=8$ $\mathscr{T}_{n}$ 0.036 0.072 0.076 0.020 0.076 0.112 0.028 0.064 0.080
$\mathscr{T}_{n}^{*}$ 0.012 0.044 0.088 0.004 0.036 0.076 0.016 0.040 0.086
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.112 0.152 0.048 0.100 0.124 0.044 0.108 0.132
$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.088 0.016 0.040 0.094 0.008 0.058 0.082
$h=4$ $\mathscr{T}_{n}$ 0.048 0.096 0.108 0.012 0.072 0.108 0.056 0.116 0.136
$\mathscr{T}_{n}^{*}$ 0.016 0.044 0.088 0.004 0.038 0.082 0.012 0.048 0.088
$h=8$ $\mathscr{T}_{n}$ 0.032 0.056 0.080 0.028 0.076 0.096 0.028 0.064 0.072
$\mathscr{T}_{n}^{*}$ 0.024 0.036 0.072 0.008 0.036 0.080 0.024 0.036 0.088
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.092 0.136 0.064 0.112 0.172 0.056 0.152 0.200
$\mathscr{T}_{n}^{*}$ 0.008 0.044 0.084 0.004 0.028 0.080 0.016 0.056 0.116
$h=4$ $\mathscr{T}_{n}$ 0.036 0.076 0.104 0.020 0.060 0.084 0.052 0.096 0.112
$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.052 0.012 0.024 0.056 0.008 0.044 0.080
$h=8$ $\mathscr{T}_{n}$ 0.028 0.064 0.092 0.028 0.056 0.092 0.016 0.044 0.094
$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.064 0.012 0.040 0.064 0.008 0.020 0.048
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.128 0.164 0.064 0.116 0.164 0.040 0.112 0.132
$\mathscr{T}_{n}^{*}$ 0.004 0.028 0.076 0.016 0.040 0.072 0.024 0.048 0.092
$h=4$ $\mathscr{T}_{n}$ 0.020 0.076 0.092 0.036 0.088 0.116 0.028 0.088 0.116
$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.048 0.012 0.040 0.084 0.000 0.036 0.088
$h=8$ $\mathscr{T}_{n}$ 0.024 0.052 0.076 0.012 0.044 0.072 0.024 0.048 0.060
$\mathscr{T}_{n}^{*}$ 0.012 0.048 0.090 0.008 0.050 0.086 0.000 0.020 0.072

### v2

@@
 Abstract
-We present simple to implement Wald-type statistics that deliver a general nonparametric inference theory for linear restrictions on varying coefficients in a range of spatial autoregressive models.
-Our theory covers error dependence of a general form, allows for a degree of misspecification robustness via nonparametric spatial weights and permits inference on both varying regression and spatial coefficients.
-One application of our method finds evidence for constant returns to scale in the production function of the Chinese nonmetal mineral industry, while another finds a nonlinear impact of the distance to the employment center on housing prices in Boston.
-A simulation study confirms that our tests perform well in finite-samples.
+We present simple to implement Wald-type statistics that deliver a general nonparametric inference theory for linear restrictions on varying coefficients in a range of regression models allowing for cross-sectional or spatial dependence.
+We provide a general central limit theorem that covers a broad range of error spatial dependence structures, allows for a degree of misspecification robustness via nonparametric spatial weights and permits inference on both varying regression and spatial dependence parameters.
+Using our method, we first uncover evidence of constant returns to scale in the Chinese nonmetal mineral industry’s production function, and then show that Boston house prices respond nonlinearly to proximity to employment centers.
+A simulation study confirms that our tests perform very well in finite samples.
 JEL codes: C21, C31
-Keywords: Spatial autoregression, varying coefficients, inference
+Keywords: Spatial autoregression, varying coefficients, nonparametric inference
 1.  Introduction
-This paper develops a general framework for inference on varying coefficient spatial autoregressive (SAR) models that is as easy to implement in practice as familiar Wald tests.
-SAR models are fast becoming a workhorse for researchers working in settings where economic agents interact via economic or social links.
+This paper develops a general framework for inference on varying coefficient regression models that is as easy to implement in practice as familiar Wald tests for finite-dimensional parameters.
+We also allow for spatial autoregressive (SAR) structure to permit interaction across the cross-section of economic agents.
+SAR models are a popular and parsimonious tool for researchers working in settings where economic agents interact via economic or social links.
 Such links can be geographic proximity, or proximity in some more general sense such as a social network.
-The SAR model can accommodate these types of general links easily, and this goes some way to explain its appeal.
+The SAR model can accommodate these types of general links easily, and this goes some way to explain its appeal, see e.g.
+Helmers and Patnam (2014) and Hsieh and van Kippersluis (2018) for examples of applications to peer effects.
 No surprise then that the baseline SAR model, due to Cliff and Ord (1981), has become much studied by econometricians.
-This model has long been popular with regional scientists, see Anselin (1988) for an early book length treatment; important econometric contributions that established the foundation for rigorous econometric analysis of such models include Kelejian and Prucha (1998) and Lee (2004).
-In keeping with the growing importance of SAR models, various strands of the literature have explored how to widen their scope.
+This model has long been popular with regional scientists, see Anselin (1988) for an early book length treatment; important econometric contributions that established the foundation for rigorous econometric analysis of such models include Kelejian and Prucha (1998), Lee (2004) and Kuersteiner and Prucha (2020).
+In keeping with the growing importance of SAR models and their ability to parsimoniously control for cross-sectional dependence in multiple regression designs, various strands of the literature have explored how to widen their scope.
 One such is the focus on varying coefficient SAR models, wherein model parameters are allowed to be unknown functions of some observed economic variable(s).
-This permits a flexible effect of the regressors and spatial dependence on the outcome variable, and indeed varying coefficient models are abundant in statistics and econometrics (see e.g.
-Fan and Zhang, 1999, 2008).
+This permits a flexible effect of both the regressors and spatial dependence on the outcome variable, and indeed varying coefficient models are abundant in statistics and econometrics (see e.g.
+Fan and Zhang (1999, 2008)).
 In the SAR context, varying coefficient models have been studied by Sun et al.
@@
 In this paper, we show how inference on varying coefficients in a range of SAR models can be conducted just like familiar parametric Wald tests for linear restrictions.
+Standard varying coefficient regression without cross-sectional dependence is naturally covered as a special case.
 The key idea is that if series approximations are used for the nonparametric varying coefficients then the series coefficients can be employed for very easy practical inference.
@@
 (2012), Gupta (2018) and Korolev (2019), amongst others.
-[…]
+Naturally our methods are also applicable to the usual varying coefficient regression model with no cross-sectional dependence, see e.g.
+Ahmad et al.
+(2005), which is a simply a particular case of our theory.
+To explain the intuition of our approach, suppose a sample of $n$ observations is available.
+With our method, the researcher can reduce the econometric problem of inference on an infinite-dimensional object to inference on a growing number of series coefficients, say $p$.
+In practice, this simply means writing down the usual Wald test statistic $\mathscr{W}$ for $p$ linear restrictions, observing that $p\rightarrow\infty$ as $n\rightarrow\infty$ if it is the length of a series approximation, and appealing to a theorem that establishes $(\mathscr{W}-p)/\sqrt{2p}\overset{d}{\rightarrow}N(0,1)$ as $n\rightarrow\infty$, under the null hypothesis.
+The idea stems from the fact that as $p\rightarrow\infty$, a $(\chi^{2}_{p}-p)/\sqrt{2p}$ random variable approaches a standard normal variate, see e.g.
+de Jong and Bierens (1994), Hong and White (1995), Gupta (2018) and Gupta and Seo (2023).
 Using this series expansion approach, we theoretically justify Wald-type test statistics for inference on varying coefficients.
-We cover a baseline high-order SAR model with varying regression coefficients but constant spatial lag parameters and extend the model to allow for spatial error dependence.
+We first cover a baseline high-order SAR model with varying regression coefficients but constant spatial lag parameters and extend the model to allow for spatial error dependence.
 The baseline model is similar to the one estimated by Sun et al.
@@
 Our approach is also shown to work for models with varying coefficients on the spatial lags, in the spirit of Malikov and Sun (2017).
-We then apply our methods on two different studies where varying coefficients have been shown to play an important role.
+We then apply our methods in two different settings where varying coefficients have been shown to play an important role.
 In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry.
-This can be seen as a follow-up application to that of Li et al.
-(2002), who sugges that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e.
+This can be seen as a follow-up study to Li et al.
+(2002), who suggest that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e.
 output elasticities of capital and labor sum to unity) for most observed managerial expense levels.
-However, they do not have a test for CRS and do not model spatial dependence.
-Our analysis confirms their observations as we cannot reject the CRS hypothesis using their model and in much richer settings.
-This suggests that CRS technology may be a salient feature of this industry.
-Our second study is an illustration on the famous Boston housing data.
+Our analysis confirms their results as we cannot reject the CRS hypothesis using a more general version of their model that permits spatial dependence.
+This suggests that CRS technology may well be a salient feature of this industry.
+Our second study re-examines the Boston house price data of Harrison Jr and Rubinfeld (1978).
 We estimate a model similar to Sun et al.
@@
 We formally test for the varying location effects of these features using our statistics and indeed find them to be statistically significant.
-In a set of Monte Carlo simulation studies, we show that our tests have excellent finite-sample performance.
-In addition to tests that are based on the asymptotic normal approximation, we also propose a bootstrap version of our tests; the bootstrap tests perform particularly well for small sample sizes.
-We experiment with a range of SAR models and spatial dependence structures to demonstrate the full range of applicability and generality of our approach.
-The next section introduces our method in the context of a baseline higher-order SAR model.
-Section 3 extends the model to allow for spatial error dependence of Kelejian and Prucha (2007) type, while Section 4 allows for nonparametric spatial weights (Pinkse et al., 2002).
-In Section 5 we allow the coefficient on the spatial lag to also vary (Malikov and Sun, 2017) and show how our approach can be adapted to this setting.
-Section 6 presents our empirical applications to the Chinese nonmetal mineral industry and Boston house prices, while Section 7 contains the simulation study.
-Proofs are in the appendix, and a supplementary online appendix contains a number of additional simulation results.
-2.   Baseline spatial autoregression
+In Monte Carlo simulation studies, we experiment with a range of SAR models and spatial dependence structures to demonstrate the applicability and generality of our approach.
+Results show that our tests have excellent finite-sample performances.
+The rest of the paper is as follows: Section 2 introduces our method in a baseline regression model allowing for higher-order SAR structure.
+Section 3 extends the model to allow for spatial error dependence of the Kelejian and Prucha (2007) type.
+Section 4 allows for nonparametric spatial weights, as in Pinkse et al.
+(2002), and thereby permits some degree of weight matrix robustness.
+Section 5 shows how our approach can be adapted to the setting where the spatial lag coefficient varies, like in Malikov and Sun (2017).
+Section 6 applies our methods to the Chinese nonmetal mineral industry and Boston house prices.
+Section 7 presents the simulation study, and Section 8 concludes.
+Proofs are in the appendix, while additional empirical and simulation results are in the online appendix.
+2.  Baseline higher-order spatial autoregression
 Consider a vector of unknown functions $\delta\left(\cdot\right)=\left(\delta_{1}\left(\cdot\right),\ldots,\delta_{d_{\delta}}\left(\cdot\right)\right)^{\prime}$, and the model
 $y_{in}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{in,j}^{\prime}y_{n}+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (2.1)
-where $y_{n}$ is the $n\times 1$ vector with typical element $y_{in}$, $w_{in,j}^{\prime}$ is the $i$th row of the spatial weight matrix $W_{jn}$, $j=1,\ldots,d_{\lambda}$, $d_{\lambda}\rightarrow\infty$ as $n\rightarrow\infty$, and $\epsilon_{in}$ is $i.i.d.$ with mean 0 and unit variance.
-Also, $\left(x_{in},p_{in},z_{in}\right)\in\mathcal{X}\times\mathcal{P}\times\mathcal{Z}\subseteq\mathbb{R}^{d_{\beta}\times d_{\delta}\times d_{z}}$, and $d_{\beta}\rightarrow 0$ as $n\rightarrow\infty$ but $d_{z}$ and $d_{\delta}$ are fixed.
+where $y_{n}$ is the $n\times 1$ vector with typical element $y_{in}$, $w_{in,j}^{\prime}$ is the $i$-th row of the spatial weight matrix $W_{jn}$, $j=1,\ldots,d_{\lambda}$, $d_{\lambda}\rightarrow\infty$ as $n\rightarrow\infty$, and $\epsilon_{in}$ is $i.i.d.$ with mean 0 and unit variance.
+Also, $\left(x_{in},p_{in},z_{in}\right)\in\mathcal{X}\times\mathcal{P}\times\mathcal{Z}\subseteq\mathbb{R}^{d_{\beta}\times d_{\delta}\times d_{z}}$, and $d_{\beta}\rightarrow\infty$ as $n\rightarrow\infty$ but $d_{z}$ and $d_{\delta}$ are fixed.
 For all $i=1,\ldots,n$, $z_{in}$ is throughout the paper uncorrelated with $\epsilon_{in}$ while $x_{in}$ is allowed to be correlated with $\epsilon_{in}$.
-Permitting $d_{\lambda}$ and $d_{\beta}$ to diverge with sample size allows a flexible modelling approach, with the model becoming richer as more observations become available to the practitioner.
-This approach has been studied in statistics since at least the work of Huber (1973), but in spirit the sequence of experiments considered by Le Cam (1960) provides an even earlier reference.
+It is worth noting that our results also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed but we state our theorems for the more complex case where they diverge.
+Allowing $d_{\lambda}$ and $d_{\beta}$ to diverge with sample size allows a flexible modeling approach, which has been studied in statistics since at least the work of Huber (1973), but in spirit the sequence of experiments considered by Le Cam (1960) provides an even earlier reference.
 Econometricians have also studied such models in least squares and GMM settings, see e.g.
 Andrews (1985), Koenker and Machado (1999) and Cattaneo et al.
-(2018), to name just three references.
-In a SAR setting, Gupta and Robinson (2015), Gupta (2018) and Gupta and Qu (2024) emphasize the advantages of this approach, especially when clustered data naturally imply asymptotic regimes in which $d_{\lambda}$ diverges with $n$.
-It is worth noting that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed but we state our theorems for the more complex case where they diverge.
-We now mostly drop $n$ subscripting, but occasionally remind the reader of the dependence on sample size of various quantities.
-The paper takes the spatial weight matrices $W_{j}$ to be non-stochastic everywhere except in Section 4.
-In all sections these weight matrices are also uniformly bounded in row and column sums, a commonly used restriction to control spatial dependence (Kelejian and Prucha, 1998; Lee, 2004).
+(2018), to name a few.
+In a SAR setting, Gupta and Robinson (2015) and Gupta (2018) emphasize the advantages of this approach, especially when clustered data imply asymptotic regimes where $d_{\lambda}$ diverges with $n$.
+We now drop $n$ subscripting, but occasionally remind the reader of the dependence on sample size of various quantities.
+We take spatial weight matrices $W_{j}$ to be non-stochastic everywhere except in Section 4.
+These weight matrices are also uniformly bounded in row and column sums, a commonly used restriction to control spatial dependence (Kelejian and Prucha, 1998; Lee, 2004).
 This property is assumed to hold almost surely in Section 4, where the spatial weights are stochastic.
@@
 Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$.
-Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$, with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$.
+Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$ with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$.
 Then we can write (2.1) as
 $y_{i}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{i,j}^{\prime}y+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (2.3)
-with $u_{i}=r_{i}+\epsilon_{i}$, where $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error.
+where $u_{i}=r_{i}+\epsilon_{i}$ and $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error.
 Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$.
-Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can be written in matrix notation as
+Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can then be written as
 $y=Q\lambda+X\beta+\Psi\alpha+u=L\xi+u,$ (2.4)
-where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$.
-The null hypothesis to approximate $H_{0}^{true}$ is
+where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. The null to approximate $H_{0}^{true}$ is
 $H_{0}:\alpha=0,$ (2.5)
@@
 $\hat{\xi}\equiv\left(\hat{\lambda}^{\prime},\hat{\beta}^{\prime},\hat{\alpha}^{\prime}\right)^{\prime}=\left(L^{\prime}{P_{K_{1}}}L\right)^{-1}L^{\prime}{P_{K_{1}}}y,$ (2.6)
-where $P_{K_{1}}=K_{1}\left(K_{1}^{\prime}K_{1}\right)^{-1}K_{1}^{\prime}$, with $K_{1}$ being an $n\times J_{1}$ instrument matrix with $J_{1}\geq d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$ but having the same asymptotic order i.e.
-${J_{1}}/{d_{\xi}}$ tends to a constant that is at least unity, as $n\rightarrow\infty$.
+where $P_{K_{1}}=K_{1}\left(K_{1}^{\prime}K_{1}\right)^{-1}K_{1}^{\prime}$ and $K_{1}$ is an $n\times J_{1}$ instrument matrix with $J_{1}\geq d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$ but the same asymptotic order, i.e.
+${J_{1}}/{d_{\xi}}$ tends to a constant at least unity, as $n\rightarrow\infty$.
 Let $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\lambda}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix.
-.
 Then, the test statistic is
@@
 2.2 Asymptotic properties
-We begin with restrictions on the moments of various objects, noting that throughout the paper $C$ denotes a generic positive constant, arbitrarily large but independent of $n$.
-Let $k_{ri}$ be the elements of the instrumental variable matrix, which corresponds to $K_{1},K_{2}$ and $K_{3}$ as defined in this section, 4 and 5, respectively, and $l_{ri}$ denote the elements of $L$.
+We begin with restrictions on moments of various objects, noting that throughout the paper $C$ denotes a generic positive constant, arbitrarily large but independent of $n$.
+Let $k_{ri}$ be elements of the instrumental variable matrix, corresponding to $K_{1},K_{2}$ and $K_{3}$ as defined in the following sections, and $l_{ri}$ denote the elements of $L$.
 Assumption 1.
@@
 $\mathrm{E}\left(l_{ri}^{2}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
-Assumption 1 is a general assumption that controls the approximation error.
+Assumption 1 controls the approximation error.
 Sufficient conditions for various cases can be found in Chen (2007).
-We avoid using specific rates for the approximation error because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions.
+We avoid using specific rates because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions.
 Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition.
 Assumption 3 imposes moment conditions on regressors and instruments.
-To ease notation, we introduce the following definition which captures boundedness and non-multicollinearity properties of random or fixed matrices of growing dimension.
+To ease notation, we introduce the following definition that captures boundedness and non-multicollinearity of random or fixed matrices of growing dimension.
 Definition 1.
@@
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}}{n}\rightarrow 0,$ (2.8)
-as $n\rightarrow\infty$,
+as $n\rightarrow\infty$, we have
 $\mathbb{W}_{1}-\frac{\epsilon^{\prime}\mathcal{M}_{1}\epsilon-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
-We take this opportunity to remind the reader that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
+We remind the reader that results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
 Assumption 5.
 $\left(I_{n}-\sum_{j=1}^{d_{\lambda}}\lambda_{j}W_{j}\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$.
-Define $H_{\ell 1}\equiv H_{\ell 1,n}:\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$, with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proof of the next theorem.
+Define $H_{\ell 1}\equiv H_{\ell 1,n}:\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$, with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in detail in the proofs.
 This sequence of local alternatives features a $d_{\alpha}^{1/4}$ damping factor that accounts for the cost of our nonparametric approach, and has been found in similar problems by Hong and White (1995) and Gupta (2018), amongst others.
 We can now state the main theorem of the section.
-Theorem 2.2.
+Theorem $2.$
 Under Assumptions 1-5 and
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}\rightarrow 0,$ (2.9)
-as $n\rightarrow\infty$, the following hold:
-[…]
+as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{1}\overset{d}{\rightarrow}N(0,1)$.
 (ii) $\mathbb{W}_{1}$ provides a consistent test.
-[…]
+(iii) Under the sequence of local alternatives $H_{\ell 1}$, $\mathbb{W}_{1}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
 3.  SAR with error spatial dependence
@@
 $\epsilon_{i}\equiv\epsilon_{in}=\sum_{j=1}^{n}b_{ij}v_{j},$ (3.1)
-where the $v_{j}$ are $i.i.d$ with mean 0 and variance 1, and the $b_{ij}\equiv b_{ijn}$ can depend on $n$.
+where the $v_{j}$ are $i.i.d$ with mean 0 and variance 1, and $b_{ij}\equiv b_{ijn}$ can depend on $n$.
 This is the Kelejian and Prucha (2007) approach to error spatial dependence, and indeed we will employ the spatial heteroskedasticity and autocorrelation consistent (SHAC) method pioneered in that paper.
-Our imposition that $z_{i}$ be uncorrelated with $\epsilon_{i}$ now, and subsequently, requires that $z_{i}$ is uncorrelated with all the $v_{j}$, $j=1,\ldots,n$.
+This type of linear process assumption is widely employed in the literature, see e.g.
+Robinson (2011), Robinson and Thawornkaiwong (2012), Hidalgo and Schafgans (2017) and Conley et al.
+(2023).
 3.1 Test statistic
@@
 Assumption 8.
-$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel function that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
+$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
 Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g.
-Kelejian and Prucha (2007), Delgado and Robinson (2015) and Gupta and Qu (2024) for similar assumptions.
+Kelejian and Prucha (2007) and Delgado and Robinson (2015) for similar assumptions.
 Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g.
@@
 Then, the $(r,s)$-th element of $\hat{\Xi}_{1}$ is
-$\hat{\Xi}_{1,rs}=n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{ri}k_{sj}\hat{u}_{j}\hat{u}_{j}\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),r,s=1,\ldots,J_{1}.$ (3.3)
+$\hat{\Xi}_{1,rs}=n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{ri}k_{sj}\hat{u}_{i}\hat{u}_{j}\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),r,s=1,\ldots,J_{1},$ (3.3)
+where $\hat{u}_{i}$ are elements of the estimation residual vector $\hat{u}=y-L\hat{\xi}$.
 3.2 Asymptotic properties
-In all subsequent lemmas and theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9.
-In Lemma 3.1 and Theorems 3.1-3.2, which follow, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$.
+In all subsequent lemmas/theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9.
+In Lemma 3.1 and Theorems 3.1-3.2, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$.
 Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e.
@@
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\xi}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (3.4)
-Then,
-$\big{\|}\hat{\Xi}_{1}-\Xi_{1}\big{\|}=o_{p}(1).$
-The condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $d_{\xi}/n\rightarrow 0$ by Assumption 9(a).
-Now, define $\mathcal{M}_{2}=\frac{1}{n}B^{\prime}K_{1}\mathcal{V}_{2}K_{1}^{\prime}B$, with $\mathcal{V}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$.
-Then we have the following theorem.
+Then, $\big\|\hat{\Xi}_{1}-\Xi_{1}\big\|=o_{p}(1)$.
+Define $\mathcal{M}_{2}=\frac{1}{n}B^{\prime}K_{1}\mathcal{V}_{2}K_{1}^{\prime}B$, with $\mathcal{V}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$.
+We have the following theorem.
 Theorem 3.1.
@@
 $\mathbb{W}_{2}-\frac{v^{\prime}\mathcal{M}_{2}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
-To establish the asymptotic distribution of $\mathbb{W}_{2}$, the following assumption is imposed to check a Lyapunov condition, as also e.g. by Delgado and Robinson (2015) and Gupta and Qu (2024).
+To derive the asymptotic distribution of $\mathbb{W}_{2}$, we need the following assumption, as also in Delgado and Robinson (2015).
 Assumption 10.
@@
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}+n^{\eta-1}d_{\xi}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,$ (3.5)
-as $n\rightarrow\infty$, the following hold:
-[…]
+as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{2}\overset{d}{\rightarrow}N(0,1)$.
 (ii) $\mathbb{W}_{2}$ provides a consistent test.
-[…]
-Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $\frac{d_{\xi}^{3}}{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
+(iii) Under the sequence of local alternatives $H_{\ell 2}$, $\mathbb{W}_{2}\overset{d}{\rightarrow}N(2^{-1/2},1)$.
+Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies ${d_{\xi}^{3}}/{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
 4.  Misspecification robustness
-We now consider a degree of misspecification robustness in (2.1), which allows $\epsilon_{i}$ to have the structure in Section 3.
-Focusing on the case $d_{\lambda}=1$ for notational simplicity, and writing $W_{1}=W$, (2.1) imposed the parametric form $\lambda Wy$ and took $W$ as known in the “spatial lag” term.
-We now allow the spatial lag in outcomes to take the nonparametric form $Gy$, where $G$ has elements $g\left(d_{rs}\right)$, $r,s=1,\ldots,n$, for some vector of exogenous (independent of $v_{j}$, $j=1,\ldots,n$) economic distance measures $d_{rs}$.
+Consider a degree of misspecification robustness in (2.1), where $\epsilon_{i}$ has the structure in Section 3.
+Focusing on the case $d_{\lambda}=1$ for notational simplicity, and writing $W_{1}=W$, (2.1) imposed the parametric form $\lambda Wy$ and took $W$ as known in this ``spatial lag'' term.
+We now allow this to take the nonparametric form $Gy$, where $G$ has elements $g\left(d_{rs}\right)$, $r,s=1,\ldots,n$, for some unknown function $g(\cdot)$, and a vector of exogenous (independent of $v_{j}$, $j=1,\ldots,n$) economic distance measures $d_{rs}$.
 This idea was introduced by Pinkse et al.
@@
 4.1 Test statistic
-We approximate the elements $g\left(d_{ij}\right)$ of $G$ with a series of basis functions. Following Pinkse et al. (2002), we suppose
+Following Pinkse et al. (2002), we approximate $g\left(d_{ij}\right)$ with a series of basis functions
 $g(d_{ij})=\sum_{l=1}^{\infty}\tau_{l}e_{l}(d_{ij}),$ (4.1)
-where $\tau_{l}$ are unknown coefficients, and the $e_{l}$ form a basis of the function space to which $g$ belongs.
+where $\tau_{l}$ are unknown coefficients, and $e_{l}$ form a basis of the function space to which $g(\cdot)$ belongs.
 Let $c_{i}^{\prime}$ be the $i$-th row of the $n\times d_{\tau}$ matrix $\mathfrak{C}$ with typical $(i,l)$-th element $\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $\tau=\left(\tau_{1},\ldots,\tau_{d_{\tau}}\right)^{\prime}$, $d_{\tau}\rightarrow\infty$ as $n\rightarrow\infty$.
-Then we generalize (2.1) to
+Then (2.1) is extended to
 $y_{in}=g_{in}^{\prime}y+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (4.2)
-and (2.3) to
+where $g_{in}\equiv g_{i}$ is the $i$-th row of $G$, and, once again dropping $n$-subscripting, accordingly extend (2.3) to
 $y_{i}=c_{i}^{\prime}\tau+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (4.3)
 with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$.
-In the matrix notation, we now have
+In matrix notation, we now have
 $y=\mathfrak{C}\tau+X\beta+\Psi\alpha+u=F\theta+u,$ (4.4)
-where $F=[\mathfrak{C},X,\Psi]$ and $\theta=\left(\tau^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. Our test is implemented using the 2SLS estimator
+where $F=[\mathfrak{C},X,\Psi]$ and $\theta=\left(\tau^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. Our test is based on the 2SLS estimator
 $\hat{\theta}=\left(F^{\prime}{P_{K_{2}}}F\right)^{-1}F^{\prime}{P_{K_{2}}}y,$ (4.5)
@@
 $\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\theta}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (4.7)
-Then,
-$\bigg{\|}\hat{\Xi}_{2}-\Xi_{2}\bigg{\|}=o_{p}(1).$
+Then, $\big\|\hat{\Xi}_{2}-\Xi_{2}\big\|=o_{p}(1)$.
 4.2 Asymptotic properties
@@
 $\mathbb{W}_{3}-\frac{v^{\prime}\mathcal{M}_{3}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
-Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix.
-The following assumption generalizes Assumption 5 to the setting of this section.
+Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proofs.
+The following assumption generalizes Assumption 5.
 Assumption 14.
-$\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $d_{rs},r,s=1,\ldots,n$.
+$\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely and uniformly over the support of the distance measure vector.
 Theorem 4.2.
@@
 $\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\theta}^{3}}{n}+n^{\eta-1}d_{\theta}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,$ (4.8)
-as $n\rightarrow\infty$, the following hold:
-[…]
-(2) $\mathbb{W}_{3}$ provides a consistent test.
-[…]
+as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N(0,1)$.
+(ii) $\mathbb{W}_{3}$ provides a consistent test.
+(iii) Under the sequence of local alternatives $H_{\ell 3}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
 5.  Varying spatial coefficients
-In this section we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but now allow the spatial lag parameters $\lambda_{j}$ to vary.
-The model we consider is closely related to Malikov and Sun (2017) and Sun and Malikov (2018), with the difference that we do not study panel data settings.
-Instead, we allow $d_{\beta}\rightarrow\infty$ and general spatial dependence in the disturbances.
-Our model is also related to the one studied by Sun (2024), wherein the model includes a nonparametric nuisance regression function but does not feature dependent disturbances.
+Now we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but allow spatial lag parameters $\lambda_{j}$ to vary.
+The model is closely related to Malikov and Sun (2017) and Sun and Malikov (2018), with the difference that we do not study panel data settings.
+Instead, we allow $d_{\beta}\rightarrow\infty$ and general spatial dependence in the errors.
+Our model also relates to Sun (2024), wherein the model includes a nonparametric nuisance regression function but does not feature dependent errors.
 Thus, our model complements the extant literature.
@@
 where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$.
-Defining $\Lambda_{j}(z)=diag\left(\lambda_{j}\left(z_{1}\right),\ldots,\lambda_{j}\left(z_{n}\right)\right)$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
+Defining $\Lambda_{j}(z)=diag(\lambda_{j}\left(z_{1}\right),\ldots,\\ \lambda_{j}\left(z_{n}\right))$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
 $y=S(z)^{-1}\left(X\beta+t(z)+\epsilon\right),$ (5.3)
@@
 $H_{0}:\vartheta=0,$ (5.4)
-where $\vartheta$ is either $\mu$ or $\alpha$ and our test will be implemented using the 2SLS estimator
+where $\vartheta$ is either $\mu$ or $\alpha$ and our test will again be implemented using the 2SLS estimator
 $\hat{\gamma}=\left(G^{\prime}{P_{K_{3}}}G\right)^{-1}G^{\prime}{P_{K_{3}}}y,$ (5.5)
@@
 Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$.
-Then, the test statistic is
+The test statistic is
 $\mathbb{W}_{4}=\frac{n\hat{\vartheta}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}-d_{\vartheta}}{\sqrt{2d_{\vartheta}}},$ (5.6)
-where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$.
-Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
+where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$. Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
 Assumption 15.
@@
 Assumption 17.
-$n^{-1}K_{3}^{\prime}K_{3}$ and $\sup_{z}n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1.
+$n^{-1}K_{3}^{\prime}K_{3}$ and $n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1, the latter uniformly in $z\in\mathcal{Z}$.
 Lemma 5.1.
@@
 $\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (5.7)
-Then, we have
-$\bigg{\|}\hat{\Xi}_{3}-\Xi_{3}\bigg{\|}=o_{p}(1).$
+Then, we have $\big\|\hat{\Xi}_{3}-\Xi_{3}\big\|=o_{p}(1)$.
 5.2 Asymptotic properties
@@
 $\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\gamma}^{3}}{n}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,$ (5.8)
-as $n\rightarrow\infty$, the following hold:
-[…]
+as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{4}\overset{d}{\rightarrow}N(0,1)$.
 (ii) $\mathbb{W}_{4}$ provides a consistent test.
-[…]
+(iii) Under the sequence of local alternatives $H_{\ell 4}$, $\mathbb{W}_{4}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
 6.  Empirical applications
 6.1 Testing for the CRS in a production function
-In this subsection, we are interested in studying whether the production function in China’s nonmetal mineral manufacturing industry has a constant returns to scale (CRS) technology.
+In this subsection, we are interested in testing whether the production function in China's nonmetal mineral manufacturing industry has a constant returns to scale (CRS) technology.
 CRS refers to a scenario where a simultaneous proportional increase in all inputs results in an identical proportional increase in output.
-Returns to scale is a classical concept in economics to assess the efficiency of a production function, and it has accordingly received much attention in the empirical literature (for recent works, see Ackerberg et al., 2015; Basu and Bundick, 2017; Attanasio et al., 2020; Combes et al., 2021).
+Returns to scale is a classical concept in economics to assess the efficiency of a production function, and it has accordingly received much attention (see Ackerberg et al.
+(2015); Basu and Bundick (2017); Attanasio et al.
+(2020); Combes et al.
+(2021)).
 Existing works, however, mostly do not account for cross-sectional dependence between firms.
-The role of cross-firms correlation has recently been highlighted empirically by Iyoha (2023), who shows that this can capture spillover effects and would lead to model misspecification if left unaccounted for.
-We consider models that generalize Li et al.
-(2002)’s production function by incorporating spatial dependence in the analysis, thereby allowing for cross-sectional dependence.
-Their analysis centers on a Cobb-Douglas style production function that allows the elasticities of capital and labor to be nonparametrically varying with management expenses.
+The role of cross-firm correlation has recently been highlighted empirically by Iyoha (2023), who shows that this can capture spillover effects and would lead to model misspecification if left unaccounted for.
+Our models generalize Li et al.
+(2002)'s production function by incorporating spatial dependence.
+Their analysis centers on a Cobb-Douglas production function that allows elasticities of capital and labor to be nonparametrically varying with management expenses.
 Management expenses are costs that are indirectly associated with output production, such as research and development (R&D), equipment upgrades and employee training.
-In this framework, CRS corresponds to the elasticities of capital and labor summing to one.
-Using data from the Third Industrial Census of China (conducted by the National Statistical Bureau in 1995) on the non-metal mineral manufacturing industry, Li et al.
+CRS corresponds to the elasticities of capital and labor summing to one.
+Using data from the Third Industrial Census of China (conducted by the National Statistical Bureau in 1995), Li et al.
 (2002) find that ignoring the varying nature of these coefficients leads to an underestimation of the returns to scale.
-Moreover, they suggest their estimate of the returns to scale of the technology is close to being constant for most values of managerial expense, but they do not have a test for the CRS hypothesis.
-Our data comes from the Chinese Industrial Enterprise Database for the year of 2014 as this is the most recent information available in the database.
-We use data on all firms in the nonmetal mineral manufacturing industry with the survey code 31, resulting in a sample of 7,355 observations.
+They suggest their estimate of returns to scale of the technology is close to being constant for most values of managerial expense, but they do not have a test for the CRS hypothesis.
+Our data comes from the Chinese Industrial Enterprise Database in 2014 as it is the most recent information available.
+We use all firms in the nonmetal mineral manufacturing industry with the survey code 31, resulting in a sample of 7,355 observations.
 As advocated by Li et al.
-(2002), the appeal for focusing on this industry is that production functions of firms are expected to be homogeneous since it has a very small proportion of foreign entity ownership.
-This is important because evidence suggests that production performance may vary across ownership types (Murakami et al., 1994).
-Table 1 introduces all variables after taking logarithms.
-Our variables are output ($y_{i}$), capital ($p_{i1}$), labor ($p_{i2}$), and management expenses ($z_{i}$).
-As in Li et al.
-(2002), we include a parametric benchmark where the Cobb-Douglas production function has a quadratic form of management expenses without the varying coefficient entering additively, namely:
+(2002), production functions of firms in this industry are expected to be homogeneous since it has a very small proportion of foreign entity ownership.
+This is important because evidence suggests that production performance may vary across ownership types (Murakami et al.
+(1994)).
+Variables are in logs: output ($y_{i}$), capital ($p_{i1}$), labor ($p_{i2}$), and management expenses ($z_{i}$). As in Li et al. (2002), we include a parametric benchmark model:
 $y_{i}=\delta_{0}+\delta_{1}p_{i1}+\delta_{2}p_{i2}+\beta_{1}z_{i}+\beta_{2}z_{i}^{2}+\epsilon_{i}.$ (6.1)
-The elasticities of capital and labor are $\delta_{1}$ and $\delta_{2}$, respectively.
-The semiparametric production function where the elasticities of of capital and labor vary with management expenses is:
+Elasticities of capital and labor are $\delta_{1}$ and $\delta_{2}$, respectively. The semiparametric model where the elasticities of of capital and labor vary with management expenses is:
 $y_{i}=\delta_{0}(z_{i})+\delta_{1}(z_{i})p_{i1}+\delta_{2}(z_{i})p_{i2}+\epsilon_{i}.$
-Based on the CRS hypothesis, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
+Based on the CRS, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
 $y_{i}^{*}=\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i},$
-where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate
-$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.2)
-where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for $k=1,...,h/2$.
-We approximate the unknown functions by polynomial and trigonometric basis functions of different orders.
-We define the series approximation based Wald statistic derived from this model with no spatial autoregressive component as $\mathbb{W}_{0}$.
+where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate the following model with no SAR structure:
+$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
+where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for k=1,,h/2k=1,...,h/2.
+We approximate the unknown functions by polynomial and trigonometric functions of different orders.
+We define the series approximation based Wald statistic derived from this model as $\mathbb{W}_{0}$.
 Next, we incorporate the SAR structure and consider:
@@
 We estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
-$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.3)
+$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
 with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$.
+Let $d_{ij,m}^{\ast}$ be the geographical distance between firms $i$ and $j$, and $d_{m}$ represent the 10th percentile of all geographical distances.
 We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
-$\begin{split}&w_{ij}=[W_{n}^{p}]_{ij}=\begin{cases}1\quad\text{iffirms}\ i\ \text{and}\ j\ \text{are in the same prefecture-level city},\\ 0\quad\text{otherwise},\end{cases}\\ &w_{ij}=[W_{n}^{d}]_{ij}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise},\end{cases}\end{split}$
-where $d_{ij,m}^{\ast}$ measures the geographical distance between firms $i$ and $j$, and $d_{m}$ represents the 10th percentile of all geographical distances.
-For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$.
+$\begin{split}w_{ij}^{p}=\begin{cases}1\quad\text{if}\ i\ \text{and}\ j\ \text{are in the same city},\\ 0\quad\text{otherwise}.\end{cases}w_{ij}^{d}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise}.\end{cases}\end{split}$
+For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$.
 For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
-Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for $k=1,...,h$, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function $k=1,...,h/2$ for $h=2,4$.
-To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function to estimate $\mathscr{K}\left({d_{ij,m}^{\ast}}/{d_{m}}\right)$.
-For $\mathbb{W}_{3}$, we denote $\widehat{w}_{ij,l}=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$ if $i\neq j$ for $l=1,...,h$.
-Then, $c_{il}=\sum_{j=1}^{n}\widehat{w}_{ij,l}y_{j}$ for $l=1,...,h$.
-For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$.
+Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for k=1,,hk=1,...,h, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function k=1,,h/2k=1,...,h/2 for $h=2,4$.
+To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function.
+For $\mathbb{W}_{3}$, let $\mathcal{E}_{l}$ be the matrix whose typical entry is, for $i\neq j$ and l=1,,hl=1,...,h, $e_{l}(d_{ij})=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$.
+Then, $c_{il}=\sum_{j=1}^{n}e_{l}(d_{ij})y_{j}$ for l=1,,hl=1,...,h.
+For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=\frac{1}{h}[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\frac{1}{h}\sin(\frac{z_{i}}{2l})$ as the trigonometric function for l=1,,hl=1,...,h.
 Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$.
-For the IVs, we use $K_{1}=[W\Psi,X,\Psi]$ with respect to $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$, and $K_{2}=[X,\Psi]$ and $K_{3}=[H\Psi,X,\Psi]$ with respect to $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$.
-The results from estimating the parametric model can be found in Table 2, and the results for other models with different basis functions are in Tables 3-6.
-In the latter group of tables, we define $\mathscr{T}_{n}$ as the asymptotic p-values corresponding to $\mathbb{W}_{i}$, $i=0,...,4$.
-It is well known that nonparametric tests, which rely on their asymptotic normal null distributions, may not perform well with finite samples.
-As an alternative, we provide bootstrap p-values that are denoted by $\mathscr{T}_{n}^{\ast}$ basing on 150 bootstrap replications.
-The details of the bootstrap procedure are given in Section 7.1.
-Our results show that, in all models, the hypothesis of CRS cannot be rejected.
+For the IVs, we use $K_{1}=K_{3}=[X,\Psi,W_{n}^{p}X,W_{n}^{d}X]$ with respect to $\mathbb{W}_{1}$, $\mathbb{W}_{2}$ and $\mathbb{W}_{4}$, and $K_{2}=[X,\Psi,\mathcal{E}_{1}X,...,\mathcal{E}_{h}X]$ with respect to $\mathbb{W}_{3}$.
+Results with different basis functions are in Table 1.
+In all models, the hypothesis of CRS cannot be rejected.
 These findings corroborate those in Li et al.
@@
 A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
-We end this application with some remarks that, in addition to inference on returns to scale, our analysis also reveals other similarities with prior works.
-Notably, from Table 7, we observe that the parametric method underestimates the returns to scale compared to the semiparametric method.
-[^1] The returns to scale in a semiparametric model can be computed using $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$.
+We end this application with some remarks: (1) In addition to inference on CRS, our analysis also reveals other similarities with prior works.
+Table 2 shows that the parametric method underestimates the returns to scale compared to the semiparametric method.
+[^1] The returns to scale in a semiparametric model is computed by $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$.
 The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses.
-To illustrate this, take the semiparametric estimates with spatial weights $W_{n}^{d}$ and produce Figure 1 and Figure 2.
-The former shows that the output elasticity of capital is increasing in management expenses, and the latter shows the output elasticity of labor is decreasing in management expenses.
-Intuitively, by the nature of management expenses in production, the elasticities may be expected to increase with both inputs.
-However, the decreasing elasticities of labor is a known empirical fact in this industry due to the practice of “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest.
-We also note that the returns to scale is relatively flat in the semiparametric model and is within the 95% bounds for most $z_{i}$, as illustrated in Figure 3.
+Figures 1 (a)-(b) illustrates this with the semiparametric estimates with spatial weights $W_{n}^{d}$.
+They show that the output elasticity of capital is increasing in management expenses, while it is decreasing for labor.
+The decreasing elasticity of labor is a known empirical fact in this industry due to “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest.
+(2) The returns to scale are relatively flat in the semiparametric model and are within the 95% bounds for most $z_{i}$, as illustrated in Figure 1 (c).
 Further discussions on these findings can be found in Li et al.
 (2002).
-The results we have described here are not unique to the estimates from $W_{n}^{d}$.
-Other choices of spatial weights give the same conclusion; we simply selected one to illustrate our findings.
-6.2   Testing the impact of the distance to the employment center on housing price
-In this subsection, we use the Boston house price dataset, which has been used in several spatial studies, to illustrate another application of our proposed test.
-The dataset [^2] The related shp file can be downloaded at https://github.com/simonbrewer/geog6000/blob/main/datafiles/boston.tr.zip.
-Data descriptions are in Table 8.
-consists of the median value of owner-occupied homes in 506 census tracts in the Boston Standard Metropolitan Statistical Area in 1970 (Harrison Jr and Rubinfeld, 1978).
-Specifically, we explore how some factors such as the per capita crime rate by town (denoted by CRIM), average number of rooms per dwelling (RM), full-value property-tax rate per \$10,000 dollar (TAX), percentage of lower socioeconomic status population (LSTAT) and index of accessibility to radial highways (RAD) affect the median value of owner-occupied homes in $1000$s (MEDV), and whether the effects of these factors vary over location or not.
-DIS, which represents the weighted distances between a property and the five Boston employment centers that are associated with higher housing prices due to closer proximity to the city center, will drive the varying coefficient.
-The motivation for considering spatial autocorrelation in house prices is straightforward, as housing prices in a specific location or region are commonly influenced by the prices of nearby houses due to factors such as environmental and geographic considerations.
+(3) Our results are not unique to the estimates from $W_{n}^{d}$.
+Other choices of spatial weights give the same conclusion.
+6.2 Testing the impact of the distance on house prices
+We use the Boston house price dataset  [^2] The dataset consists the median value of owner-occupied homes in 506 census tracts in the Boston Standard Metropolitan Statistical Area in 1970 (Harrison Jr and Rubinfeld (1978)), which is available at https://github.com/simonbrewer/geog6000/blob/main/datafiles/boston.tr.zip.
+to explore how some factors affect the median value of owner-occupied homes in $1000$s (MEDV), and whether the effects of these factors vary over location.
+Factors include the per capita crime rate by town (denoted by CRIM), average number of rooms per dwelling (RM), full-value property-tax rate per \$10,000 dollar (TAX), percentage of lower socioeconomic status population (LSTAT) and index of accessibility to radial highways (RAD).
+The location variable DIS, which represents the weighted distances between a property and the five Boston employment centers, will drive the varying coefficient.
+Let $y_{i},x_{i1},x_{i2},p_{i1},p_{i2},p_{i3},z_{i}$ denote the logarithm of MEDV, RAD, LSTAT, CRIM, RM, TAX and DIS respectively.
+As house prices in a specific location or region are commonly influenced by the prices of nearby houses due to factors such as environmental and geographic considerations, it is natural to consider the spatial autocorrelation.
 [^3] There are also some studies that use the SAR semiparametric model with the Boston house price data, for example, Malikov and Sun (2017) and Li et al.
@@
 (2014) proposed a semiparametric coefficient varying spatial model and derived the distribution theory of their estimator.
-While they do not have inference results on functions of the random coefficients, they have a model selection procedure to determine the parametric and nonparametric components in their model.
+While they do not have inference results on functions of the random coefficients, they have a model selection procedure to determine the parametric and nonparametric components.
 We thus take the model they selected as our starting point.
-Let $y_{i},x_{i1},x_{i2},p_{i1},p_{i2},p_{i3},z_{i}$ denote the log of MEDV, RAD, LSTAT, CRIM, RM, TAX and DIS respectively.
 Guided by model selection results, Sun et al.
-(2014) argue that DIS doesn’t affect LSTAT and RAD.
-Intuitively, for LSTAT, the geographical distribution of employment opportunities does not necessarily correlate with a community’s socio-economic status; a community might be far from employment centers but could have good public transportation links or a high number of remote work opportunities, mitigating the impact on the community’s lower socioeconomic status.
+(2014) argue that DIS doesn't affect LSTAT and RAD.
+Intuitively, for LSTAT, the geographical distribution of employment opportunities does not necessarily correlate with a community's socio-economic status; a community might be far from employment centers but could have good public transportation links or a high number of remote work opportunities, mitigating the impact on the community's lower socioeconomic status.
 For RAD, the accessibility to highways might depend more on urban planning and infrastructure investment rather than the proximity to employment centers.
-The baseline model, based on Sun et al. (2014), is:
+The baseline model, based on Sun et al.
+(2014), is:
 $y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
-We focus on a more general version of their model, by allowing the spatial lag coefficient to be varying with DIS as well:
+We consider a more general version where the spatial lag coefficient varies with DIS:
 $y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
 Therefore, we estimate:
-$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2}\\ &y_{i}=\sum\limits_{j=1}^{n}g(d_{ij})y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4},\end{split}$
-and the estimated models are
-$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4}.\end{split}$ (6.4)
+$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{4}.\end{split}$
 We consider two row-normalized spatial weight matrices.
 The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017).
-The second, denoted by $W_{n}^{t}$, indicates whether two locations are in the same tract.
-For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$.
+The second, $W_{n}^{t}$, indicates whether two locations are in the same tract.
+For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$.
 For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
-From Tables 9-12, we find that $log(DIS)$ indeed affects $p_{i1}$, $p_{i2}$ and $p_{i3}$, and has a significantly nonlinear impact on neighboring house prices.
-All test statistics show a rejection of the null $H_{0}:\alpha=0$.
-The bootstrap p-value also confirms this conclusion.
-In addition, testing with $\mathbb{W}_{4}$ also leads to a rejection of the null $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices.
-Furthermore, we estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$.
-From Table 13, we thus find that the parametric method underestimates the spatial coefficient.
-Our results, which are based on more general models, thus complement those in Sun et al.
+Results are listed in Table 3.
+All tests reject the null $H_{0}:\alpha=0$.
+In addition, testing with $\mathbb{W}_{4}$ leads to a rejection of $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices.
+We further estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$.
+As shown in Panel E of Table 3, the parametric method underestimates the spatial coefficient.
+Thus our results based on more general models complement those in Sun et al.
 (2014).
-7.   Monte Carlo Simulation
-We present a basic set of Monte Carlo results in the main body of the paper.
-Because our theory covers such a wide range of models, which include several important special cases, we have performed a wide range of simulations for them too.
-However, the sheer number of tables we have makes it impractical to include them all in the main body of the paper, and we provide references to an online supplementary appendix.
+7.  Monte Carlo
+We present a basic set of Monte Carlo results in this section and a wider range of simulations in the online appendix. All experiments use 1000 replications.
 7.1 Basic setting
-Taking $n=100,300,500,$ we choose two specifications to generate $y$ with $d_{\delta}=1$:
-$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$ (7.1)
+Taking $n=200,500,900,$ we choose two specifications to generate $y$ with $d_{\delta}=1$:
+$y_{i}=\sum_{k=1}^{d_{\lambda}}\lambda_{k}\sum\limits_{j=1}^{n}w_{kij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$
 and
-$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},i=1,\ldots,n,$ (7.2)
-[…]
-Kelejian and Prucha (2007) consider the kernel function with a distance measure $d_{ij,m}^{*}$ and a threshold $d_{m}$.
-In estimation, we use the Epanechnikov kernel function, we choose the Euclidean distance $d_{ij}$ mentioned in $W_{1}$ to represent $d_{ij,m}^{*}$, and $d_{m}$ is considered as the quantile value of 10% of all distances.
-If $d_{ij,m}^{*}>d_{m}$, $\mathscr{K}\left(\frac{d_{ij,m}^{*}}{d_{m}}\right)=0$.
-We also consider $d_{\lambda}=2$ with $W_{1}$, $W_{2}$, and $\lambda=(0.5,0.3)^{\prime}$.
-The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[W_{1}X,W_{2}X,X,\Psi]$.
-We calculate the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%.
-We consider 1000 Monte Carlo replications for all experiments and $B=150$ bootstrap replications.
-We define $\mathscr{S}_{n}$ as the test statistic corresponding to $\mathbb{W}_{i}$ for $i=1,\ldots,4$ and $\mathscr{S}^{\ast}_{n}$ as the bootstrap version of $\mathscr{S}_{n}$.
-The bootstrap version of our test, e.g., focusing on the setting described in Section 3, can be carried out by taking these steps:
-  • Compute $\widehat{\epsilon}=y-L\hat{\xi}$ under null or alternative and obtain $\widetilde{\epsilon}=(I_{n}-\frac{1}{n}l_{n}l_{n}^{\prime})\widehat{\epsilon}$, where $l_{n}$ is the $n\times 1$ column vector of ones.
-  • Sample randomly with replacement $n$ times from elements of $\widetilde{\epsilon}$ to obtain $\tilde{\epsilon}^{*}$.
-  • Generate the bootstrap sample $\hat{y}=\left(I_{n}-\sum\limits_{j=1}^{d_{\lambda}}\hat{\lambda}_{j}W_{j}\right)^{+}\left([X,\Psi](\hat{\beta},\hat{\alpha})^{\prime}+\tilde{\epsilon}^{*}\right)$, where $(\cdot)^{+}$denotes the Moore-Penrose generalized inverse.
-  • For the bootstrap, compute $\hat{\xi}^{\ast}$ and obtain the bootstrap residual $\hat{\epsilon}^{\ast}=\hat{y}-\hat{L}\hat{\xi}^{\ast}$ to construct $\mathscr{S}_{n}^{\ast}$, where $\hat{L}=[W_{1}\hat{y},\ldots,W_{d_{\lambda}}\hat{y},X,\Psi]$.
-  • Repeat $B$ times to obtain the sequence of bootstrap test statistics $\left\{\mathscr{S}_{nj}^{\ast}\right\}_{j=1}^{B}$.
-Calculate bootstrap p-values $\mathscr{T}_{n}^{\ast}=\frac{1}{B}\sum\limits_{j=1}^{B}\mathbf{1}\left(\mathscr{S}_{n}<\mathscr{S}_{nj}^{\ast}\right)$ and reject the null hypothesis if $\mathscr{T}_{n}^{\ast}$ is smaller than the prescribed nominal level of significance.
-As before, we denote by $\mathscr{T}_{n}$ the p-value based on the asymptotic test.
-From Tables 14-15, we find that the bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases.
-Specifically, we observe that the pairs $(n,h)=(500,8)$ deliver acceptable sizes under three disturbances.
-Moreover, we observe that the asymptotic test is powerful against local alternatives as $n$ increases from Tables S1-S2 in the supplement.
-We also consider the case of $d_{\lambda}=2$ in the supplement.
-The size performances [^5] Power performances are close to or equal to 1 in most cases so we omit the power tables but they are available on request.
-are reported in Tables S3-S4 in the supplement and the finite sample properties are similar to the case of $d_{\lambda}=1$.
+$y_{i}=\sum_{k=1}^{d_{\lambda}}\lambda_{k}\sum\limits_{j=1}^{n}w_{kij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+v_{i},\ v_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}+\epsilon_{i},i=1,\ldots,n,$
+where $\epsilon_{i}$ are $i.i.d$ with three zero mean and unit variance distributions: (V1) $N(0,1)$, (V2) $\sqrt{\frac{5}{4}}t(10)$ and (V3) $\frac{1}{4}(\chi_{8}^{2}-8)$.
+We generate $z_{i}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$, $p_{i}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-2,2]$, $x_{i1}=1$, $x_{i2}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(1,2)$ and set $\beta=(-1,1)^{\prime}$.
+For $\lambda$, we fix the total weight at 0.9 and then define a decreasing vector $\lambda^{*}=(d_{\lambda},d_{\lambda}-1,...,1)^{\prime}$.
+We then set $\lambda=0.9\left(\sum_{k=1}^{d_{\lambda}}\lambda_{k}^{*}\right)^{-1}\lambda^{*}$.
+For the spatial weights matrices, we generate $W_{k}$ as circulants.
+Specifically, let $W_{k}^{*}$ be the symmetric circulant matrix whose first-row entries are $w_{k1j}^{*}=0$ if $j=1$ or $j=k+2,\ldots,n-k$ and $w_{k1j}^{*}=1$ if $j=2,\ldots,k+1$ or $j=n-k+1,\ldots,n$.
+Thus, the weight matrix $W_{k}^{*}$ encapsulates a binary neighbourhood criterion for $k$ neighbours on either `side' of a unit.
+Now define the normalized matrix $W_{k}=\{\overline{\alpha}(W_{k}^{*})\}^{-1}W_{k}^{*},$ recalling that $\overline{\alpha}(W_{k}^{*})=2k$ for a circulant matrix.
+Furthermore, $b_{ij}$ is the typical element of $(I_{n}-\sum_{k=1}^{d_{\lambda}}\lambda_{k}W_{k})^{-1}$.
+To construct the SHAC statistics $\mathbb{W}_{2},\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we proceed as follows.
+Generate $M=d_{\lambda}$ distinct distance measures by setting $\ell=[n^{\eta}]+1$ with $\eta=3/7$ and taking $q=8$, so that each unit has at least $\ell$ neighbors, where $\eta$ is defined in Assumption 9 (a).
+[^4] From the symmetric circulant matrix $W_{d_{\lambda}}$, we build a graph and compute its pairwise shortest‐path distances in MATLAB, then apply cmdscale to generate $n\times 2$ coordinate matrix.
+We next compute the full $n\times n$ Euclidean distance matrix $D$ with the typical elements $d_{ij}$.
+For each m=1,,Mm=1,...,M, we add the actual distance matrix, $D_{m}^{*}$, used in practice with unobserved symmetric measurement noise.
+Its typical entries are $d_{ij,m}^{*}=d_{ij}+\nu_{ij,m},$ where $\nu_{ij,m}=1/2(\mu_{ij,m}+\mu_{ji,m})$ with $\mu_{ij,m}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$.
+Define for each unit $i$ the $\ell$-th nearest‐neighbor distance by $d_{(i,\ell),m}=$ the $\ell$-th smallest entry of row $i$ of $D^{*}_{m}$, excluding the diagonal.
+We then set $d_{m}=\max_{1\leq i\leq n}d_{(i,\ell),m},$ so that every unit has at least $\ell$ neighbors within distance $d_{m}$.
+Collecting these over $m=1,\dots,M$ yields the $M\times 1$ vector $d_{m}$.
+We set $\delta=0$ for the null hypothesis and $\delta(\mathrm{x})=1-\mathrm{x}^{2}$ for any $\mathrm{x}\in R$ for the alternatives.
+We choose polynomials $\psi_{j}(\mathrm{x})=\mathrm{x}^{j}$ for $j=1,\cdots,h$ and the design of trigonometric functions is in the supplement.
+The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[X,W_{1}X,...,W_{d_{\lambda}}X,\Psi]$.
+Furthermore, asy-p represents the standard-normal p-values, whereas chi-p represents the chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$, which is to improve testing performance in small samples.
+This may also help in empirical applications, although in both applications in the previous section critical values from either distribution yield the same results.
+We discuss size performance in this section and present power performance in the supplement, since under global alternatives the power of the asymptotic tests tends to one as $n\to\infty$.
+Table 4 reports the empirical sizes of $\mathbb{W}_{1}$, and Table 5 reports those of the SHAC-corrected statistic $\mathbb{W}_{2}$, at nominal 1%, 5% and 10% levels.
+When $n=200$, $\mathbb{W}_{2}$ exhibits substantially larger size distortion than $\mathbb{W}_{1}$.
+The reason is: Theorem 2.1 requires $n^{-1}\,d_{\xi}\to 0,$ whereas Theorem 3.1 requires $n^{\eta-1}\,d_{\xi}\to 0$ with $\eta=3/7$.
+In our design $d_{\xi}=d_{\lambda}+d_{\beta}+d_{h}$ with $d_{\lambda}\in\{2,4\}$, $d_{\beta}=2$, $d_{h}\in\{2,4,8\}$, so $d_{\xi}/n$ is already small at $n=200$, but $n^{\eta-1}d_{\xi}^{2}$ is relatively large.
+As $n$ increases to 500 and 900, $n^{\eta-1}d_{\xi}^{2}$ shrinks and the size of $\mathbb{W}_{2}$ converges rapidly to its nominal levels.
+Table 4-5 also show a clear pattern across the polynomial order with $h\in\{2,4,8\}$.
+For both $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$, the smallest sieve yields the most accurate sizes, the largest sieve the greatest distortion, particularly for $\mathbb{W}_{2}$ at $n=200$.
+This accords with the fact that larger $h$ increases $d_{\alpha}$ and hence $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$, which in turn slows the convergence rates $n^{-1}d_{\xi}\to 0$ for $\mathbb{W}_{1}$ and $n^{\eta-1}d_{\xi}^{2}\to 0$ for $\mathbb{W}_{2}\,$.
+A similar scenario obtains when increasing $d_{\lambda}$, since it also raises $d_{\xi}$ and worsens finite-sample size accuracy.
 7.2 Nonparametric spatial matrix
-For the nonparametric spatial matrix, we generate $w_{1ij}=\Phi\left(-b_{ij}\right)\mathbf{1}\left(c_{ij}<0.1\right)$ if $i\neq j$, and $w_{1ii}=0$, where $\Phi(\cdot)$ is the standard normal cdf, $b_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-3,3]$, and $c_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$.
-From this construction, we ensure that $W_{1}$ is sparse with no more than $10\%$ elements being nonzero.
-Then, $W=W_{1}/1.2\bar{\alpha}(W_{1})$, ensuring the existence of $(I-W)^{-1}$.
-For $\mathbb{W}_{3}$, we approximate elements in $W$ by $\widehat{w}_{ij}=\sum_{l=0}^{r}\tau_{l}[d_{ij,m}^{*}]^{l}\mathbf{1}(d_{ij,m}^{*}<d_{m})$ if $i\neq j$ with $r=2,4$, where $d_{ij,m}^{*}$ and $d_{m}$ are discussed in Section 7.1, and the IV matrix for $\mathbb{W}_{3}$ is $K_{2}=[X,\Psi]$.
-We define this DGP setting as (D3).
-Table 16 reports the simulation results at the 1%, 5% and 10% asymptotic level with $r=2,4$ and $h=2,4,8$.
-We find that there is a clear pattern of rejection rates approaching the theoretical level as sample size increases, and sizes are acceptable for $n=500$ and $h=8$.
-When $r$ increases, the results are similar.
-From Table S5, the power performances are close to or equal to 1 as $n$ increases.
-We also consider $W=W_{1}+W_{2}$ which corresponds to $d_{\lambda}=2$.
-Specifically, we set $w_{ij2}=d_{m}\exp(-d_{ij,m}^{*}/d_{m})$ if $d_{ij,m}^{*}\leq d_{m}$ and $w_{ij2}=0$ if $d_{ij,m}^{*}>d_{m}$.
-From Table S6 in the supplement, a similar pattern for size is observed.
+For the nonparametric spatial matrix, we generate $g^{*}_{ij}=\Phi\left(-b_{ij}\right)\mathbf{1}\left(c_{ij}<0.1\right)$ if $i\neq j$, and $w_{1ii}=0$, where $\Phi(\cdot)$ is the standard normal cdf, $b_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-3,3]$, and $c_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$.
+From this construction, we ensure that $G^{*}$ is sparse with no more than $10\%$ elements being nonzero.
+Then, define $G=G^{*}/1.2\bar{\alpha}(G^{*})$, ensuring the existence of $(I-G)^{-1}$.
+For $\mathbb{W}_{3}$, we approximate elements in $G$ by $\widehat{g}_{ij}=\sum_{l=0}^{d_{\tau}}\tau_{l}e_{l}(d_{ij})=\sum_{l=0}^{d_{\tau}}\tau_{l}[d_{ij,m}^{*}]^{l}\mathbf{1}(d_{ij,m}^{*}<d_{m})$ if $i\neq j$ with $d_{\tau}=2,4$, where $d_{ij,m}^{*}$ and $d_{m}$ are discussed in Section 7.1, and the IV matrix for $\mathbb{W}_{3}$ is $K_{2}=[X,\mathcal{E}_{1}X,...,\mathcal{E}_{d_{\tau}}X,\Psi]$, where $e_{l}(d_{ij})$ are the typical elements of $\mathcal{E}_{l}$.
+Table 6 reports the empirical sizes of $\mathbb{W}_{3}$ at nominal 1%, 5% and 10% levels.
+When $n=200$, $\mathbb{W}_{3}$ with $d_{\tau}=2$ yields rejection rates very close to the nominal levels, whereas the larger basis $d_{\tau}=8$ produces mild over-rejection.
+As $n$ increases to 500 and 900, both choices of $d_{\tau}$ rapidly converge to their nominal sizes.
+This scenario is similar to that of $\mathbb{W}_{2}$ and originates from the rate condition $n^{\eta-1}d_{\tau}^{2}\to 0$ in Theorem 4.1.
+In other words, it also requires an accurate approximation of the SHAC estimator.
 7.3 Varying spatial coefficient
-We now consider varying $\lambda$ such that $\lambda=\lambda(z_{i})$.
-When $d_{\lambda}=1$, we set $\lambda(z_{i})=\sin(0.75\pi\times z_{i})$, when $d_{\lambda}=2$, we set $\lambda_{1}(z_{i})=\sin(0.38\pi\times z_{i})$ and $\lambda_{2}(z_{i})=\cos(0.37\pi\times z_{i})$.
-Others are same as those in Section 7.1.
-For estimation, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$.
-Then, $h_{1il}=y_{i}\sum_{j=1}^{n}w_{1ij}\phi_{jl}(z)$ and $h_{2il}=y_{i}\sum_{j=1}^{n}w_{2ij}\phi_{jl}(z)$.
-The IV matrix is $K_{3}=[H_{1}X,X,\Psi]$ for $d_{\lambda}=1$ and $K_{3}=[H_{1}X,H_{2}X,X,\Psi]$ for $d_{\lambda}=2$.
-We define this DGP setting as (D4).
-Tables 17 and 18 report the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%.
-Sizes are acceptable as $n$ and $h$ increase, especially, the trigonometric approximation performs slightly better than polynomial approximations for Chi-square distribution when testing $\lambda=0$.
-Bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases.
-When $d_{\lambda}=2$, the results are similar, as shown in Tables S7-S8 in the supplement.
-From Tables S9 and S10 in the supplement, power improves as $h$ increases for all designs, and the test for $\lambda$ and $\delta$ have similar results.
-As for $d_{\lambda}=2$, we find that our proposed test statistics still perform well as both $n$ and $h$ increase.
-7.4   Additional simulation results in Online Supplement
-In a range of experiments reported in the online supplement, we study how large $n$ with $h$ increasing to see how size changes.
-So we set $n=1000$ with $h=12,16$, $d_{\lambda}=1,2$ and $d_{\delta}=1$.
-From Table S11 of the supplement, we find that as $h$ increases, the performance of size improves.
-We also consider tests of equality of varying coefficients and CRS hypothesis, where $d_{\delta}=2$ such that
-$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+\epsilon_{i},$ (7.3)
-and
-$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},$ (7.4)
-for $i=1,\ldots,n$.
-Our null of interest when testing equality is $H_{0}^{true}:\delta_{1}(z)=\delta_{2}(z),z\in\mathcal{Z},$ while for CRS it is $H_{0}^{true}:\delta_{1}(z)+\delta_{2}(z)=1,z\in\mathcal{Z}.$ In the first case, we reparameterize with $\theta(z)=\delta_{1}(z)-\delta_{2}(z)$, and the model is $y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i1}+p_{i2})\delta_{2}(z_{i})+\epsilon_{i}.$ In the second case, we reparameterize with $\theta(z)=\delta_{1}(z)+\delta_{2}(z)-1$ and the model is $y_{i}-p_{i1}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i2}-p_{i1})\delta_{2}(z_{i})+\epsilon_{i}.$
-The results are included in the supplement.
-Taking CRS as an example, from Tables S12-S13, bootstrap tests are closer to the nominal levels than the asymptotic tests although the sizes of both types of tests improve generally as the sample size increases.
-For the nonparametric spatial matrix, from Table S14, there is a clear pattern of rejection rates approaching the theoretical level as sample size increases.
-The results of testing equality are similar, as shown in Tables S15-S17.
+We now consider varying $\lambda$ such that $\lambda=\lambda(z_{i})$ and $d_{\lambda}=1,2$.
+When $d_{\lambda}=1$, we set $\lambda(z_{i})=0.9\sin(\pi z_{i})$, when $d_{\lambda}=2$, we set $\lambda_{1}(z_{i})=0.6\sin(\pi z_{i})$ and $\lambda_{2}(z_{i})=0.3\sin(\pi z_{i})$.
+Others are the same as those in Section 7.1.
+For estimation, we use $\phi_{il}(z)=\frac{1}{h}[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\frac{1}{h}\sin(\frac{z_{i}}{2l})$ as the trigonometric function for l=1,,hl=1,...,h.
+The IV matrix $K_{3}=K_{1}$ with $d_{\lambda}=1,2$.
+Table 7 reports the rejection probabilities of $\mathbb{W}_{4}$ at nominal 1%, 5% and 10% levels when testing $\delta(z)$.
+At $n=200$, the smallest sieve $(h=2,\,d_{\lambda}=1)$ achieves sizes close to nominal, whereas larger sieves, especially $h=8,\,d_{\lambda}=2$, over‐reject noticeably.
+As $n$ increases to 900, size distortion vanishes rapidly.
+Table S1 shows the sizes of $\mathbb{W}_{4}$ when testing $\lambda(z)$.
+In particular, at $n=200$ with $d_{\lambda}=2$ and $h=8$, the over‐rejection is especially severe.
+By $n=900$, size accuracy is largely restored.
+This pattern reflects the rate condition in Theorem 5.1, $n^{\eta-1}d_{\gamma}^{2}\to 0$, since both $h$ and $d_{\lambda}$ enter the total sieve dimension $d_{\gamma}$, slowing the approximation when they are large.
+Across all designs, the chi-p delivers empirical sizes closer to nominal levels than asy-p in small samples, while the difference dissipates as $n$ grows.
+Results for trigonometric functions are similar, and left for the online appendix.
+8.  Conclusion
+We provide a machinery for conducting tests on nonparametrically varying coefficients using ideas from parametric testing, and with similar ease of implementation.
+We allow for spatial dependence and provide for testing nonparametrically varying spatial dependence coefficients.
+It seems reasonable to conjecture that the methodology is applicable in other settings where one may be interested in testing linear restrictions on nonparametric functions.
+These provide directions for future research.
 Appendix
@@
 $\mathscr{A}=B^{\prime}\mathscr{S}B,$ (A.1)
-where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7.
-(b) The entries $s_{ij}$ of the symmetric matrix $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$.
+where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7 and $\mathscr{S}$ is a symmetric matrix.
+(b) The entries $s_{ij}$ of $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$.
 (c) $\mathscr{S}$ satisfies 1.
@@
 Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 10. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
-[…]
+${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}\overset{d}{\longrightarrow}N(0,1).$ (A.2)
 Proof.
-The proof is identical to that of Theorem 4.3 in Gupta and Qu (2024) and is omitted. ∎
+${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}=\sum_{s=1}^{n}\rho_{s}$, where
+$\rho_{s}=L^{-1}a_{ss}\left(v_{s}^{2}-1\right)+2L^{-1}\mathbf{1}(s\geq 2)v_{s}\sum_{t<s}a_{st}v_{t},$
+writing $L=\sqrt{2J}$. From Scott (1973), (A.2) follows if
+$\sum_{s=1}^{n}\mathrm{E}\rho_{s}^{4}\overset{p}{\longrightarrow}0,\text{ as }n\rightarrow\infty,$ (A.3)
+$\text{and }\sum_{s=1}^{n}\left[\mathrm{E}\left(\rho_{s}^{2}\left.{}\right|v_{t},t<s\right)-\mathrm{E}\left(\rho_{s}^{2}\right)\right]\overset{p}{\longrightarrow}0,\text{ as }n\rightarrow\infty.$ (A.4)
+We show (A.3) first. Evaluating the LHS and using Assumption A yields
+$\mathrm{E}\rho_{s}^{4}$ $\leq$ $CL^{-4}a_{ss}^{4}+CL^{-4}\sum_{t<s}a_{st}^{4}\leq CL^{-4}\left(\sum_{t\leq s}a_{st}^{2}\right)^{2}\leq CL^{-4}\left(b_{s}^{\prime}\mathscr{S}\sum_{t\leq s}b_{t}b_{t}^{\prime}\mathscr{S}b_{s}\right)^{2}$
+$\leq$ $CL^{-4}\left(b_{s}^{\prime}\mathscr{S}^{2}b_{s}\right)^{2}=CL^{-4}\sum_{i,j,k=1}^{n}b_{is}b_{ks}s_{ij}s_{kj}$
+$\leq$ $CL^{-4}\sum_{i,k=1}^{n}\left|b_{is}\right|\left|b_{ks}\right|\sum_{j=1}^{n}\left(s^{2}_{ij}+s^{2}_{kj}\right)=O_{p}\left(L^{-4}Jn^{-1}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right),$
+so $\sum_{s=1}^{n}\mathrm{E}\rho_{s}^{4}=O_{p}\left(L^{-4}Jn^{-1}\sum_{s=1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right)=O_{p}\left(L^{-4}J\right)$.
+Thus (A.3) is established.
+Note that $\mathrm{E}\left(\left.\rho_{s}^{2}\right|v_{t},t<s\right)=4L^{-2}\left\{\left(\mu_{4}-1\right)a_{ss}^{2}+2\mu_{3}\mathbf{1}(s\geq 2)\sum_{t<s}a_{st}a_{ss}v_{t}\right\}+4L^{-2}\mathbf{1}(s\geq 2)\left(\sum_{t<s}a_{st}v_{t}\right)^{2},$ where $\mathrm{E}v_{t}^{j}=\mu_{j},j=3,4$, and $\mathrm{E}\rho_{s}^{2}=4L^{-2}\left(\mu_{4}-1\right)a_{ss}^{2}+4L^{-2}\mathbf{1}(s\geq 2)\sum_{t<s}a_{st}^{2},$ so that (A.4) is bounded by
+$CL^{-2}\sum_{s=2}^{n}\sum_{t<s}a_{st}a_{ss}v_{t}+C\left\{\sum_{s=2}^{n}\left(\sum_{t<s}a_{st}v_{t}\right)^{2}-\sum_{t<s}a_{st}^{2}\right\}.$ (A.5)
+By transforming the range of summation, the square of the first term in (A.5) has expectation bounded by
+$CL^{-4}\mathrm{E}\left(\sum_{t=1}^{n-1}\sum_{s=t+1}^{n}a_{st}a_{ss}v_{t}\right)^{2}\leq CL^{-4}\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}a_{st}a_{ss}\right)^{2},$ (A.6)
+where the factor in parentheses on the RHS of (A.6) is
+$\sum_{s,r=t+1}^{n}b_{s}^{\prime}\mathscr{S}b_{s}b_{s}^{\prime}\mathscr{S}b_{t}b_{r}^{\prime}\mathscr{S}b_{r}b_{r}^{\prime}\mathscr{S}b_{t}\leq\sum_{s,r=t+1}^{n}\left|b_{s}^{\prime}\mathscr{S}b_{s}b_{r}^{\prime}\mathscr{S}b_{r}\right|\left|b_{s}^{\prime}\mathscr{S}b_{t}\right|\left|b_{r}^{\prime}\mathscr{S}b_{t}\right|$
+$\leq$ $C\sum_{s,r=t+1}^{n}\sum_{i,j,k,l=1}^{n}\left|b_{is}\right|\left|s_{ij}\right|\left|b_{jr}\right|\left|b_{ks}\right|\left|s_{lk}\right|\left|b_{kr}\right|\left|b_{s}^{\prime}\mathscr{S}b_{t}\right|\left|b_{r}^{\prime}\mathscr{S}b_{t}\right|$
+$\leq$ $C\left(\sup_{i,j}\left|s_{ij}\right|\right)^{2}\left(\sup_{s\geq 1}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{4}\sum_{s,r=t+1}^{n}\left|b_{s}^{\prime}\mathscr{S}b_{t}\right|\left|b_{r}^{\prime}\mathscr{S}b_{t}\right|$
+$=$ $O_{p}\left(J^{2}n^{-2}\left(\sum_{s=t+1}^{n}\left|b_{t}^{\prime}\mathscr{S}b_{s}\right|\right)^{2}\right)=O_{p}\left(J^{2}n^{-2}\left(\sum_{s=t+1}^{n}\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)^{2}\right),$
+where we used Assumption A.
+Now, by Assumption A, $\sum_{s=t+1}^{n}\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|=O_{p}\left(\sup_{i,j}\left|s_{ij}\right|\sup_{t}\sum_{i=1}^{n}\left|b_{it}\right|\sum_{j=1}^{n}\sum_{s=t+1}^{n}\left|b_{js}\right|\right)=O_{p}\left(J\sup_{t}\sum_{i=1}^{n}\left|b_{it}\right|\right),$ and thus (A.6) is $O_{p}\left(L^{-4}J^{4}n^{-2}\sup_{t}\left(\sum_{i=1}^{n}\left|b_{it}\right|\right)\left(\sum_{i=1}^{n}\left(\sum_{t=1}^{n-1}\left|b_{it}\right|\right)\right)\right)=O_{p}\left(L^{-4}J^{4}n^{-1}\right)$, whence the first term in (A.5) is $O_{p}\left(J^{2}n^{-1}\right)$, which is negligible.
+Again transforming the range of summation and using the elementary inequality $|a+b|^{2}\leq C\left(a^{2}+b^{2}\right)$, we can bound the square of the second term in (A.5) by
+$C\left(\sum_{t=1}^{n-1}\sum_{s=t+1}^{n}a_{st}^{2}\left(v_{t}^{2}-1\right)\right)^{2}+C\left(2\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\sum_{s=t+1}^{n}a_{st}a_{sr}v_{t}v_{r}\right)^{2}.$ (A.7)
+The expectations of the two terms in (A.7) are bounded by a constant times $e_{1}$ and $e_{2}$, respectively, where $e_{1}=\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}a_{st}^{2}\right)^{2},e_{2}=\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\left(\sum_{s=t+1}^{n}a_{st}a_{sr}\right)^{2}.$ Thus (A.7) is $O_{p}\left(e_{1}+e_{2}\right)$.
+Using Assumption A and elementary inequalities, $e_{2}$ is bounded by
+$\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\sum_{s=t+1}^{n}\sum_{u=t+1}^{n}b_{s}^{\prime}\mathscr{S}b_{t}b_{s}^{\prime}\mathscr{S}b_{r}b_{u}^{\prime}\mathscr{S}b_{t}b_{u}^{\prime}\mathscr{S}b_{r}$
+$=$ $O_{p}\left(L^{-4}\sum_{s,r,t,u=1}^{n}\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{ju}\right|\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right$.
+$\times\left.\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{ju}\right|\right)$
+$=$ $O_{p}\left(L^{-4}pn^{-1}\sum_{s,r,t=1}^{n}\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\sum_{u=1}^{n}\left|b_{ju}\right|\right)\right$.
+$\times$ $\left.\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i=1}^{n}\left|b_{it}\right|\sup_{u}\sum_{j=1}^{n}\left|b_{ju}\right|\right)$
+$=$ $O_{p}\left(L^{-4}J^{2}n^{-2}\sum_{s,r=1}^{n}\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\sum_{i=1}^{n}\left|b_{ir}\right|\sum_{j=1}^{n}\left(\sum_{u=1}^{n}\left|b_{ju}\right|\right)\right$.
+$\times$ $\left.\left(\sum_{i,j=1}^{n}\sum_{t=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\right)$
+$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sum_{i,j=1}^{n}\left(\sum_{r=1}^{n}\left|b_{ir}\right|\right)\left|s_{ij}\right|\left(\sum_{s=1}^{n}\left|b_{js}\right|\right)\left(\sup_{j}\sum_{i=1}^{n}\left|s_{ij}\right|\right)\sum_{j=1}^{n}\left|b_{js}\right|\right)$
+$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sup_{k}\sum_{i,j=1}^{n}\left|s_{ij}\right|\sum_{i=1}^{n}\left|s_{ik}\right|\right)=O_{p}\left(L^{-4}J^{2}n^{-1}\sup_{k}\sum_{i,j,\ell=1}^{n}\left|s_{ij}\right|\left|s_{\ell k}\right|\right)$
+$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sup_{k}\sum_{i,j,\ell=1}^{n}\left(s_{ij}^{2}+s_{\ell k}^{2}\right)\right)=O_{p}\left(L^{-4}J^{2}n^{-1}\sum_{i,j,\ell=1}^{n}\left(s_{ij}^{2}+s_{\ell j}^{2}\right)\right)$
+$=$ $O_{p}\left(L^{-4}J^{2}n^{-1}\sum_{i,j=1}^{n}s_{ij}^{2}\right)=O_{p}\left(L^{-4}J^{2}\sup_{j}\sum_{i=1}^{n}s_{ij}^{2}\right)=O_{p}\left(Jn^{-1}\right),$
+Similary $e_{1}$ is
+$O_{p}\left(L^{-4}\sum_{t=1}^{n-1}\left\{\sum_{s=t+1}^{n}\left(\sum_{i,j=1}^{n}\left|s_{ij}\right|\left|b_{jt}\right|\left|b_{is}\right|\right)^{2}\right\}^{2}\right)$
+$=$ $O_{p}\left(L^{-4}\left(\sup_{i,j}\left|s_{ij}\right|\right)^{4}\sum_{t=1}^{n-1}\left\{\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{2}\right\}^{2}\right)$
+$=$ $O_{p}\left((LJN)^{-4}\sum_{t=1}^{n-1}\left\{\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{2}\right\}^{2}\right)$
+$=$ $O_{p}\left((LJN)^{-4}\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right)^{2}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{4}\right)$
+$=$ $O_{p}\left((LJN)^{-4}\left(\sum_{t=1}^{n-1}\sum_{j=1}^{n}\left|b_{jt}\right|\right)\left(\sum_{s=t+1}^{n}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right$.
+$\times\left.\sup_{s}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\sup_{t}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{3}\right)$
+$=$ $O_{p}\left((L^{-4}J^{4}n^{-1}\right)=O_{p}\left(J^{2}n^{-1}\right).$
 Appendix B Proofs of Theorems
@@
 Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$.
-Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$.
+Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}\\ P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$.
 Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$.
@@
 ∎
-Proof of Theorem 2.2.
+Proof of Theorem $2.$
 This is a special case of Theorem 3.2 and therefore omitted. ∎
@@
 $\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{2}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{2}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.2)
-where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, ^$\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and ^$\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$.
+where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, $\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and $\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$.
 The remainder of the proof is similar to the proof of Theorem 5.1.
@@
 $\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{3}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{3}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.3)
-where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, ^$\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$ with $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and ^$\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$.
+where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, $\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$, $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and $\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$.
 The remainder of the proof is similar to the proof of Theorem 5.1.
@@
 $\begin{split}\mathrm{E}\left\|n^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta})\right\|^{2}&\leq n^{-2}\left\|K_{3}\right\|^{2}\mathrm{E}\left\|r_{\lambda}+r_{\delta}\right\|^{2}\\ &\leq Cn^{-1}\overline{\alpha}(n^{-1}K_{3}^{\prime}K_{3})\left(\mathrm{E}\left\|r_{\lambda}\right\|^{2}+\mathrm{E}\left\|r_{\delta}\right\|^{2}\right)=o(1),\end{split}$ (B.5)
-because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17.
-Next,
+because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17. Next,
 $\mathrm{E}\left\|n^{-1}K_{3}^{\prime}\epsilon\right\|^{2}$ $\leq$ $n^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}\mathrm{E}\epsilon_{i}^{2}+n^{-2}\sum_{i\neq j}\left\|k_{i}\right\|\left\|k_{j}\right\|\left|\mathrm{E}\epsilon_{i}\epsilon_{j}\right|$ (B.6)
@@
 $\left\|\hat{\gamma}-\gamma\right\|=O_{p}(\sqrt{d_{\gamma}/n}).$ (B.7)
-Now, observe that
-$\hat{\Xi}_{r,s}-\Xi_{r,s}=a_{1,rs}+a_{2,rs}+a_{3,rs},$ (B.8)
-where
+Now, observe that $\hat{\Xi}_{r,s}-\Xi_{r,s}=a_{1,rs}+a_{2,rs}+a_{3,rs},$ where
 $a_{1,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}(\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j})\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),$ (B.8)
@@
 where recall that $\sigma_{ij}$ is a typical element of $\Sigma$.
-We will establish $a_{l,rs}=o_{p}(1),l=1,2,3$. To begin, write $r_{i}=r_{i\lambda}+r_{i\delta}$, $g_{i}^{\prime}$ for the $i$-th row of $G$ and note that
+We will show $a_{l,rs}=o_{p}(1),l=1,2,3$.
+To begin, write $r_{i}=r_{i\lambda}+r_{i\delta}$, $g_{i}^{\prime}$ for the $i$-th row of $G$ and note that
 $\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j}=g_{i}^{\prime}(\hat{\gamma}-\gamma)g_{j}^{\prime}(\hat{\gamma}-\gamma)+g_{i}^{\prime}(\hat{\gamma}-\gamma)r_{j}+r_{i}r_{j}+g_{j}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{i}+g_{i}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{j}+\epsilon_{i}r_{j}+\epsilon_{j}r_{i}.$ (B.11)
 Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$.
-Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, also using (B.7),
-$a_{11,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|^{2}\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}^{2}}{n}\right),$
-$a_{12,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}}{n}\right),$
-$a_{13,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|r_{i}\right|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{1}{n}\right),$
-$a_{14,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
-$a_{15,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
-$a_{16,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{i}\right|\left|r_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
-$a_{17,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{j}\right|\left|r_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
+Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, writing $\pi_{M}=1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})$ and also using (B.7),
+$a_{11,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{i}\right\|\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|^{2}\leq C\ell\cdot O_{p}\left({d_{\gamma}^{2}}/{n}\right),$
+$a_{12,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left({d_{\gamma}}/{n}\right),$
+$a_{13,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left|r_{i}\right|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left({1}/{n}\right),$
+$a_{14,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({d_{\gamma}}/{n^{1/2-1/q}}\right),$
+$a_{15,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({d_{\gamma}}/{n^{1/2-1/q}}\right),$
+$a_{16,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left|\epsilon_{i}\right|\left|r_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({1}/{n^{1/2-1/q}}\right),$
+$a_{17,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\pi_{M}\left|\epsilon_{j}\right|\left|r_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left({1}/{n^{1/2-1/q}}\right),$
 where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007).
@@
 Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$.
-Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big{\|}r_{\lambda\delta}\big{\|}^{2}=o(1)$ by Assumption 16.
+Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big\|r_{\lambda\delta}\big\|^{2}=o_{p}(1)$ by Assumption 16.
 Under $H_{0}:\vartheta=0$,
-$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{4}v+A_{11}+A_{12}+2A_{21}+2A_{22})\end{split}$ (B.13)
-where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, ^$\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$ with $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and ^$\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$.
+$\begin{split}n\hat{\vartheta}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}=(v^{\prime}\widehat{\mathcal{M}}_{4}v+A_{11}+A_{12}+2A_{21}+2A_{22})\end{split}$ (B.12)
+where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, $\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$, $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and $\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$.
 Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big\|\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17.
-Thus, $\big{\|}\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big{\|}\leq\big{\|}\widehat{\mathcal{U}}_{3}^{-1}\big{\|}\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\big{\|}\mathcal{U}_{3}^{-1}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$.
-Similarly, we have ^$\big{\|}\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big{\|}=o_{p}(1)$, and ^$\big{\|}\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big{\|}=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
-Thus, we have
+Thus, $\big\|\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big\|\leq\big\|\widehat{\mathcal{U}}_{3}^{-1}\big\|\big\|\mathcal{U}_{3}-\widehat{\mathcal{U}}_{3}\big\|\big\|\mathcal{U}_{3}^{-1}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$.
+Similarly, we have $\big\|\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big\|=o_{p}(1)$, and $\big\|\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big\|=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
+Since $\mathrm{E}(v_{i}^{2})=1$, we have
 $\frac{1}{\sqrt{d_{\vartheta}}}\big|v^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v\big|\leq Cr_{n}\cdot O_{p}\left(d_{\vartheta}^{-1/2}\right)=o_{p}(1)$ (B.13)
-since $\mathrm{E}(v_{i}^{2})=1$.
-Similar to (B.14) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\alpha}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\alpha}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\alpha}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\alpha}}}|A_{22}|=o_{p}(1)$.
+Similar to (B.13) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\vartheta}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\vartheta}}}|A_{22}|=o_{p}(1)$.
 Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible.
+∎
 Proof of Theorem 5.2.
 (i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1.
-We now proceed as in the proof Theorem 4.3 of Gupta and Qu (2024) and note that the entries of $\mathscr{S}$ are,
-$s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ (B.15)
-where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so
-$\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.16)
-uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above.
-Similarly, we also observe that
-$\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.17)
-uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 17. The conclusion now follows by Theorem A.1.
-(ii) This proof is similar to that of Theorem 4.4 in Gupta and Qu (2024).
-(iii) We prove all local alternatives here.
-We define $H_{\ell i}\equiv H_{\ell i,n}$, with the corresponding statistic being $\mathbb{W}_{i}^{*}$, for $i=1,2,3,4$.
+We now note that the entries of $\mathscr{S}$ are $s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so $\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above.
+Similarly, we also observe that $\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i$.
+Thus Assumption A(b) is satisfied.
+Assumption A is satisfied to due to Assumption 17.
+The conclusion now follows by Theorem A.1.
+(ii) By Lemma 5.1, we have
+$\left\|\hat{\vartheta}-\vartheta\right\|=O_{p}\left(\sqrt{d_{\vartheta}/n}\right).$ (B.14)
+From Assumptions 15 and 17, we have $\left\|\widehat{\mathscr{D}}_{4}-\mathscr{D}_{4}\right\|=o_{p}(1)$ and
+$\left\|\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1}\right\|\leq\left\|\widehat{\mathscr{D}}_{4}^{-1}\right\|\left\|\mathscr{D}_{4}-\widehat{\mathscr{D}}_{4}\right\|\left\|\mathscr{D}_{4}^{-1}\right\|=o_{p}(1).$ (B.15)
+Write $\hat{\vartheta}=\vartheta+\Delta_{\vartheta}$ from (B.14) where $\Delta_{\vartheta}=O_{p}\left(\sqrt{d_{\vartheta}/n}\right)$.
+Then $n\hat{\vartheta}^{\prime}\hat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}=n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\vartheta+2n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}+n\Delta_{\vartheta}^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}+nR_{n},$ where $\left|R_{n}\right|=\left|\hat{\vartheta}^{\prime}\hat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}-\hat{\vartheta}^{\prime}\mathscr{D}_{4}^{-1}\hat{\vartheta}\right|=o_{p}(1)\cdot\left\|\hat{\vartheta}\right\|^{2}=o_{p}\left(\|\vartheta\|^{2}+d_{\vartheta}/n\right)=o_{p}\left(\|\vartheta\|^{2}\right)$ by (B.15).
+Using (5.8), (B.14) and bounded eigenvalues in (B.15),
+$\begin{split}&\frac{1}{\sqrt{2d_{\vartheta}}}2n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}=O\left(\frac{n}{\sqrt{d_{\vartheta}}}\right)O\left(\|\vartheta\|\right)O_{p}(1)O_{p}\left(\sqrt{d_{\vartheta}/n}\right)=o_{p}\left(\frac{n\|\vartheta\|}{\sqrt{d_{\vartheta}}}\right)\\ &\frac{1}{\sqrt{2d_{\vartheta}}}n\Delta_{\vartheta}^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}=O\left(\frac{n}{\sqrt{d_{\vartheta}}}\right)O_{p}(1)O_{p}\left(\frac{d_{\vartheta}}{n}\right)=o_{p}\left(\frac{n}{\sqrt{d_{\vartheta}}}\right)\\ &\frac{1}{\sqrt{2d_{\vartheta}}}nR_{n}=o_{p}\left(\frac{n\|\vartheta\|^{2}}{\sqrt{d_{\vartheta}}}\right).\end{split}$ (B.16)
+Note that for the sparse case, $\|\vartheta\|=O(1)$, and for the dense case, $\|\vartheta\|_{\infty}=O(1)$ and $\|\vartheta\|=\sqrt{d_{\vartheta}}$.
+Thus, in general, $\mathbb{W}_{4}=\frac{n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\vartheta-d_{\vartheta}}{\sqrt{2d_{\vartheta}}}+o_{p}\Big(\frac{n\|\vartheta\|^{2}}{\sqrt{d_{\vartheta}}}\Big)$.
+For any nonstochastic sequence $\{C_{n}\},C_{n}=o\left(\frac{n\|\vartheta\|^{2}}{\sqrt{d_{\vartheta}}}\right),\lim_{n\to\infty}P\left(\left|\mathbb{W}_{4}\right|>C_{n}\right)=1$, so that consistency follows.
+(iii) We cover all local alternatives here.
+We define $H_{\ell i}\equiv H_{\ell i,n}$, with corresponding statistic $\mathbb{W}_{i}^{*}$, for $i=1,2,3,4$.
 For $\mathbb{W}_{1}^{*}$, we consider $\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$ with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix.
-Under $H_{\ell 1}$, we have
-$\mathbb{W}_{1}^{*}=\mathbb{W}_{1}+\frac{1}{\sqrt{2d_{\alpha}}}d_{\alpha}^{\frac{1}{2}}\mathcal{A}_{1}(\frac{1}{n}\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{-1}+o_{p}(1),$
-where $\mathcal{A}_{1}=\frac{1}{n}\nu_{1n}^{\prime}\mathscr{D}_{1}^{-1}\nu_{1n}$.
+Under $H_{\ell 1}$, we have $\mathbb{W}_{1}^{*}=\mathbb{W}_{1}+\frac{1}{\sqrt{2d_{\alpha}}}d_{\alpha}^{\frac{1}{2}}\mathcal{A}_{1}\left(\frac{1}{n}\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n}\right)^{-1}+o_{p}(1),$ where $\mathcal{A}_{1}=\frac{1}{n}\nu_{1n}^{\prime}\mathscr{D}_{1}^{-1}\nu_{1n}$.
 We choose $\Gamma_{1n}=\mathscr{D}_{1}^{-1}$ and then finish the proof.
-For $\mathbb{W}_{4}^{*}$, we consider $\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{\frac{1}{2}}$.
-Under $H_{\ell 4}$, we have
-$\mathbb{W}_{4}^{*}=\mathbb{W}_{4}+\frac{1}{\sqrt{2d_{\vartheta}}}d_{\vartheta}^{\frac{1}{2}}\mathcal{A}_{4}(\frac{1}{n}\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{-1}+o_{p}(1),$
-where $\mathcal{A}_{4}=\frac{1}{n}\nu_{4n}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\nu_{4n}$, and $\big{\|}\frac{1}{n}\nu_{4n}^{\prime}(\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1})\nu_{4n}\big{\|}=o_{p}(1)$ by (5.7).
+For $\mathbb{W}_{4}^{*}$, we consider $\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/\left(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n}\right)^{\frac{1}{2}}$.
+Under $H_{\ell 4}$, we have $\mathbb{W}_{4}^{*}=\mathbb{W}_{4}+\frac{1}{\sqrt{2d_{\vartheta}}}d_{\vartheta}^{\frac{1}{2}}\mathcal{A}_{4}\left(\frac{1}{n}\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n}\right)^{-1}+o_{p}(1),$ where $\mathcal{A}_{4}=\frac{1}{n}\nu_{4n}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\nu_{4n}$, and $\left\|\frac{1}{n}\nu_{4n}^{\prime}(\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1})\nu_{4n}\right\|=o_{p}(1)$ by (5.7).
 We choose $\Gamma_{4n}=\widehat{\mathscr{D}}_{4}^{-1}$ and then finish the proof.
@@
 Appendix C Empirical Results
-Table 1: Data descriptions in Section 6.1, logarithm form.
-Variables Descriptions Mean Min Max Sd
-$y$ Value added 11.539 3.611 19.216 1.847
-$p_{1}$ Capital assets 10.280 1.609 19.160 1.980
-$p_{2}$ Average number of employ 5.647 2.021 11.865 1.130
-$z$ Management expense 7.707 2.303 16.139 1.853
-Table 2: Results of baseline model according to (6.1) in Section 6.1.
-                 $\delta_{0}$                  3.313***
-                 $\delta_{1}$                  0.554***
-                 $\delta_{2}$                  0.402***
-                 $\beta_{1}$                  0.224
-                 $\beta_{2}$                  0.018
-                 $H_{0}:\delta_{1}+\delta_{2}-1=0$
-                 t-test                  -0.238
-                 p-value                  0.812
-                 $n$                  7355
-                 The standard errors of the estimators
-                 are reported in parentheses.
-                  $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 3: Estimation and testing results in Section 6.1 using polynomial series with $h=2$.
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
+Table 1: Estimation and testing results in Section 6.1.
+Panel A: polynomial series, $h=2$
+Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.189* 0.130** 0.224*** 0.256*** 0.451*** 0.457*** 0.499***
 (0.112) (0.062) (0.014) (0.081) (0.157) (0.042) (0.028)
-$\tau_{2}/\mu_{2}$0.095 0.405*** 0.484***
+$\tau_{2}/\mu_{2}$ -0.095 0.405*** 0.484***
 (0.126) (0.047) (0.030)
-$\beta_{0}$ 1.073** 0.167* 0.153** 0.167*** 0.153*** -0.485*** -0.093*** 0.077***
-(0.514) (0.100) (0.076) (0.016) (0.020) (0.013) (0.004) (0.003)
-$\alpha_{11}$ 1.254** -0.036*** -0.032*** -0.036 -0.032 -0.023 0.005 0.004
-(0.551) (0.007) (0.004) (0.027) (0.035) (0.975) (0.033) (0.026)
-$\alpha_{12}$ 0.375 -0.015 0.017 -0.015 0.017 -0.032 0.015*** 0.027***
-(3.048) (0.035) (0.019) (0.010) (0.036) (0.050) (0.002) (0.002)
-$\alpha_{21}$0.194 -0.004 -0.004 -0.004 -0.004 -0.002 0.000 0.003
-(0.786) (0.020) (0.033) (0.005) (0.012) (0.111) (0.001) (0.002)
-$\alpha_{22}$0.023 0.092*** 0.088*** 0.092*** 0.088*** 0.032*** -0.017*** 0.012***
-(0.077) (0.013) (0.013) (0.004) (0.029) (0.008) (0.000) (0.000)
-$\alpha_{31}$0.046 0.031 0.019 0.031 0.019* 0.011 0.001 0.003
-(0.891) (0.017) (0.010) (0.022) (0.011) (0.136) (0.018) (0.014)
-$\alpha_{32}$0.062 0.004 0.005 0.004 0.005 0.059 -0.033 0.026
-(0.092) (0.023) (0.023) (0.001) (0.004) (0.060) (0.031) (0.051)
 $\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
 $H_{0}:\alpha_{3}=0$ 1.593 3.311 3.280 1.419 4.020 0.226 1.716 2.589
-$\mathscr{T}_{n}$ 0.451 0.191 0.194 0.492 0.134 0.893 0.424 0.274
-$\mathscr{T}_{n}^{*}$ 0.767 0.213 0.167 0.413 0.167 0.867 0.433 0.213
-$H_{0}:\mu=0$ 68.365 92.106
-$\mathscr{T}_{n}$ $<$0.001 $<$0.001
-$\mathscr{T}_{n}^{*}$ 0.007 0.007
-$n$ 7355
-The standard errors of estimators are reported in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 4: Estimation and testing results in Section 6.1 using trigonometric series with $h=2$.
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
+$H_{0}:\mu=0$ 68.365*** 92.106***
+Panel B: trigonometric series, $h=2$
+Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.065 0.162* 0.065** 0.162*** 0.582*** 0.607*** 0.297***
 (0.052) (0.087) (0.033) (0.014) (0.098) (0.166) (0.088)
-$\tau_{2}/\mu_{2}$0.276*** 0.147** 0.494***
+$\tau_{2}/\mu_{2}$ -0.276*** 0.147** 0.494***
 (0.041) (0.074) (0.088)
-$\beta_{0}$ 1.111*** -0.381*** -0.382** -0.381*** -0.382*** -0.187*** 0.747*** 1.076***
-(0.020) (0.136) (0.162) (0.042) (0.021) (0.012) (0.245) (0.380)
-$\alpha_{11}$ 0.949*** 0.167*** 0.160*** 0.167*** 0.160*** 0.125*** -0.253*** -0.024***
-(0.150) (0.023) (0.025) (0.014) (0.012) (0.016) (0.025) (0.004)
-$\alpha_{12}$ 0.861*** -0.049 -0.028** -0.049 -0.028 -0.124*** 0.146*** 0.034***
-(0.151) (0.033) (0.012) (0.035) (0.032) (0.008) (0.040) (0.007)
-$\alpha_{21}$ 0.001 0.0096 0.094 0.096*** 0.094*** 0.014*** -0.091*** 0.021***
-(0.007) (0.112) (0.124) (0.003) (0.003) (0.002) (0.019) (0.005)
-$\alpha_{22}$ 0.002 0.026 0.024 0.026*** 0.024*** -0.018*** 0.020*** 0.009***
-(0.007) (0.069) (0.043) (0.004) (0.003) (0.002) (0.005) (0.003)
-$\alpha_{31}$ 0.017 0.085 0.068 0.085 0.068 0.001 -0.020 0.019
-(0.018) (0.246) (0.191) (0.105) (0.064) (0.001) (0.017) (0.012)
-$\alpha_{32}$ 0.009 0.033 0.025 0.033 0.025 -0.003 0.047 0.021
-(0.022) (0.090) (0.084) (0.002) (0.002) (0.003) (0.007) (0.006)
 $\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
 $H_{0}:\alpha_{3}=0$ 0.777 1.178 0.877 1.086 0.943 1.575 1.707 2.169
-$\mathscr{T}_{n}$ 0.678 0.555 0.645 0.581 0.624 0.455 0.426 0.338
-$\mathscr{T}_{n}^{*}$ 0.660 0.533 0.680 0.553 0.600 0.467 0.267 0.333
-$H_{0}:\mu=0$ 321.305 406.511
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.013 0.027
-$n$ 7355
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 5: Estimation and testing results in Section 6.1 using polynomial series with $h=4$.
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
+$H_{0}:\mu=0$ 321.305*** 406.511***
+Panel C: polynomial series, $h=4$
+Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.330*** 0.265** 0.330*** 0.065*** 0.444*** 0.236*** 0.313***
 (0.095) (0.134) (0.019) (0.024) (0.135) (0.043) (0.0969)
-$\tau_{2}/\mu_{2}$0.013 0.219*** 0.330***
+$\tau_{2}/\mu_{2}$ -0.013 0.219*** 0.330***
 (0.045) (0.052) (0.064)
@@
 (0.012) (0.082) (0.126)
-$\tau_{4}/\mu_{4}$0.088 0.189* 0.096*
+$\tau_{4}/\mu_{4}$ -0.088 0.189* 0.096*
 (0.068) (0.110) (0.051)
-$\beta_{0}$ 1.866 0.040*** 0.005*** 0.040*** 0.005 -0.036 0.075 -0.278***
-(4.659) (0.011) (0.001) (0.011) (0.023) (0.159) (0.221) (0.062)
-$\alpha_{11}$ 1.471 0.002*** 0.000*** 0.002 0.000 0.000 0.003 -0.006
-(4.108) (0.000) (0.000) (1.759) (2.686) (10.471) (26.990) (3.119)
-$\alpha_{12}$1.067 0.005*** 0.002*** 0.005 0.002 0.010 0.009 -0.005
-(5.407) (0.001) (0.000) (0.329) (0.588) (2.398) (5.235) (0.903)
-$\alpha_{13}$ 1.186 0.013*** 0.007*** 0.013 0.007 0.035 0.014 0.010
-(8.484) (0.001) (0.002) (0.028) (0.053) (0.232) (0.466) (0.100)
-$\alpha_{14}$0.020 0.121*** -0.004 0.121*** -0.004*** -0.174*** 0.180*** -0.046***
-(0.288) (0.021) (0.042) (0.001) (0.002) (0.008) (0.016) (0.004)
-$\alpha_{21}$0.032 0.015*** 0.003*** 0.015 0.003 0.000 0.006 -0.004
-(0.676) (0.004) (0.001) (0.180) (0.462) (0.861) (3.899) (0.134)
-$\alpha_{22}$0.029 0.041*** 0.020*** 0.041 0.020 0.004 0.013 0.022
-(2.100) (0.008) (0.006) (0.031) (0.100) (0.190) (0.642) (0.071)
-$\alpha_{23}$ 0.003 -0.010 0.092*** -0.010*** 0.092*** -0.011 -0.018 0.019*
-(0.213) (0.019) (0.026) (0.002) (0.009) (0.017) (0.046) (0.010)
-$\alpha_{24}$ $0$0.005*** -0.003** -0.005*** -0.003*** 0.005*** -0.001 -0.001***
-(0.007) (0.000) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000)
-$\alpha_{31}$ 0.005 -0.005 -0.002*** -0.005 -0.002 -0.001 -0.002 0.006
-(0.012) (0.061) (0.001) (0.344) (0.556) (1.561) (3.200) (0.827)
-$\alpha_{32}$0.009 -0.008 -0.009 -0.008 -0.009 -0.009 -0.002 0.004
-(0.042) (0.042) (0.053) (0.068) (0.124) (0.428) (0.543) (0.297)
-$\alpha_{33}$ 0.007 0.059 -0.041 0.059 -0.041 -0.032 0.029 0.031
-(0.032) (0.096) (0.113) (0.046) (0.033) (0.041) (0.040) (0.031)
-$\alpha_{34}$0.002 0.001 0.016*** 0.001 0.016 0.008 -0.004 -0.003
-(0.026) (0.073) (0.157) (0.006) (0.062) (0.011) (0.083) (0.044)
 $\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
 $H_{0}:\alpha_{3}=0$ 0.075 1.707 1.464 2.401 1.808 0.030 0.098 0.385
-$\mathscr{T}_{n}$ 0.963 0.426 0.481 0.301 0.405 0.985 0.952 0.825
-$\mathscr{T}_{n}^{*}$ 0.947 0.440 0.500 0.287 0.467 0.933 0.887 0.793
-$H_{0}:\mu=0$ 101.231 122.398
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.007 0.007
-$n$ 7355
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 6: Estimation and testing results in Section 6.1 using trigonometric series with $h=4$.
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
+$H_{0}:\mu=0$ 101.231*** 122.398***
+Panel D: trigonometric series, $h=4$
+Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.224*** 0.256*** 0.224*** 0.256*** 0.196** 0.123* 0.136***
@@
 (0.018) (0.145) (0.140)
-$\tau_{4}/\mu_{4}$0.011 0.107* 0.138*
+$\tau_{4}/\mu_{4}$ -0.011 0.107* 0.138*
 (0.124) 0.026** 0.035***
-$\beta_{0}$ 1.108*** 0.060** 0.004** 0.060*** 0.004*** -0.067*** (0.008) (0.009)
-(0.021) (0.031) (0.002) (0.009) (0.001) (0.025) (0.055) (0.077)
-$\alpha_{11}$ 0.995*** 0.167*** 0.153*** 0.167*** 0.153*** -0.415*** 0.143*** 0.150***
-(0.162) (0.010) (0.011) (0.016) (0.020) (0.014) (0.016) (0.014)
-$\alpha_{12}$ 0.904*** -0.036*** -0.032*** -0.036 -0.032 0.735*** -0.032 -0.022
-(0.164) (0.005) (0.006) (0.027) (0.035) (0.019) (0.023) (0.021)
-$\alpha_{13}$ 0.084 -0.015** 0.017*** -0.015 0.017 -0.198*** -0.003 -0.026***
-(0.158) (0.007) (0.005) (0.010) (0.036) (0.006) (0.008) (0.006)
-$\alpha_{14}$0.251 -0.004*** -0.004 -0.004 -0.004 -0.160*** -0.002 -0.019***
-(0.160) (0.001) (0.003) (0.005) (0.012) (0.007) (0.005) (0.004)
-$\alpha_{21}$ 0.001 0.092*** 0.088*** 0.092*** 0.088*** -0.006*** 0.037*** 0.037***
-(0.007) (0.015) (0.016) (0.004) (0.029) (0.001) (0.003) (0.003)
-$\alpha_{22}$ 0.001 0.031*** 0.019*** 0.031*** 0.019*** -0.037*** 0.013*** 0.014***
-(0.008) (0.012) (0.003) (0.002) (0.003) (0.003) (0.002) (0.002)
-$\alpha_{23}$ 0.000 0.004 0.035 0.004*** 0.035*** 0.014*** 0.007*** -0.003***
-(0.007) (0.021) (0.022) (0.001) (0.004) (0.001) (0.001) (0.001)
-$\alpha_{24}$ 0.005 0.003 0.006 0.003*** 0.006*** 0.003*** 0.002*** 0.002***
-(0.008) (0.004) (0.008) (0.000) (0.001) (0.001) (0.000) (0.000)
-$\alpha_{31}$ 0.018 0.087 0.068 0.087 0.068 -0.027 0.041 0.041
-(1.320) (1.065) (1.012) (0.127) (0.126) (0.863) (0.079) (0.166)
-$\alpha_{32}$ 0.020 0.029 0.025 0.029 0.025 -0.081 0.013 0.006
-(0.685) (2.392) (2.003) (0.268) (0.212) (0.062) (0.091) (0.099)
-$\alpha_{33}$ 0.003 0.028 0.005 0.028 0.005 0.019 0.010 0.009
-(0.331) (5.691) (4.001) (0.316) (0.305) (0.891) (1.063) (0.085)
-$\alpha_{34}$ 0.005 -0.022 -0.020 -0.022 -0.020 0.019 -0.011 -0.013
-(0.369) (0.355) (0.093) (0.287) (0.365) (1.605) (0.062) (0.076)
 $\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
 $H_{0}:\alpha_{3}=0$ 0.290 1.848 1.497 0.038 1.443 0.034 0.375 1.022
-$\mathscr{T}_{n}$ 0.865 0.397 0.473 0.981 0.486 0.983 0.829 0.600
-$\mathscr{T}_{n}^{*}$ 0.873 0.393 0.453 0.953 0.467 0.927 0.833 0.580
-$H_{0}:\mu=0$ 86.301 90.233
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.027 0.007
-$n$ 7355
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+$H_{0}:\mu=0$ 86.301*** 90.233***
+Standard errors in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
 Table 2: Comparison of parametric method and semiparametric method in Section 6.1.
-No $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
+Without $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
 parametric $\hat{\delta}_{1}$ 0.5536 0.5438 0.5330 0.5989 0.5996 0.5960
 $\hat{\delta}_{2}$ 0.4022 0.3794 0.3725 0.3145 0.3226 0.3099
-$\hat{\delta}_{1}$+$\hat{\delta}_{2}$ 0.9558 0.9232 0.9055 0.9134 0.9222 0.9060
+$\hat{\delta}_{1}$$\hat{\delta}_{2}$ 0.9558 0.9232 0.9055 0.9134 0.9222 0.9060
 polynomial, $h=2$ $\hat{\delta}_{1}$ 0.6228 0.6488 0.6179 0.6263 0.6051 0.6108
@@
 $\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9913 0.9598 0.9816 0.9195 0.9938 0.9768
-$n$ 7355
-The result in the first column does not consider any spatial weight matrix.
-Figure 1: Comparison of parametric method and semiparametric method in Section 6.1, capital.
-Figure 2: Comparison of parametric method and semiparametric method in Section 6.1, labor.
-Figure 3: Comparison of parametric method and semiparametric method in Section 6.1, CRS.
-Table 8: Data descriptions in Section 6.2.
-Variables Descriptions Mean Min Mean SD
-$y$ log(MEDV) 3.035 1.609 3.912 0.408
-$x_{1}$ log(LSTAT) 2.371 0.548 3.637 0.601
-$x_{2}$ log(RAD) 1.868 0 3.178 0.875
-$p_{1}$ log(CRIM) -0.780 -5.064 4.488 2.162
-$p_{2}$ log(RM) 1.832 1.270 2.172 0.112
-$p_{3}$ log(TAX) 5.931 5.231 6.567 0.396
-$z$ log(DIS) 1.188 0.122 2.495 0.540
-Table 9: Estimation and testing results in Section 6.2 using polynomial series with $h=2$.
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
+(a) Capital
+(b) Labor
+(c) CRS
+Figure 1: Cobb-Douglas versus semiparametric production.
+Table 3: Estimation and testing results in Section 6.2
+Panel A: polynomial series $h=2$
+Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.238*** 0.219*** 0.238*** 0.219*** 0.214** 0.512*** 0.468***
@@
 (0.129) (0.049) (0.057)
-$\beta_{0}$ 4.081*** 3.987*** 4.081*** 3.987*** -0.003 0.132*** 0.022*
-(0.267) (0.272) (0.147) (0.156) (0.006) (0.018) (0.012)
-$\beta_{1}$0.432*** -0.432*** -0.432*** -0.432*** -0.143* -0.053*** -0.162
-(0.025) (0.024) (0.027) (0.025) (0.075) (0.004) (0.175)
-$\beta_{2}$ 0.166*** 0.154*** 0.166*** 0.154*** 0.086* 0.025*** 0.044**
-(0.024) (0.023) (0.018) (0.017) (0.049) (0.001) (0.020)
-$\alpha_{11}$0.044*** -0.041*** -0.044*** -0.041* -0.056* -0.033** -0.017***
-(0.006) (0.006) (0.006) (0.024) (0.030) (0.001) (0.004)
-$\alpha_{12}$ 0.028*** 0.025*** 0.028*** 0.025*** 0.043* 0.022*** 0.014***
-(0.004) (0.004) (0.005) (0.005) (0.022) (0.001) (0.003)
-$\alpha_{21}$ 0.042** 0.046** 0.042** 0.046*** 0.115** -0.001 -0.068***
-(0.019) (0.018) (0.019) (0.006) (0.051) (0.002) (0.011)
-$\alpha_{22}$0.012 -0.015 -0.012 -0.015 0.136*** 0.008*** 0.035***
-(0.012) (0.011) (0.009) (0.022) (0.049) (0.001) (0.004)
-$\alpha_{31}$$0**$027*** -$0**$027*** $0*$005*** -0.019***
-(0.007) (0.007) (0.006) (0.002) (0.030) (0.002) (0.003)
-$\alpha_{32}$ 0.076* 0.070* 0.076*** $0**$084*** 0.082*** 0.101***
-(0.040) (0.040) (0.003) (0.004) (0.015) (0.021) (0.011)
 Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 183.122 91.459 123.439
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
-$H_{0}:\mu=0$ 32.336 35.141
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.007 0.013
-$n$ 506
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 10: Estimation and testing results in Section 6.2 using trigonometric series with $h=2$.
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.254*** 0.225*** 0.254*** 0.225*** 0.259*** 0.177*** 0.121**
-(0.049) (0.044) (0.028) (0.023) (0.069) (0.027) (0.057)
-$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
-(0.149) (0.040) (0.083)
-$\beta_{0}$ 3.527*** 3.547*** 3.527*** 3.547*** 0.001 0.567*** 0.675***
-(0.236) (0.229) (0.175) (0.139) (0.009) (0.091) (0.028)
-$\beta_{1}$0.410*** -0.415*** -0.410*** -0.415*** -0.235*** -0.396*** -0.456***
-(0.024) (0.023) (0.023) (0.024) (0.064) (0.021) (0.032)
-$\beta_{2}$ 0.156*** 0.146*** 0.156*** 0.146*** 0.170*** 0.154*** 0.170***
-(0.024) (0.023) (0.017) (0.017) (0.031) (0.018) (0.020)
-$\alpha_{11}$0.056* -0.035* -0.056*** -0.035*** -0.024* -0.006** -0.007*
-(0.040) (0.021) (0.003) (0.003) (0.012) (0.003) (0.004)
-$\alpha_{12}$0.023*** -0.022*** -0.023*** -0.022*** -0.163*** -$0**$024***
-(0.005) (0.005) (0.006) (0.005) (0.027) (0.005) (0.005)
-$\alpha_{21}$ 0.054*** 0.049*** 0.054*** 0.049*** 0.241** 0.037*** 0.048***
-(0.010) (0.010) (0.008) (0.007) (0.097) (0.009) (0.011)
-$\alpha_{22}$0.016 -0.007 -0.016*** -0.007*** 0.212*** 0.106** 0.132**
-(0.018) (0.018) (0.003) (0.002) (0.026) (0.052) (0.065)
-$\alpha_{31}$0.027*** -0.023*** -0.027*** -0.023*** -0.017 -0.025*** -0.030***
-(0.004) (0.004) (0.002) (0.002) (0.036) (0.002) (0.002)
-$\alpha_{32}$ 0.010* 0.009* 0.010*** 0.009** 0.622*** 0.053*** 0.065***
-(0.006) (0.005) (0.004) (0.004) (0.015) (0.004) (0.005)
-Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 38.897 34.423 229.115 325.327 266.301 175.332 144.982
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.020 0.007 0.007 0.020 0.007 0.013 0.007
-$H_{0}:\mu=0$ 28.883 21.326
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.027 0.007
-$n$ 506
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 11: Estimation and testing results in Section 6.2 using polynomial series with $h=4$.
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
+$H_{0}:\alpha=0$ 40.041*** 33.913*** 290.165*** 33.913*** 183.122*** 91.459*** 123.439***
+$H_{0}:\mu=0$ 32.336*** 35.141***
+Panel B: polynomial series $h=4$
+Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.239*** 0.269*** 0.239** 0.269*** 0.162** 0.169** 0.169***
@@
 (0.035) (0.060) (0.053)
-$\beta_{0}$ 3.285*** 2.925*** 3.285 2.925*** -$0**$092* -0.440**
-(0.464) (0.444) (7.140) (0.347) (0.003) (0.047) (0.211)
-$\beta_{1}$0.399*** -0.386*** -0.399*** -0.386*** -0.516*** 0.021* -0.105***
-(0.026) (0.025) (0.039) (0.025) (0.092) (0.012) (0.013)
-$\beta_{2}$ 0.158*** 0.137*** 0.158*** 0.137*** 0.313*** 0.019* 0.052***
-(0.025) (0.025) (0.023) (0.018) (0.057) (0.011) (0.013)
-$\alpha_{11}$ 0.041 0.031 0.041 0.031 0.048* 0.034** 0.006**
-(0.053) (0.050) (0.067) (0.039) (0.027) (0.015) (0.003)
-$\alpha_{12}$0.224* -0.185* -0.224* -0.185** -0.120*** -0.072** -0.012*
-(0.118) (0.112) (0.135) (0.086) (0.024) (0.034) (0.006)
-$\alpha_{13}$ 0.228*** 0.193** 0.228*** 0.193*** -0.055* 0.050*** 0.080*
-(0.087) (0.083) (0.087) (0.062) (0.028) (0.016) (0.046)
-$\alpha_{14}$0.064*** -0.055*** -0.064*** -0.055*** 0.058*** -0.011** -0.018*
-(0.021) (0.020) (0.018) (0.015) (0.018) (0.005) (0.010)
-$\alpha_{21}$0.706*** -$0**$706*** -$0**$073 0.019 0.003**
-(0.240) (0.226) (0.226) (0.114) (0.073) (0.011) (0.001)
-$\alpha_{22}$ 1.574*** 1.596*** 1.574*** 1.596*** -0.014 -0.026* -0.006**
-(0.484) (0.453) (0.460) (0.229) (0.045) (0.015) (0.002)
-$\alpha_{23}$1.040*** -1.078*** -1.040*** -1.078*** -0.020 0.056*** 0.030***
-(0.312) (0.291) (0.295) (0.147) (0.016) (0.014) (0.009)
-$\alpha_{24}$ 0.217*** 0.230*** 0.217*** 0.230*** -0.057* $0**$045**
-(0.065) (0.060) (0.060) (0.030) (0.030) (0.003) (0.018)
-$\alpha_{31}$ 0.262*** 0.264*** 0.262*** 0.264*** 0.068*** -$0**$017***
-(0.084) (0.079) (0.080) (0.041) (0.015) (0.003) (0.006)
-$\alpha_{32}$0.572*** -0.577*** -0.572*** -0.577*** 0.032 0.019*** 0.035**
-(0.166) (0.155) (0.164) (0.079) (0.027) (0.007) (0.014)
-$\alpha_{33}$ 0.371*** 0.381*** 0.371*** 0.381*** -0.026 -0.089* -0.021***
-(0.106) (0.099) (0.105) (0.050) (0.035) (0.046) (0.004)
-$\alpha_{34}$0.076*** -$0**$076*** -0.080*** 0.029* 0.011* 0.038***
-(0.022) (0.020) (0.021) (0.010) (0.015) (0.006) (0.007)
 $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 33.876 20.486 392.020 928.892 601.967 394.776 352.386
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.007 0.013 0.013 0.007 0.020 0.013 0.013
-$H_{0}:\mu=0$ 56.124 66.981
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.027 0.007
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 12: Estimation and testing results in Section 6.2 using trigonometric series with $h=4$.
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
+$H_{0}:\alpha=0$ 33.876*** 20.486*** 392.020*** 928.892*** 601.967*** 394.776*** 352.386***
+$H_{0}:\mu=0$ 56.124*** 66.981***
+Panel C: trigonometric series $h=2$
+Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
 $\lambda/\tau_{1}/\mu_{1}$ 0.230*** 0.275*** 0.230*** 0.275*** 0.211*** 0.039 0.043***
@@
 (0.031) (0.190) (0.041)
-$\beta_{0}$ 3.207*** 2.762*** 3.207*** 2.762*** -0.005 0.001** 0.001***
+Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
+$H_{0}:\alpha=0$ 40.041*** 33.913*** 290.165*** 33.913*** 258.704*** 91.459*** 123.439***
+$H_{0}:\mu=0$ 43.155*** 46.397***
+Panel D: trigonometric series $h=4$
+Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
+$\lambda/\tau_{1}/\mu_{1}$ 0.254*** 0.225*** 0.254*** 0.225*** 0.259*** 0.177*** 0.121**
 (0.049) (0.044) (0.028) (0.023) (0.069) (0.027) (0.057)
-$\beta_{1}$0.392*** -0.376*** -0.392*** -0.376*** -0.296*** -0.257** -0.200***
-(0.026) (0.024) (0.031) (0.024) (0.061) (0.111) (0.025)
-$\beta_{2}$ 0.146*** 0.124*** 0.146*** 0.124*** -0.046 0.103 0.069***
-(0.025) (0.024) (0.019) (0.016) (0.036) (0.299) (0.010)
-$\alpha_{11}$0.089*** -0.082*** -0.089*** -0.082*** 0.234*** -0.074** -0.060***
-(0.027) (0.025) (0.019) (0.021) (0.078) (0.032) (0.017)
-$\alpha_{12}$ 0.201*** 0.175*** 0.201*** 0.175*** 0.355*** 0.162** 0.068*
-(0.068) (0.064) (0.046) (0.045) (0.093) (0.071) (0.038)
-$\alpha_{13}$$0**$084*** -$0**$084*** -0.315*** -0.075** -0.020
-(0.031) (0.029) (0.022) (0.019) (0.095) (0.033) (0.017)
-$\alpha_{14}$$0**$081*** -$0**$081*** 0.236*** -0.073*** -0.054***
-(0.026) (0.025) (0.019) (0.021) (0.066) (0.022) (0.017)
-$\alpha_{21}$ 0.086*** 0.104*** 0.086*** 0.104*** -0.283*** 0.064*** 0.107***
-(0.031) (0.029) (0.021) (0.016) (0.085) (0.019) (0.011)
-$\alpha_{22}$0.611*** -0.688*** -0.611*** -0.688*** -0.005 -0.394 -0.506***
-(0.179) (0.165) (0.170) (0.146) (0.062) (1.827) (0.138)
-$\alpha_{23}$ 0.354*** 0.396*** 0.354*** 0.396*** 0.305*** 0.229*** 0.278***
-(0.103) (0.095) (0.100) (0.086) (0.108) (0.071) (0.080)
-$\alpha_{24}$ 0.025 0.050 0.025 $0**$042 0.036 0.091***
-(0.036) (0.034) (0.019) (0.014) (0.033) (0.076) (0.011)
-$\alpha_{31}$0.027** -0.027** -0.027*** -0.027*** 0.128*** -0.025*** -0.035***
-(0.013) (0.012) (0.009) (0.007) (0.032) (0.006) (0.005)
-$\alpha_{32}$ 0.216*** 0.238*** 0.216*** 0.238*** -0.138 0.150* 0.217***
-(0.061) (0.056) (0.057) (0.048) (0.086) (0.088) (0.043)
-$\alpha_{33}$0.123*** -0.136*** -0.123*** -0.136*** -0.047 -0.086** -0.121***
-(0.035) (0.033) (0.033) (0.029) (0.046) (0.038) (0.025)
-$\alpha_{34}$0.004 -0.012 -0.004 -0.012** -0.014 -0.009 -0.030***
-(0.013) (0.012) (0.008) (0.005) (0.026) (0.018) (0.005)
+$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
+(0.149) (0.040) (0.083)
 Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 258.704 91.459 123.439
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
-$H_{0}:\mu=0$ 43.155 46.397
-$\mathscr{T}_{n}$ $<0.01$ $<0.01$
-$\mathscr{T}_{n}^{*}$ 0.027 0.007
-$n$ 506
-The standard errors of estimators are reported in parentheses.
-$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 13: Comparison of parametric method and semiparametric method in Section 6.2.
+$H_{0}:\alpha=0$ 38.897*** 34.423*** 229.115*** 325.327*** 266.301*** 175.332*** 144.982***
+$H_{0}:\mu=0$ 28.883*** 21.326***
+Panel E: Comparison of parametric method and semiparametric method in Section 6.2
 polynomial, $h=2$ polynomial, $h=4$ trigonometric, $h=2$ trigonometric, $h=4$
@@
 0.3136 0.3371 0.3252 0.3406 0.3074 0.3221 0.2996 0.3125
+Standard errors in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
 Appendix D Simulation Results
-Table 14: Empirical sizes for $\mathbb{W}_{1}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
-polynomial series $n=100$ $n=300$ $n=500$
+Table 4: Empirical sizes for $\mathbb{W}_{1}$, rejection probabilities at 1%, 5% and 10% levels.
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.072 0.102 0.038 0.072 0.094 0.048 0.076 0.092
-$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.088 0.029 0.069 0.118 0.019 0.078 0.127
-$h=4$ $\mathscr{T}_{n}$ 0.044 0.090 0.118 0.032 0.080 0.110 0.036 0.082 0.110
-$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.088 0.078 0.127 0.137 0.019 0.108 0.147
-$h=8$ $\mathscr{T}_{n}$ 0.058 0.132 0.184 0.032 0.106 0.146 0.038 0.064 0.094
-$\mathscr{T}_{n}^{*}$ 0.049 0.088 0.137 0.028 0.127 0.137 0.019 0.098 0.108
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.100 0.132 0.026 0.076 0.110 0.048 0.082 0.110
-$\mathscr{T}_{n}^{*}$ 0.029 0.137 0.167 0.020 0.088 0.098 0.020 0.059 0.118
-$h=4$ $\mathscr{T}_{n}$ 0.048 0.100 0.142 0.024 0.066 0.100 0.026 0.066 0.100
-$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.186 0.019 0.078 0.118 0.019 0.078 0.137
-$h=8$ $\mathscr{T}_{n}$ 0.060 0.120 0.178 0.042 0.074 0.120 0.018 0.054 0.104
-$\mathscr{T}_{n}^{*}$ 0.038 0.176 0.196 0.039 0.078 0.147 0.020 0.078 0.137
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.062 0.108 0.140 0.028 0.058 0.092 0.026 0.056 0.098
-$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.008 0.020 0.039 0.019 0.088 0.118
-$h=4$ $\mathscr{T}_{n}$ 0.072 0.116 0.170 0.020 0.066 0.116 0.036 0.078 0.108
-$\mathscr{T}_{n}^{*}$ 0.049 0.078 0.127 0.029 0.069 0.088 0.015 0.108 0.147
-$h=8$ $\mathscr{T}_{n}$ 0.094 0.168 0.216 0.026 0.062 0.112 0.034 0.078 0.136
-$\mathscr{T}_{n}^{*}$ 0.058 0.127 0.186 0.029 0.108 0.167 0.014 0.088 0.137
-trigonometric series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.035 0.075 0.115 0.033 0.063 0.098 0.042 0.074 0.107
+chi-p 0.013 0.051 0.112 0.006 0.049 0.096 0.010 0.056 0.106
+$h=4$ asy-p 0.033 0.074 0.112 0.031 0.074 0.107 0.038 0.077 0.114
+chi-p 0.013 0.056 0.106 0.017 0.050 0.105 0.010 0.055 0.106
+$h=8$ asy-p 0.033 0.073 0.117 0.025 0.060 0.091 0.011 0.060 0.096
+chi-p 0.014 0.057 0.108 0.011 0.046 0.088 0.008 0.040 0.091
+7.1 $h=2$ asy-p 0.036 0.077 0.111 0.049 0.086 0.109 0.034 0.071 0.089
+chi-p 0.015 0.048 0.109 0.014 0.062 0.107 0.007 0.044 0.087
+$h=4$ asy-p 0.038 0.090 0.126 0.052 0.092 0.125 0.027 0.070 0.094
+chi-p 0.015 0.059 0.118 0.015 0.070 0.117 0.005 0.045 0.094
+$h=8$ asy-p 0.029 0.069 0.101 0.023 0.072 0.105 0.030 0.075 0.106
+chi-p 0.012 0.055 0.097 0.009 0.053 0.097 0.013 0.054 0.102
+7.1 $h=2$ asy-p 0.042 0.072 0.111 0.037 0.085 0.113 0.025 0.059 0.091
+chi-p 0.012 0.057 0.106 0.011 0.059 0.112 0.004 0.039 0.088
+$h=4$ asy-p 0.038 0.082 0.125 0.031 0.082 0.116 0.025 0.069 0.102
+chi-p 0.013 0.056 0.114 0.012 0.058 0.111 0.007 0.046 0.096
+$h=8$ asy-p 0.023 0.066 0.108 0.031 0.067 0.098 0.027 0.069 0.104
+chi-p 0.011 0.043 0.100 0.012 0.048 0.095 0.011 0.052 0.103
+$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.064 0.098 0.044 0.070 0.104 0.046 0.070 0.092
-$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.078 0.029 0.039 0.088 0.019 0.048 0.118
-$h=4$ $\mathscr{T}_{n}$ 0.052 0.092 0.122 0.030 0.076 0.140 0.032 0.074 0.116
-$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.069 0.039 0.108 0.118 0.019 0.058 0.137
-$h=8$ $\mathscr{T}_{n}$ 0.068 0.138 0.180 0.038 0.110 0.134 0.034 0.066 0.094
-$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.137 0.029 0.118 0.157 0.019 0.038 0.108
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.084 0.112 0.038 0.084 0.118 0.048 0.084 0.114
-$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.176 0.029 0.059 0.108 0.010 0.039 0.088
-$h=4$ $\mathscr{T}_{n}$ 0.050 0.108 0.152 0.022 0.066 0.104 0.032 0.070 0.114
-$\mathscr{T}_{n}^{*}$ 0.069 0.118 0.186 0.049 0.098 0.127 0.019 0.058 0.127
-$h=8$ $\mathscr{T}_{n}$ 0.064 0.108 0.162 0.032 0.086 0.124 0.022 0.052 0.080
-$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.020 0.108 0.157 0.019 0.058 0.157
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.060 0.090 0.120 0.034 0.070 0.112 0.022 0.062 0.098
-$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.098 0.020 0.039 0.078 0.018 0.068 0.137
-$h=4$ $\mathscr{T}_{n}$ 0.064 0.094 0.164 0.022 0.076 0.114 0.032 0.070 0.106
-$\mathscr{T}_{n}^{*}$ 0.078 0.098 0.147 0.039 0.088 0.098 0.017 0.068 0.118
-$h=8$ $\mathscr{T}_{n}$ 0.098 0.158 0.206 0.024 0.060 0.102 0.032 0.078 0.120
-$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.039 0.127 0.167 0.016 0.068 0.127
-Table 15: Empirical sizes for $\mathbb{W}_{2}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
-polynomial series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.037 0.084 0.123 0.037 0.071 0.101 0.033 0.081 0.111
+chi-p 0.012 0.055 0.120 0.011 0.049 0.098 0.008 0.060 0.106
+$h=4$ asy-p 0.040 0.086 0.141 0.034 0.072 0.101 0.030 0.073 0.112
+chi-p 0.011 0.058 0.131 0.011 0.052 0.099 0.014 0.046 0.104
+$h=8$ asy-p 0.020 0.053 0.086 0.021 0.056 0.089 0.024 0.045 0.090
+chi-p 0.009 0.042 0.083 0.008 0.038 0.081 0.008 0.035 0.088
+7.1 $h=2$ asy-p 0.048 0.083 0.110 0.045 0.083 0.115 0.036 0.069 0.100
+chi-p 0.016 0.061 0.108 0.015 0.062 0.113 0.014 0.055 0.099
+$h=4$ asy-p 0.049 0.093 0.129 0.046 0.095 0.125 0.035 0.084 0.117
+chi-p 0.015 0.070 0.121 0.017 0.070 0.119 0.012 0.051 0.115
+$h=8$ asy-p 0.031 0.066 0.106 0.020 0.058 0.086 0.016 0.041 0.086
+chi-p 0.013 0.049 0.103 0.005 0.042 0.085 0.003 0.033 0.086
+7.1 $h=2$ asy-p 0.037 0.074 0.107 0.048 0.078 0.111 0.038 0.080 0.103
+chi-p 0.016 0.049 0.106 0.014 0.060 0.107 0.010 0.055 0.101
+$h=4$ asy-p 0.039 0.089 0.113 0.035 0.071 0.107 0.032 0.071 0.110
+chi-p 0.014 0.069 0.110 0.015 0.057 0.099 0.011 0.047 0.103
+$h=8$ asy-p 0.033 0.063 0.096 0.019 0.037 0.069 0.016 0.052 0.088
+chi-p 0.015 0.050 0.090 0.007 0.032 0.062 0.006 0.032 0.088
+Table 5: Empirical sizes for $\mathbb{W}_{2}$, rejection probabilities at 1%, 5% and 10% levels.
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.088 0.200 0.248 0.094 0.156 0.212 0.090 0.190 0.266
-$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.108 0.049 0.088 0.127 0.020 0.039 0.069
-$h=4$ $\mathscr{T}_{n}$ 0.060 0.150 0.206 0.084 0.170 0.218 0.086 0.172 0.204
-$\mathscr{T}_{n}^{*}$ 0.010 0.039 0.039 0.029 0.049 0.069 0.049 0.127 0.157
-$h=8$ $\mathscr{T}_{n}$ 0.062 0.134 0.194 0.062 0.136 0.174 0.068 0.144 0.202
-$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.108 0.049 0.078 0.098 0.020 0.127 0.137
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.102 0.142 0.182 0.054 0.096 0.152 0.052 0.106 0.136
-$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.118 0.059 0.108 0.127 0.020 0.078 0.176
-$h=4$ $\mathscr{T}_{n}$ 0.098 0.158 0.212 0.042 0.090 0.136 0.036 0.086 0.114
-$\mathscr{T}_{n}^{*}$ 0.020 0.069 0.118 0.069 0.118 0.137 0.029 0.108 0.147
-$h=8$ $\mathscr{T}_{n}$ 0.120 0.210 0.282 0.066 0.122 0.162 0.034 0.088 0.144
-$\mathscr{T}_{n}^{*}$ 0.010 0.059 0.088 0.049 0.127 0.157 0.029 0.069 0.167
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.068 0.118 0.142 0.024 0.054 0.084 0.028 0.078 0.092
-$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.000 0.020 0.088 0.069 0.069 0.108
-$h=4$ $\mathscr{T}_{n}$ 0.054 0.106 0.164 0.024 0.062 0.098 0.048 0.076 0.108
-$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.069 0.020 0.059 0.088 0.039 0.098 0.127
-$h=8$ $\mathscr{T}_{n}$ 0.062 0.138 0.174 0.028 0.058 0.092 0.032 0.084 0.132
-$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.029 0.069 0.098 0.039 0.078 0.127
-trigonometric series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.037 0.071 0.099 0.033 0.061 0.095 0.048 0.085 0.124
+chi-p 0.008 0.055 0.096 0.007 0.040 0.092 0.016 0.061 0.124
+$h=4$ asy-p 0.046 0.083 0.109 0.030 0.064 0.097 0.046 0.078 0.118
+chi-p 0.024 0.057 0.108 0.010 0.046 0.093 0.016 0.061 0.115
+$h=8$ asy-p 0.079 0.111 0.138 0.031 0.059 0.079 0.026 0.047 0.082
+chi-p 0.053 0.105 0.132 0.012 0.046 0.078 0.011 0.036 0.085
+7.1 $h=2$ asy-p 0.022 0.050 0.079 0.024 0.056 0.084 0.034 0.075 0.102
+chi-p 0.004 0.034 0.079 0.010 0.035 0.080 0.003 0.047 0.102
+$h=4$ asy-p 0.040 0.071 0.104 0.027 0.057 0.086 0.026 0.066 0.104
+chi-p 0.016 0.052 0.099 0.009 0.040 0.082 0.007 0.043 0.098
+$h=8$ asy-p 0.069 0.102 0.130 0.025 0.049 0.072 0.017 0.042 0.089
+chi-p 0.047 0.086 0.124 0.011 0.041 0.072 0.009 0.029 0.091
+7.1 $h=2$ asy-p 0.029 0.069 0.087 0.035 0.070 0.104 0.028 0.072 0.106
+chi-p 0.010 0.041 0.085 0.005 0.051 0.104 0.010 0.048 0.104
+$h=4$ asy-p 0.039 0.073 0.103 0.035 0.066 0.102 0.034 0.067 0.107
+chi-p 0.018 0.062 0.098 0.009 0.048 0.095 0.008 0.051 0.101
+$h=8$ asy-p 0.069 0.113 0.146 0.022 0.053 0.083 0.019 0.046 0.092
+chi-p 0.049 0.094 0.143 0.012 0.039 0.077 0.008 0.036 0.092
+$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.082 0.160 0.212 0.106 0.198 0.246 0.084 0.176 0.232
-$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.078 0.108 0.127 0.020 0.088 0.118
-$h=4$ $\mathscr{T}_{n}$ 0.072 0.146 0.208 0.076 0.160 0.204 0.086 0.150 0.208
-$\mathscr{T}_{n}^{*}$ 0.029 0.118 0.118 0.029 0.039 0.078 0.029 0.069 0.098
-$h=8$ $\mathscr{T}_{n}$ 0.086 0.144 0.194 0.076 0.154 0.204 0.084 0.152 0.184
-$\mathscr{T}_{n}^{*}$ 0.039 0.108 0.137 0.029 0.078 0.098 0.049 0.108 0.118
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.086 0.134 0.180 0.056 0.112 0.140 0.056 0.108 0.144
-$\mathscr{T}_{n}^{*}$ 0.039 0.078 0.118 0.039 0.108 0.137 0.029 0.049 0.147
-$h=4$ $\mathscr{T}_{n}$ 0.102 0.188 0.240 0.040 0.100 0.140 0.052 0.092 0.138
-$\mathscr{T}_{n}^{*}$ 0.020 0.059 0.069 0.078 0.137 0.167 0.049 0.108 0.147
-$h=8$ $\mathscr{T}_{n}$ 0.112 0.192 0.270 0.052 0.106 0.160 0.032 0.076 0.132
-$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.049 0.029 0.167 0.176 0.049 0.118 0.167
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.096 0.120 0.026 0.064 0.098 0.026 0.072 0.102
-$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.069 0.020 0.059 0.078 0.059 0.098 0.127
-$h=4$ $\mathscr{T}_{n}$ 0.040 0.104 0.150 0.022 0.066 0.104 0.034 0.068 0.118
-$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.118 0.049 0.088 0.088 0.059 0.088 0.127
-$h=8$ $\mathscr{T}_{n}$ 0.056 0.132 0.158 0.022 0.064 0.096 0.024 0.076 0.130
-$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.078 0.029 0.078 0.127 0.049 0.088 0.108
-Table 16: Empirical sizes for $\mathbb{W}_{3}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ at 1%, 5% and 10% level.
-polynomial series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.036 0.064 0.085 0.036 0.062 0.088 0.040 0.074 0.097
+chi-p 0.015 0.046 0.083 0.011 0.045 0.088 0.018 0.052 0.097
+$h=4$ asy-p 0.046 0.089 0.116 0.037 0.060 0.092 0.040 0.074 0.104
+chi-p 0.028 0.063 0.113 0.019 0.050 0.086 0.018 0.052 0.097
+$h=8$ asy-p 0.097 0.139 0.170 0.042 0.061 0.083 0.024 0.054 0.086
+chi-p 0.065 0.122 0.166 0.028 0.053 0.077 0.006 0.043 0.088
+7.1 $h=2$ asy-p 0.028 0.053 0.082 0.026 0.055 0.075 0.032 0.071 0.099
+chi-p 0.013 0.032 0.081 0.006 0.031 0.074 0.010 0.048 0.096
+$h=4$ asy-p 0.044 0.079 0.109 0.027 0.054 0.073 0.032 0.063 0.095
+chi-p 0.023 0.058 0.106 0.011 0.043 0.072 0.010 0.043 0.087
+$h=8$ asy-p 0.070 0.110 0.140 0.028 0.054 0.073 0.026 0.053 0.088
+chi-p 0.050 0.094 0.137 0.018 0.044 0.070 0.006 0.042 0.090
+7.1 $h=2$ asy-p 0.029 0.052 0.070 0.027 0.061 0.086 0.027 0.054 0.077
+chi-p 0.010 0.039 0.067 0.009 0.039 0.085 0.009 0.037 0.074
+$h=4$ asy-p 0.055 0.093 0.128 0.027 0.058 0.089 0.031 0.060 0.089
+chi-p 0.029 0.071 0.122 0.012 0.036 0.086 0.014 0.043 0.082
+$h=8$ asy-p 0.090 0.128 0.157 0.029 0.053 0.075 0.022 0.044 0.085
+chi-p 0.061 0.112 0.150 0.015 0.040 0.068 0.012 0.035 0.086
+Table 6: Empirical sizes for $\mathbb{W}_{3}$, rejection probabilities at 1%, 5% and 10% levels.
+$d_{\tau}=2$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-$r=2$ 7.1 $h=2$ 0.032 0.096 0.164 0.040 0.096 0.116 0.014 0.056 0.096
-$h=4$ 0.028 0.056 0.084 0.024 0.052 0.084 0.036 0.060 0.080
-$h=8$ 0.016 0.036 0.052 0.024 0.036 0.048 0.024 0.040 0.092
-7.1 $h=2$ 0.040 0.088 0.148 0.044 0.096 0.120 0.020 0.060 0.112
-$h=4$ 0.040 0.088 0.104 0.032 0.100 0.120 0.012 0.044 0.066
-$h=8$ 0.016 0.048 0.064 0.008 0.016 0.028 0.028 0.056 0.092
-7.1 $h=2$ 0.040 0.104 0.116 0.024 0.076 0.104 0.048 0.092 0.124
-$h=4$ 0.028 0.052 0.088 0.032 0.048 0.068 0.024 0.072 0.104
-$h=8$ 0.020 0.048 0.052 0.024 0.056 0.072 0.004 0.024 0.052
-$r=4$ 7.1 $h=2$ 0.032 0.120 0.156 0.040 0.084 0.108 0.016 0.052 0.106
-$h=4$ 0.036 0.064 0.092 0.032 0.076 0.088 0.020 0.040 0.076
-$h=8$ 0.012 0.032 0.040 0.012 0.032 0.048 0.004 0.032 0.044
-7.1 $h=2$ 0.032 0.100 0.136 0.032 0.076 0.092 0.048 0.100 0.144
-$h=4$ 0.044 0.096 0.100 0.052 0.092 0.116 0.004 0.028 0.056
-$h=8$ 0.008 0.044 0.056 0.008 0.036 0.052 0.024 0.052 0.080
-7.1 $h=2$ 0.056 0.092 0.128 0.048 0.080 0.132 0.060 0.100 0.132
-$h=4$ 0.028 0.060 0.084 0.008 0.044 0.056 0.020 0.068 0.120
-$h=8$ 0.032 0.044 0.060 0.024 0.060 0.084 0.020 0.048 0.064
-trigonometric series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.037 0.071 0.099 0.033 0.061 0.095 0.048 0.085 0.124
+chi-p 0.008 0.055 0.096 0.007 0.040 0.092 0.016 0.061 0.124
+$h=4$ asy-p 0.046 0.083 0.109 0.030 0.064 0.097 0.046 0.078 0.118
+chi-p 0.024 0.057 0.108 0.010 0.046 0.093 0.016 0.061 0.115
+$h=8$ asy-p 0.079 0.111 0.138 0.031 0.059 0.079 0.026 0.047 0.082
+chi-p 0.053 0.105 0.132 0.012 0.046 0.078 0.011 0.036 0.085
+7.1 $h=2$ asy-p 0.022 0.050 0.079 0.024 0.056 0.084 0.034 0.075 0.102
+chi-p 0.004 0.034 0.079 0.010 0.035 0.080 0.003 0.047 0.102
+$h=4$ asy-p 0.040 0.071 0.104 0.027 0.057 0.086 0.026 0.066 0.104
+chi-p 0.016 0.052 0.099 0.009 0.040 0.082 0.007 0.043 0.098
+$h=8$ asy-p 0.069 0.102 0.130 0.025 0.049 0.072 0.017 0.042 0.089
+chi-p 0.047 0.086 0.124 0.011 0.041 0.072 0.009 0.029 0.091
+7.1 $h=2$ asy-p 0.029 0.069 0.087 0.035 0.070 0.104 0.028 0.072 0.106
+chi-p 0.010 0.041 0.085 0.005 0.051 0.104 0.010 0.048 0.104
+$h=4$ asy-p 0.039 0.073 0.103 0.035 0.066 0.102 0.034 0.067 0.107
+chi-p 0.018 0.062 0.098 0.009 0.048 0.095 0.008 0.051 0.101
+$h=8$ asy-p 0.069 0.113 0.146 0.022 0.053 0.083 0.019 0.046 0.092
+chi-p 0.049 0.094 0.143 0.012 0.039 0.077 0.008 0.036 0.092
+$d_{\tau}=4$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-$r=2$ 7.1 $h=2$ 0.028 0.092 0.160 0.030 0.096 0.116 0.012 0.060 0.106
-$h=4$ 0.020 0.060 0.084 0.024 0.052 0.084 0.028 0.056 0.072
-$h=8$ 0.008 0.032 0.048 0.020 0.032 0.048 0.024 0.040 0.072
-7.1 $h=2$ 0.044 0.096 0.152 0.044 0.096 0.120 0.014 0.058 0.116
-$h=4$ 0.040 0.080 0.108 0.032 0.104 0.116 0.012 0.040 0.088
-$h=8$ 0.024 0.044 0.064 0.008 0.016 0.028 0.032 0.060 0.088
-7.1 $h=2$ 0.036 0.100 0.112 0.024 0.076 0.104 0.048 0.092 0.132
-$h=4$ 0.016 0.052 0.092 0.028 0.048 0.068 0.028 0.068 0.104
-$h=8$ 0.012 0.036 0.048 0.032 0.056 0.068 0.004 0.028 0.052
-$r=4$ 7.1 $h=2$ 0.032 0.116 0.156 0.040 0.084 0.108 0.018 0.050 0.102
-$h=4$ 0.032 0.060 0.088 0.032 0.076 0.088 0.016 0.028 0.080
-$h=8$ 0.012 0.032 0.048 0.012 0.032 0.048 0.004 0.032 0.098
-7.1 $h=2$ 0.032 0.104 0.136 0.032 0.076 0.092 0.014 0.062 0.104
-$h=4$ 0.040 0.100 0.108 0.052 0.092 0.116 0.008 0.028 0.076
-$h=8$ 0.024 0.056 0.064 0.008 0.040 0.056 0.024 0.056 0.110
-7.1 $h=2$ 0.060 0.096 0.136 0.028 0.080 0.132 0.014 0.048 0.102
-$h=4$ 0.036 0.064 0.088 0.008 0.044 0.056 0.024 0.068 0.112
-$h=8$ 0.020 0.044 0.060 0.020 0.056 0.080 0.016 0.048 0.094
-Table 17: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\delta(z)$.
-polynomial series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.036 0.064 0.085 0.036 0.062 0.088 0.040 0.074 0.097
+chi-p 0.015 0.046 0.083 0.011 0.045 0.088 0.018 0.052 0.097
+$h=4$ asy-p 0.046 0.089 0.116 0.037 0.060 0.092 0.040 0.074 0.104
+chi-p 0.028 0.063 0.113 0.019 0.050 0.086 0.018 0.052 0.097
+$h=8$ asy-p 0.097 0.139 0.170 0.042 0.061 0.083 0.024 0.054 0.086
+chi-p 0.065 0.122 0.166 0.028 0.053 0.077 0.006 0.043 0.088
+7.1 $h=2$ asy-p 0.028 0.053 0.082 0.026 0.055 0.075 0.032 0.071 0.099
+chi-p 0.013 0.032 0.081 0.006 0.031 0.074 0.010 0.048 0.096
+$h=4$ asy-p 0.044 0.079 0.109 0.027 0.054 0.073 0.032 0.063 0.095
+chi-p 0.023 0.058 0.106 0.011 0.043 0.072 0.010 0.043 0.087
+$h=8$ asy-p 0.070 0.110 0.140 0.028 0.054 0.073 0.026 0.053 0.088
+chi-p 0.050 0.094 0.137 0.018 0.044 0.070 0.006 0.042 0.090
+7.1 $h=2$ asy-p 0.029 0.052 0.070 0.027 0.061 0.086 0.027 0.054 0.077
+chi-p 0.010 0.039 0.067 0.009 0.039 0.085 0.009 0.037 0.074
+$h=4$ asy-p 0.055 0.093 0.128 0.027 0.058 0.089 0.031 0.060 0.089
+chi-p 0.029 0.071 0.122 0.012 0.036 0.086 0.014 0.043 0.082
+$h=8$ asy-p 0.090 0.128 0.157 0.029 0.053 0.075 0.022 0.044 0.085
+chi-p 0.061 0.112 0.150 0.015 0.040 0.068 0.012 0.035 0.086
+Table 7: Empirical sizes for $\mathbb{W}_{4}$, rejection prob. at 1%, 5% and 10% levels. Testing $\delta(z)$.
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.088 0.128 0.064 0.104 0.156 0.044 0.112 0.152
-$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.064 0.004 0.022 0.074 0.012 0.036 0.082
-$h=4$ $\mathscr{T}_{n}$ 0.044 0.080 0.116 0.016 0.048 0.080 0.040 0.076 0.096
-$\mathscr{T}_{n}^{*}$ 0.000 0.024 0.076 0.000 0.040 0.072 0.012 0.036 0.084
-$h=8$ $\mathscr{T}_{n}$ 0.008 0.024 0.040 0.004 0.008 0.012 0.004 0.016 0.032
-$\mathscr{T}_{n}^{*}$ 0.020 0.064 0.072 0.008 0.020 0.076 0.024 0.052 0.090
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.048 0.124 0.160 0.032 0.060 0.108 0.060 0.136 0.172
-$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.024 0.004 0.012 0.016 0.016 0.028 0.060
-$h=4$ $\mathscr{T}_{n}$ 0.060 0.112 0.132 0.028 0.068 0.104 0.044 0.096 0.128
-$\mathscr{T}_{n}^{*}$ 0.000 0.012 0.036 0.000 0.000 0.020 0.008 0.012 0.024
-$h=8$ $\mathscr{T}_{n}$ 0.016 0.044 0.056 0.008 0.032 0.056 0.012 0.040 0.068
-$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.076 0.004 0.038 0.072 0.006 0.044 0.072
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.104 0.140 0.052 0.116 0.168 0.028 0.104 0.156
-$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.068 0.000 0.060 0.076 0.000 0.068 0.094
-$h=4$ $\mathscr{T}_{n}$ 0.004 0.016 0.024 0.000 0.004 0.012 0.044 0.068 0.108
-$\mathscr{T}_{n}^{*}$ 0.024 0.072 0.100 0.016 0.044 0.080 0.012 0.040 0.088
-$h=8$ $\mathscr{T}_{n}$ 0.000 0.028 0.048 0.012 0.032 0.068 0.016 0.036 0.068
-$\mathscr{T}_{n}^{*}$ 0.006 0.032 0.084 0.000 0.034 0.078 0.004 0.040 0.076
-trigonometric series $n=100$ $n=300$ $n=500$
+7.1 $h=2$ asy-p 0.065 0.080 0.129 0.053 0.086 0.110 0.050 0.072 0.090
+chi-p 0.022 0.053 0.124 0.027 0.069 0.109 0.021 0.059 0.089
+$h=4$ asy-p 0.042 0.085 0.112 0.071 0.106 0.130 0.047 0.072 0.098
+chi-p 0.039 0.071 0.103 0.026 0.090 0.126 0.029 0.061 0.092
+$h=8$ asy-p 0.169 0.124 0.159 0.019 0.060 0.074 0.022 0.039 0.071
+chi-p 0.144 0.101 0.154 0.025 0.054 0.071 0.016 0.036 0.080
+7.1 $h=2$ asy-p 0.076 0.084 0.144 0.056 0.078 0.090 0.041 0.060 0.070
+chi-p 0.031 0.055 0.137 0.028 0.071 0.087 0.017 0.053 0.069
+$h=4$ asy-p 0.134 0.085 0.133 0.075 0.097 0.115 0.054 0.077 0.093
+chi-p 0.099 0.064 0.125 0.015 0.065 0.111 0.014 0.060 0.091
+$h=8$ asy-p 0.174 0.124 0.153 0.053 0.081 0.094 0.025 0.045 0.080
+chi-p 0.142 0.103 0.150 0.022 0.072 0.092 0.015 0.036 0.083
+7.1 $h=2$ asy-p 0.032 0.059 0.126 0.058 0.080 0.098 0.051 0.077 0.095
+chi-p 0.018 0.039 0.122 0.033 0.067 0.094 0.017 0.062 0.093
+$h=4$ asy-p 0.072 0.094 0.141 0.060 0.100 0.129 0.055 0.084 0.104
+chi-p 0.095 0.070 0.137 0.041 0.077 0.122 0.016 0.070 0.101
+$h=8$ asy-p 0.063 0.089 0.151 0.047 0.067 0.083 0.025 0.049 0.076
+chi-p 0.109 0.070 0.147 0.020 0.063 0.078 0.016 0.036 0.088
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
 1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.104 0.144 0.044 0.092 0.116 0.048 0.116 0.160
-$\mathscr{T}_{n}^{*}$ 0.004 0.022 0.068 0.016 0.020 0.076 0.004 0.034 0.088
-$h=4$ $\mathscr{T}_{n}$ 0.012 0.044 0.076 0.032 0.076 0.116 0.052 0.116 0.132
-$\mathscr{T}_{n}^{*}$ 0.004 0.024 0.074 0.004 0.030 0.080 0.004 0.038 0.078
-$h=8$ $\mathscr{T}_{n}$ 0.000 0.004 0.004 0.000 0.004 0.008 0.008 0.020 0.032
-$\mathscr{T}_{n}^{*}$ 0.016 0.036 0.078 0.008 0.020 0.080 0.016 0.028 0.082
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.108 0.132 0.032 0.084 0.108 0.040 0.112 0.148
-$\mathscr{T}_{n}^{*}$ 0.000 0.008 0.024 0.000 0.000 0.012 0.008 0.024 0.056
-$h=4$ $\mathscr{T}_{n}$ 0.048 0.076 0.116 0.044 0.072 0.104 0.036 0.088 0.128
-$\mathscr{T}_{n}^{*}$ 0.008 0.012 0.028 0.000 0.004 0.012 0.000 0.016 0.044
-$h=8$ $\mathscr{T}_{n}$ 0.012 0.060 0.072 0.012 0.024 0.040 0.016 0.040 0.116
-$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.072 0.004 0.038 0.068 0.008 0.040 0.092
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.128 0.168 0.044 0.084 0.128 0.040 0.108 0.152
-$\mathscr{T}_{n}^{*}$ 0.004 0.030 0.064 0.000 0.044 0.088 0.000 0.046 0.094
-$h=4$ $\mathscr{T}_{n}$ 0.016 0.024 0.040 0.008 0.056 0.080 0.044 0.104 0.136
-$\mathscr{T}_{n}^{*}$ 0.044 0.088 0.100 0.012 0.030 0.068 0.008 0.040 0.094
-$h=8$ $\mathscr{T}_{n}$ 0.032 0.044 0.060 0.008 0.036 0.052 0.020 0.040 0.068
-$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.082 0.004 0.028 0.062 0.006 0.032 0.068
-Table 18: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\lambda(z)$.
-polynomial series $n=100$ $n=300$ $n=500$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.136 0.168 0.048 0.096 0.140 0.060 0.124 0.156
-$\mathscr{T}_{n}^{*}$ 0.024 0.048 0.088 0.020 0.048 0.084 0.004 0.036 0.082
-$h=4$ $\mathscr{T}_{n}$ 0.040 0.068 0.096 0.040 0.080 0.100 0.032 0.088 0.132
-$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.094 0.012 0.032 0.076 0.004 0.052 0.092
-$h=8$ $\mathscr{T}_{n}$ 0.020 0.056 0.100 0.024 0.068 0.084 0.040 0.096 0.116
-$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.088 0.012 0.036 0.068 0.020 0.036 0.100
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.076 0.108 0.036 0.104 0.136 0.072 0.124 0.168
-$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.052 0.012 0.032 0.076 0.020 0.052 0.104
-$h=4$ $\mathscr{T}_{n}$ 0.032 0.108 0.152 0.028 0.076 0.120 0.060 0.108 0.136
-$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.068 0.008 0.020 0.068 0.020 0.040 0.092
-$h=8$ $\mathscr{T}_{n}$ 0.044 0.088 0.100 0.020 0.040 0.106 0.012 0.040 0.096
-$\mathscr{T}_{n}^{*}$ 0.020 0.060 0.080 0.012 0.028 0.036 0.012 0.028 0.040
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.136 0.192 0.048 0.116 0.148 0.072 0.184 0.216
-$\mathscr{T}_{n}^{*}$ 0.008 0.048 0.080 0.012 0.040 0.068 0.020 0.064 0.108
-$h=4$ $\mathscr{T}_{n}$ 0.028 0.064 0.108 0.032 0.076 0.092 0.072 0.144 0.204
-$\mathscr{T}_{n}^{*}$ 0.008 0.032 0.072 0.008 0.028 0.056 0.024 0.056 0.128
-$h=8$ $\mathscr{T}_{n}$ 0.036 0.072 0.076 0.020 0.076 0.112 0.028 0.064 0.080
-$\mathscr{T}_{n}^{*}$ 0.012 0.044 0.088 0.004 0.036 0.076 0.016 0.040 0.086
-trigonometric series $n=100$ $n=300$ $n=500$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.112 0.152 0.048 0.100 0.124 0.044 0.108 0.132
-$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.088 0.016 0.040 0.094 0.008 0.058 0.082
-$h=4$ $\mathscr{T}_{n}$ 0.048 0.096 0.108 0.012 0.072 0.108 0.056 0.116 0.136
-$\mathscr{T}_{n}^{*}$ 0.016 0.044 0.088 0.004 0.038 0.082 0.012 0.048 0.088
-$h=8$ $\mathscr{T}_{n}$ 0.032 0.056 0.080 0.028 0.076 0.096 0.028 0.064 0.072
-$\mathscr{T}_{n}^{*}$ 0.024 0.036 0.072 0.008 0.036 0.080 0.024 0.036 0.088
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.092 0.136 0.064 0.112 0.172 0.056 0.152 0.200
-$\mathscr{T}_{n}^{*}$ 0.008 0.044 0.084 0.004 0.028 0.080 0.016 0.056 0.116
-$h=4$ $\mathscr{T}_{n}$ 0.036 0.076 0.104 0.020 0.060 0.084 0.052 0.096 0.112
-$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.052 0.012 0.024 0.056 0.008 0.044 0.080
-$h=8$ $\mathscr{T}_{n}$ 0.028 0.064 0.092 0.028 0.056 0.092 0.016 0.044 0.094
-$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.064 0.012 0.040 0.064 0.008 0.020 0.048
-7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.128 0.164 0.064 0.116 0.164 0.040 0.112 0.132
-$\mathscr{T}_{n}^{*}$ 0.004 0.028 0.076 0.016 0.040 0.072 0.024 0.048 0.092
-$h=4$ $\mathscr{T}_{n}$ 0.020 0.076 0.092 0.036 0.088 0.116 0.028 0.088 0.116
-$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.048 0.012 0.040 0.084 0.000 0.036 0.088
-$h=8$ $\mathscr{T}_{n}$ 0.024 0.052 0.076 0.012 0.044 0.072 0.024 0.048 0.060
-$\mathscr{T}_{n}^{*}$ 0.012 0.048 0.090 0.008 0.050 0.086 0.000 0.020 0.072
+7.1 $h=2$ asy-p 0.086 0.117 0.135 0.061 0.086 0.099 0.028 0.053 0.076
+chi-p 0.061 0.101 0.132 0.036 0.075 0.098 0.016 0.041 0.082
+$h=4$ asy-p 0.122 0.163 0.187 0.063 0.084 0.111 0.041 0.059 0.068
+chi-p 0.091 0.144 0.184 0.044 0.072 0.107 0.030 0.048 0.082
+$h=8$ asy-p 0.145 0.195 0.227 0.043 0.065 0.089 0.028 0.047 0.058
+chi-p 0.109 0.177 0.225 0.030 0.057 0.083 0.018 0.038 0.081
+7.1 $h=2$ asy-p 0.083 0.113 0.141 0.057 0.081 0.098 0.034 0.051 0.067
+chi-p 0.052 0.101 0.136 0.036 0.067 0.096 0.018 0.042 0.085
+$h=4$ asy-p 0.117 0.154 0.183 0.060 0.090 0.113 0.028 0.046 0.063
+chi-p 0.087 0.138 0.176 0.045 0.078 0.109 0.019 0.040 0.081
+$h=8$ asy-p 0.134 0.188 0.218 0.041 0.067 0.086 0.027 0.046 0.068
+chi-p 0.105 0.167 0.214 0.032 0.060 0.085 0.016 0.036 0.081
+7.1 $h=2$ asy-p 0.093 0.131 0.148 0.045 0.066 0.091 0.036 0.058 0.069
+chi-p 0.057 0.114 0.145 0.030 0.056 0.087 0.020 0.048 0.087
+$h=4$ asy-p 0.135 0.177 0.203 0.064 0.088 0.107 0.030 0.047 0.076
+chi-p 0.090 0.163 0.200 0.043 0.074 0.104 0.018 0.038 0.084
+$h=8$ asy-p 0.148 0.202 0.228 0.060 0.080 0.092 0.022 0.036 0.072
+chi-p 0.115 0.179 0.227 0.039 0.077 0.089 0.016 0.030 0.082

### v3

@@
 A simulation study confirms that our tests perform very well in finite samples.
-JEL codes: C21, C31
-Keywords: Spatial autoregression, varying coefficients, nonparametric inference
-1.  Introduction
+Keywords: Spatial autoregression, varying coefficients, inference
+JEL classification: C21, C31
+1 Introduction
 This paper develops a general framework for inference on varying coefficient regression models that is as easy to implement in practice as familiar Wald tests for finite-dimensional parameters.
@@
 In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry.
-This can be seen as a follow-up study to Li et al.
-(2002), who suggest that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e.
-output elasticities of capital and labor sum to unity) for most observed managerial expense levels.
-Our analysis confirms their results as we cannot reject the CRS hypothesis using a more general version of their model that permits spatial dependence.
-This suggests that CRS technology may well be a salient feature of this industry.
+This re-visits the empirical study in Li et al.
+(2002), who showed that output elasticities of capital and labor vary with managerial expense.
+They also noted that the returns to scale, which is the sum of output elasticities of inputs, are close to one at many managerial expense levels suggesting CRS, but they did not have a test for this hypothesis.
+Using a newer dataset, our methods can test for CRS while allowing for spatial dependence.
+We do not reject the CRS hypothesis in all settings, which suggests that CRS technology is a salient feature of this industry.
 Our second study re-examines the Boston house price data of Harrison Jr and Rubinfeld (1978).
@@
 Section 7 presents the simulation study, and Section 8 concludes.
-Proofs are in the appendix, while additional empirical and simulation results are in the online appendix.
-2.  Baseline higher-order spatial autoregression
+Proofs and additional simulation results are in the online appendix.
+2 Baseline higher-order spatial autoregression
 Consider a vector of unknown functions $\delta\left(\cdot\right)=\left(\delta_{1}\left(\cdot\right),\ldots,\delta_{d_{\delta}}\left(\cdot\right)\right)^{\prime}$, and the model
@@
 Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$.
-Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$ with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$.
+Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$, with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$.
 Then we can write (2.1) as
 $y_{i}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{i,j}^{\prime}y+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (2.3)
-where $u_{i}=r_{i}+\epsilon_{i}$ and $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error.
+with $u_{i}=r_{i}+\epsilon_{i}$, where $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error.
 Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$.
-Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can then be written as
+Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can be written in matrix notation as
 $y=Q\lambda+X\beta+\Psi\alpha+u=L\xi+u,$ (2.4)
-where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. The null to approximate $H_{0}^{true}$ is
+where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. The null hypothesis to approximate $H_{0}^{true}$ is
 $H_{0}:\alpha=0,$ (2.5)
@@
 Let $k_{ri}$ be elements of the instrumental variable matrix, corresponding to $K_{1},K_{2}$ and $K_{3}$ as defined in the following sections, and $l_{ri}$ denote the elements of $L$.
-Assumption 1.
+Assumption 2.1.
 $\sup_{i\geq 1}\mathrm{E}\left(r_{i}^{2}\right)=o(n^{-1})$.
-Assumption 2.
+Assumption 2.2.
 $\mathrm{E}\left|\epsilon_{i}\right|^{q}<C$, for some $q>4$.
-Assumption 3.
+Assumption 2.3.
 $\mathrm{E}\left(l_{ri}^{2}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
-Assumption 1 controls the approximation error.
+Assumption 2.1 controls the approximation error.
 Sufficient conditions for various cases can be found in Chen (2007).
 We avoid using specific rates because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions.
-Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition.
-Assumption 3 imposes moment conditions on regressors and instruments.
+Assumption 2.2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition.
+Assumption 2.3 imposes moment conditions on regressors and instruments.
 To ease notation, we introduce the following definition that captures boundedness and non-multicollinearity of random or fixed matrices of growing dimension.
@@
 $\limsup_{n\rightarrow\infty}\overline{\alpha}(A)<\infty\text{ and }\liminf_{n\rightarrow\infty}\underline{\alpha}(A)>0.$
-Assumption 4.
+Assumption 2.4.
 $n^{-1}K_{1}^{\prime}K_{1}$, $n^{-1}L^{\prime}K_{1}K_{1}^{\prime}L$ and $n^{-1}\Psi^{\prime}\Psi$ have 1.
@@
 Theorem 2.1.
-Under $H_{0}$, Assumptions 1-4 and
+Under $H_{0}$, Assumptions 2.1-2.4 and
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}}{n}\rightarrow 0,$ (2.8)
@@
 We remind the reader that results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
-Assumption 5.
+Assumption 2.5.
 $\left(I_{n}-\sum_{j=1}^{d_{\lambda}}\lambda_{j}W_{j}\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$.
@@
 Theorem $2.$
-Under Assumptions 1-5 and
+Under Assumptions 2.1-2.5 and
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}\rightarrow 0,$ (2.9)
@@
 (iii) Under the sequence of local alternatives $H_{\ell 1}$, $\mathbb{W}_{1}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
-3.  SAR with error spatial dependence
+3 SAR with error spatial dependence
 Having demonstrated our methods with a baseline model, we now consider (2.1) but relax the $i.i.d$ assumption on $\epsilon_{i}$ to
@@
 We introduce the following assumptions:
-Assumption 6.
+Assumption 3.1.
 $\Sigma$ has 1.
-Assumption 7.
+Assumption 3.2.
 $\sup_{i\geq 1}\sum_{j=1}^{n}\left|b_{ij}\right|+\sup_{j\geq 1}\sum_{i=1}^{n}\left|b_{ij}\right|<\infty$.
-Assumption 8.
-$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
-Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g.
+Assumption 3.3.
+$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel function that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
+Assumption 3.2 restricts the spatial dependence in the errors to a manageable degree, see e.g.
 Kelejian and Prucha (2007) and Delgado and Robinson (2015) for similar assumptions.
-Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g.
+Assumption 3.3 is a standard assumption on kernels in the HAC setting, see e.g.
 Kelejian and Prucha (2007).
@@
 Also let $\sigma_{ij}$ be a typical element of $\Sigma$.
-Assumption 9.
-(a) $\mathrm{E}(\ell^{2})=o\left(n^{2\eta}\right)$ where $\eta<\frac{1}{2}(q-2)/(q-1)$ with $q>4$ in Assumption 2.
+Assumption 3.4.
+(a) $\mathrm{E}(\ell^{2})=o\left(n^{2\eta}\right)$ where $\eta<\frac{1}{2}(q-2)/(q-1)$ with $q>4$ in Assumption 2.2.
 (b) $\sum_{j=1}^{n}\left|\sigma_{ij}\right|d^{\chi}_{ij,1}<C$ for some $\chi\geq 1$.
@@
 3.2 Asymptotic properties
-In all subsequent lemmas/theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9.
+In all subsequent lemmas/theorems, $q$ is defined in Assumption 2.2 and $\eta$ is defined in Assumption 3.4.
 In Lemma 3.1 and Theorems 3.1-3.2, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$.
@@
 Lemma 3.1.
-Let Assumptions 1, 3, 4, 6-9 hold, and
+Let Assumptions 2.1, 2.3, 2.4, 3.1-3.4 hold, and
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\xi}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (3.4)
@@
 Theorem 3.1.
-Let (3.4) hold. Then, under $H_{0}$, Assumptions 1, 3, 4, 6-9, as $n\rightarrow\infty$,
+Let (3.4) hold. Then, under $H_{0}$, Assumptions 2.1, 2.3, 2.4, 3.1-3.4, as $n\rightarrow\infty$,
 $\mathbb{W}_{2}-\frac{v^{\prime}\mathcal{M}_{2}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
 To derive the asymptotic distribution of $\mathbb{W}_{2}$, we need the following assumption, as also in Delgado and Robinson (2015).
-Assumption 10.
+Assumption 3.5.
 $\mathrm{E}\left|v_{i}\right|^{s}<C$, for some $s\geq 8$.
@@
 Theorem 3.2.
-Under Assumptions 1, 3-10, with
+Under Assumptions 2.1, 2.3-3.5, with
 $\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}+n^{\eta-1}d_{\xi}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,$ (3.5)
@@
 Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies ${d_{\xi}^{3}}/{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
-4.  Misspecification robustness
+4 Misspecification robustness
 Consider a degree of misspecification robustness in (2.1), where $\epsilon_{i}$ has the structure in Section 3.
@@
 4.1 Test statistic
-Following Pinkse et al. (2002), we approximate $g\left(d_{ij}\right)$ with a series of basis functions
+Following Pinkse et al. (2002), we approximate $g\left(d_{ij}\right)$ of $G$ with a series of basis functions
 $g(d_{ij})=\sum_{l=1}^{\infty}\tau_{l}e_{l}(d_{ij}),$ (4.1)
@@
 $y_{in}=g_{in}^{\prime}y+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (4.2)
-where $g_{in}\equiv g_{i}$ is the $i$-th row of $G$, and, once again dropping $n$-subscripting, accordingly extend (2.3) to
+and (2.3) to
 $y_{i}=c_{i}^{\prime}\tau+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (4.3)
 with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$.
-In matrix notation, we now have
+In the matrix notation, we now have
 $y=\mathfrak{C}\tau+X\beta+\Psi\alpha+u=F\theta+u,$ (4.4)
@@
 Denote the elements of $F$ as $f_{ri}$. We have following assumptions.
-Assumption 11.
+Assumption 4.1.
 $\mathrm{E}\left(f^{2}_{ri}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
-Assumption 12.
+Assumption 4.2.
 $\sup_{i\geq 1}\mathrm{E}\left(r_{iG}^{2}\right)=o(n^{-1})$ and $\sup_{i\geq 1}\mathrm{E}\left(r_{i\delta}^{2}\right)=o(n^{-1})$.
-Assumption 13.
+Assumption 4.3.
 $n^{-1}K_{2}^{\prime}K_{2}$ and $n^{-1}F^{\prime}K_{2}K_{2}^{\prime}F$ have 1.
 Lemma 4.1.
-Let Assumptions 6-9, 11-13 hold, and
+Let Assumptions 3.1-3.4, 4.1-4.3 hold, and
 $\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\theta}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (4.7)
@@
 Theorem 4.1.
-Let (4.7) hold. Then, under $H_{0}$, Assumptions 6-9, 11-13, as $n\rightarrow\infty$,
+Let (4.7) hold. Then, under $H_{0}$, Assumptions 3.1-3.4, 4.1-4.3, as $n\rightarrow\infty$,
 $\mathbb{W}_{3}-\frac{v^{\prime}\mathcal{M}_{3}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
 Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proofs.
-The following assumption generalizes Assumption 5.
-Assumption 14.
+The following assumption generalizes Assumption 2.5.
+Assumption 4.4.
 $\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely and uniformly over the support of the distance measure vector.
 Theorem 4.2.
-Under Assumptions 6-14, with
+Under Assumptions 3.1-4.4, with
 $\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\theta}^{3}}{n}+n^{\eta-1}d_{\theta}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,$ (4.8)
-as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N(0,1)$.
-(ii) $\mathbb{W}_{3}$ provides a consistent test.
-(iii) Under the sequence of local alternatives $H_{\ell 3}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
-5.  Varying spatial coefficients
+as $n\rightarrow\infty$, the following hold: (1) Under $H_{0}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N(0,1)$.
+(2) $\mathbb{W}_{3}$ provides a consistent test.
+(3) Under the sequence of local alternatives $H_{\ell 3}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
+5 Varying spatial coefficients
 Now we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but allow spatial lag parameters $\lambda_{j}$ to vary.
@@
 where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$.
-Defining $\Lambda_{j}(z)=diag(\lambda_{j}\left(z_{1}\right),\ldots,\\ \lambda_{j}\left(z_{n}\right))$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
+Defining $\Lambda_{j}(z)=diag\left(\lambda_{j}\left(z_{1}\right),\ldots,\lambda_{j}\left(z_{n}\right)\right)$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
 $y=S(z)^{-1}\left(X\beta+t(z)+\epsilon\right),$ (5.3)
@@
 Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$.
-The test statistic is
+Then, the test statistic is
 $\mathbb{W}_{4}=\frac{n\hat{\vartheta}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}-d_{\vartheta}}{\sqrt{2d_{\vartheta}}},$ (5.6)
 where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$. Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
-Assumption 15.
+Assumption 5.1.
 $\sup_{z\in\mathcal{Z}}\mathrm{E}\left(g^{2}_{ri}(z)\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
-Assumption 16.
+Assumption 5.2.
 $\sup_{i\geq 1}\mathrm{E}\left(r_{i\lambda}^{2}\right)=o(n^{-1})$ and $\sup_{i\geq 1}\mathrm{E}\left(r_{i\delta}^{2}\right)=o(n^{-1})$.
-Assumption 17.
+Assumption 5.3.
 $n^{-1}K_{3}^{\prime}K_{3}$ and $n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1, the latter uniformly in $z\in\mathcal{Z}$.
 Lemma 5.1.
-Let Assumptions 6-9, and 15 -17 hold, and
+Let Assumptions 3.1-3.4, and 5.1 -5.3 hold, and
 $\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (5.7)
@@
 Theorem 5.1.
-Let (5.7) hold. Then, under $H_{0}$, Assumptions 6-9, 15-17, as $n\rightarrow\infty$,
+Let (5.7) hold. Then, under $H_{0}$, Assumptions 3.1-3.4, 5.1-5.3, as $n\rightarrow\infty$,
 $\mathbb{W}_{4}-\frac{v^{\prime}\mathcal{M}_{4}v-d_{\vartheta}}{\sqrt{2d_{\vartheta}}}=o_{p}(1).$
-The following assumption is the appropriate version of Assumptions 5 and 14 for the model considered in this section.
-Assumption 18.
+The following assumption is the appropriate version of Assumptions 2.5 and 4.4 for the model considered in this section.
+Assumption 5.4.
 $\left(I_{n}-S(z)\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $z$.
@@
 Theorem 5.2.
-Under Assumptions 6-10 and 15-18, suppose that
+Under Assumptions 3.1-3.5 and 5.1-5.4, suppose that
 $\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\gamma}^{3}}{n}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,$ (5.8)
@@
 (iii) Under the sequence of local alternatives $H_{\ell 4}$, $\mathbb{W}_{4}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
-6.  Empirical applications
+6 Empirical applications
 6.1 Testing for the CRS in a production function
@@
 $y_{i}=\delta_{0}(z_{i})+\delta_{1}(z_{i})p_{i1}+\delta_{2}(z_{i})p_{i2}+\epsilon_{i}.$
-Based on the CRS, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
+Based on the CRS hypothesis, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
 $y_{i}^{*}=\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i},$
 where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate the following model with no SAR structure:
-$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
+$\begin{split}&y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
 where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for k=1,,h/2k=1,...,h/2.
@@
 We define the series approximation based Wald statistic derived from this model as $\mathbb{W}_{0}$.
-Next, we incorporate the SAR structure and consider:
-$y_{i}^{*}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\beta_{0}+\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i}.$
-We estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
-$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
-with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$.
+Next, we estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
+$y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i}.$ (6.2)
+We compute the Wald statistics as follows: $\mathbb{W}_{1}$ from (6.2) with $i.i.d$ errors $\epsilon_{i}$; $\mathbb{W}_{2}$ from (6.2) with $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j};$ $\mathbb{W}_{3}$ from (6.2) replacing $\lambda\sum_{j=1}^{n}w_{ij}y_{j}$ by $\sum_{l=1}^{d_{\tau}}\tau_{l}c_{il}$ and setting $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j};$ $\mathbb{W}_{4}$ from (6.2) with $\lambda\sum_{j=1}^{n}w_{ij}y_{j}$ replaced by $\sum_{m=1}^{d_{\mu}}\mu_{m}\mathfrak{H}_{im}$ and $\varepsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j}$, where $\mathfrak{H}_{im}$ is the $(i,m)$-th entry of the $n\times d_{\mu}$ matrix $H(z)$ defined in Section 5.
+The coefficient on $p_{i1}$ remains the measure of returns to scale technology even with a SAR feature, although this notion differs from the overall returns to scale that combines technological and other spillovers effects.
 Let $d_{ij,m}^{\ast}$ be the geographical distance between firms $i$ and $j$, and $d_{m}$ represent the 10th percentile of all geographical distances.
-We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
-$\begin{split}w_{ij}^{p}=\begin{cases}1\quad\text{if}\ i\ \text{and}\ j\ \text{are in the same city},\\ 0\quad\text{otherwise}.\end{cases}w_{ij}^{d}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise}.\end{cases}\end{split}$
-For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$.
-For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
+We use two different row-normalized spatial weight matrices, denoted
+$\begin{split}w_{ij}^{p}=[W_{n}^{p}]_{ij}=\begin{cases}1\quad\text{if}\ i\ \text{and}\ j\ \text{ in same city},\\ 0\quad\text{otherwise}.\end{cases}w_{ij}^{d}=[W_{n}^{d}]_{ij}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise}.\end{cases}\end{split}$
+For the test statistics, $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$.
+For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{d_{\mu}})^{\prime}$.
 Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for k=1,,hk=1,...,h, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function k=1,,h/2k=1,...,h/2 for $h=2,4$.
 To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function.
-For $\mathbb{W}_{3}$, let $\mathcal{E}_{l}$ be the matrix whose typical entry is, for $i\neq j$ and l=1,,hl=1,...,h, $e_{l}(d_{ij})=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$.
-Then, $c_{il}=\sum_{j=1}^{n}e_{l}(d_{ij})y_{j}$ for l=1,,hl=1,...,h.
-For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=\frac{1}{h}[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\frac{1}{h}\sin(\frac{z_{i}}{2l})$ as the trigonometric function for l=1,,hl=1,...,h.
-Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$.
-For the IVs, we use $K_{1}=K_{3}=[X,\Psi,W_{n}^{p}X,W_{n}^{d}X]$ with respect to $\mathbb{W}_{1}$, $\mathbb{W}_{2}$ and $\mathbb{W}_{4}$, and $K_{2}=[X,\Psi,\mathcal{E}_{1}X,...,\mathcal{E}_{h}X]$ with respect to $\mathbb{W}_{3}$.
-Results with different basis functions are in Table 1.
+For $\mathbb{W}_{3}$, let $\mathcal{E}_{l}$ be the matrix whose typical entry is, for $i\neq j$ and $l=1,...,d_{\tau}$, $e_{l}(d_{ij})=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$.
+Then, $c_{il}=\sum_{j=1}^{n}e_{l}(d_{ij})y_{j}$.
+For $\mathbb{W}_{4}$, we use $\phi_{im}(z)=\frac{1}{d_{\mu}}[\frac{2}{\pi}\tanh(z_{i})]^{m}$ as the polynomial function and $\phi_{im}(z)=\frac{1}{d_{\mu}}\sin(\frac{z_{i}}{2m})$ as the trigonometric function for $m=1,...,d_{\mu}$.
+Then, $\mathfrak{H}_{im}=(\sum_{j=1}^{n}w_{ij}y_{j})\phi_{im}(z)$, where $w_{ij}$ denotes the $(i,j)$-th entry of either $W_{n}^{p}$ or $W_{n}^{d}$.
+For the IVs, set $K_{1}=K_{3}=[X,\Psi,W_{n}^{p}X]$ when the specification uses $W_{n}^{p}$, and $K_{1}=K_{3}=[X,\Psi,W_{n}^{d}X]$ when it uses $W_{n}^{d}$, for $\mathbb{W}_{1}$, $\mathbb{W}_{2}$, and $\mathbb{W}_{4}$, and $K_{2}=[X,\Psi,\mathcal{E}_{1}X,...,\mathcal{E}_{d_{\tau}}X]$ with respect to $\mathbb{W}_{3}$.
+We set $d_{\mu}=d_{\tau}=h$, and the results with different basis functions are in Table 2.
 In all models, the hypothesis of CRS cannot be rejected.
@@
 (2002).
-A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
+These non-rejection results are economically meaningful.
+One can conclude that the CRS technology is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
 We end this application with some remarks: (1) In addition to inference on CRS, our analysis also reveals other similarities with prior works.
 Table 1 shows that the parametric method underestimates the returns to scale compared to the semiparametric method.
-[^1] The returns to scale in a semiparametric model is computed by $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$.
-The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses.
-Figures 1 (a)-(b) illustrates this with the semiparametric estimates with spatial weights $W_{n}^{d}$.
+[^1] The returns to scale in a semiparametric model are computed by summing $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$.The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses.
+Figures S1 (a)-(b) in the online appendix illustrate this with the semiparametric estimates with spatial weights $W_{n}^{d}$.
 They show that the output elasticity of capital is increasing in management expenses, while it is decreasing for labor.
 The decreasing elasticity of labor is a known empirical fact in this industry due to “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest.
-(2) The returns to scale are relatively flat in the semiparametric model and are within the 95% bounds for most $z_{i}$, as illustrated in Figure 1 (c).
+(2) The returns to scale are relatively flat in the semiparametric model and are within the 95% bounds for most $z_{i}$, as illustrated in Figure S1 (c) in the online appendix.
 Further discussions on these findings can be found in Li et al.
@@
 Other choices of spatial weights give the same conclusion.
+Table 1: Comparison of parametric method and semiparametric method in Section 6.1.
+Without $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
+parametric $\hat{\delta}_{1}$ 0.5536 0.5438 0.5330 0.5989 0.5996 0.5960
+$\hat{\delta}_{2}$ 0.4022 0.3794 0.3725 0.3145 0.3226 0.3099
+$\hat{\delta}_{1}$$\hat{\delta}_{2}$ 0.9558 0.9232 0.9055 0.9134 0.9222 0.9060
+polynomial, $h=2$ $\hat{\delta}_{1}$ 0.6228 0.6488 0.6179 0.6263 0.6051 0.6108
+$\hat{\delta}_{2}$ 0.3664 0.3212 0.3691 0.2930 0.3502 0.3220
+$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9892 0.9699 0.9869 0.9194 0.9553 0.9328
+trigonometric, $h=2$ $\hat{\delta}_{1}$ 0.6233 0.6257 0.6149 0.6284 0.6298 0.6806
+$\hat{\delta}_{2}$ 0.3517 0.3099 0.3381 0.3040 0.3377 0.2816
+$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9750 0.9356 0.9530 0.9324 0.9675 0.9621
+polynomial, $h=4$ $\hat{\delta}_{1}$ 0.6840 0.6390 0.5432 0.6077 0.6213 0.6342
+$\hat{\delta}_{2}$ 0.3089 0.3395 0.4345 0.3252 0.3471 0.3025
+$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9929 0.9785 0.9777 0.9330 0.9683 0.9368
+trigonometric, $h=4$ $\hat{\delta}_{1}$ 0.6792 0.6896 0.6814 0.6152 0.6289 0.6569
+$\hat{\delta}_{2}$ 0.3121 0.2702 0.3002 0.3043 0.3649 0.3199
+$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9913 0.9598 0.9816 0.9195 0.9938 0.9768
 6.2 Testing the impact of the distance on house prices
@@
 $y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
-Therefore, we estimate:
-$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{4}.\end{split}$
+Let $d_{\tau}=d_{\mu}=h$, we estimate:
+$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}\mathfrak{H}_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{4}.\end{split}$
 We consider two row-normalized spatial weight matrices.
@@
 As shown in Panel E of Table 3, the parametric method underestimates the spatial coefficient.
-Thus our results based on more general models complement those in Sun et al.
+Thus, our results based on more general models complement those in Sun et al.
 (2014).
-7.  Monte Carlo
+7 Monte Carlo Simulation
 We present a basic set of Monte Carlo results in this section and a wider range of simulations in the online appendix. All experiments use 1000 replications.
@@
 To construct the SHAC statistics $\mathbb{W}_{2},\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we proceed as follows.
-Generate $M=d_{\lambda}$ distinct distance measures by setting $\ell=[n^{\eta}]+1$ with $\eta=3/7$ and taking $q=8$, so that each unit has at least $\ell$ neighbors, where $\eta$ is defined in Assumption 9 (a).
+Generate $M=d_{\lambda}$ distinct distance measures by setting $\ell=[n^{\eta}]+1$ with $\eta=3/7$ and taking $q=8$, so that each unit has at least $\ell$ neighbors, where $\eta$ is defined in Assumption 3.4 (a).
 [^4] From the symmetric circulant matrix $W_{d_{\lambda}}$, we build a graph and compute its pairwise shortest‐path distances in MATLAB, then apply cmdscale to generate $n\times 2$ coordinate matrix.
@@
 We set $\delta=0$ for the null hypothesis and $\delta(\mathrm{x})=1-\mathrm{x}^{2}$ for any $\mathrm{x}\in R$ for the alternatives.
-We choose polynomials $\psi_{j}(\mathrm{x})=\mathrm{x}^{j}$ for $j=1,\cdots,h$ and the design of trigonometric functions is in the supplement.
+We choose polynomials $\psi_{j}(\mathrm{x})=\mathrm{x}^{j}$ for $j=1,\cdots,h$ and the design of trigonometric functions is in the online appendix.
 The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[X,W_{1}X,...,W_{d_{\lambda}}X,\Psi]$.
@@
 This may also help in empirical applications, although in both applications in the previous section critical values from either distribution yield the same results.
-We discuss size performance in this section and present power performance in the supplement, since under global alternatives the power of the asymptotic tests tends to one as $n\to\infty$.
+We discuss size performance in this section and present power performance in the online appendix, since under global alternatives the power of the asymptotic tests tends to one as $n\to\infty$.
 Table 4 reports the empirical sizes of $\mathbb{W}_{1}$, and Table 5 reports those of the SHAC-corrected statistic $\mathbb{W}_{2}$, at nominal 1%, 5% and 10% levels.
@@
 Results for trigonometric functions are similar, and left for the online appendix.
-8.  Conclusion
+8 Conclusion
 We provide a machinery for conducting tests on nonparametrically varying coefficients using ideas from parametric testing, and with similar ease of implementation.
@@
 These provide directions for future research.
-Appendix
+A Supplement to
+Wald inference on varying coefficients
+Abhimanyu Gupta
+Department of Economics, Queen's University, Dunning Hall, 94 University Avenue, Kingston, Ontario K7L 3N6, Canada and Department of Economics, University of Essex, Wivenhoe Park, Colchester, CO4 3SQ, UK.
+Email: abhimanyu.g@queensu.ca.
+Xi Qu
+Department of Economics, Antai College of Economics and Management, Shanghai Jiao Tong University, 1954 Huashan Road, Shanghai, 200030, China PRC. Email: xiqu@sjtu.edu.cn.
+Sorawoot Srisuma
+Department of Economics, National University of Singapore, 1 Arts Link, 117570, Singapore.
+Jiajun Zhang
+International Business School, Shanghai University of International Business and Economics, 201620, China PRC. E-mail: jiajun30@suibe.edu.cn.
 Appendix A General Central Limit Theorem
@@
 $\mathscr{A}=B^{\prime}\mathscr{S}B,$ (A.1)
-where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7 and $\mathscr{S}$ is a symmetric matrix.
+where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 3.2 and $\mathscr{S}$ is a symmetric matrix.
 (b) The entries $s_{ij}$ of $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$.
@@
 Theorem A.1.
-Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 10. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
+Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 3.5. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
 ${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}\overset{d}{\longrightarrow}N(0,1).$ (A.2)
-Proof.
-${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}=\sum_{s=1}^{n}\rho_{s}$, where
+(Proof). ${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}=\sum_{s=1}^{n}\rho_{s}$, where
 $\rho_{s}=L^{-1}a_{ss}\left(v_{s}^{2}-1\right)+2L^{-1}\mathbf{1}(s\geq 2)v_{s}\sum_{t<s}a_{st}v_{t},$
@@
 $\sum_{t=1}^{n-1}\sum_{r=1}^{t-1}\sum_{s=t+1}^{n}\sum_{u=t+1}^{n}b_{s}^{\prime}\mathscr{S}b_{t}b_{s}^{\prime}\mathscr{S}b_{r}b_{u}^{\prime}\mathscr{S}b_{t}b_{u}^{\prime}\mathscr{S}b_{r}$
-$=$ $O_{p}\left(L^{-4}\sum_{s,r,t,u=1}^{n}\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{ju}\right|\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\right$.
-$\times\left.\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{ju}\right|\right)$
+$=$ $O_{p}\left(L^{-4}\sum_{s,r,t,u=1}^{n}\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{ju}\right|\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{js}\right|\times\sum_{i,j=1}^{n}\left|b_{it}\right|\left|s_{ij}\right|\left|b_{ju}\right|\right)$
 $=$ $O_{p}\left(L^{-4}pn^{-1}\sum_{s,r,t=1}^{n}\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\left|b_{js}\right|\right)\left(\sum_{i,j=1}^{n}\left|b_{ir}\right|\left|s_{ij}\right|\sum_{u=1}^{n}\left|b_{ju}\right|\right)\right$.
@@
 $=$ $O_{p}\left((LJN)^{-4}\sum_{t=1}^{n-1}\left(\sum_{s=t+1}^{n}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right)^{2}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{4}\right)$
-$=$ $O_{p}\left((LJN)^{-4}\left(\sum_{t=1}^{n-1}\sum_{j=1}^{n}\left|b_{jt}\right|\right)\left(\sum_{s=t+1}^{n}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\right$.
-$\times\left.\sup_{s}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\sup_{t}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{3}\right)$
+$=$ $O_{p}\left((LJN)^{-4}\left(\sum_{t=1}^{n-1}\sum_{j=1}^{n}\left|b_{jt}\right|\right)\left(\sum_{s=t+1}^{n}\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\times\sup_{s}\left(\sum_{i=1}^{n}\left|b_{is}\right|\right)^{2}\sup_{t}\left(\sum_{j=1}^{n}\left|b_{jt}\right|\right)^{3}\right)$
 $=$ $O_{p}\left((L^{-4}J^{4}n^{-1}\right)=O_{p}\left(J^{2}n^{-1}\right).$
+$\blacksquare$
 Appendix B Proofs of Theorems
-Proof of Theorem 2.1.
-Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$.
-Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}\\ P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$.
+(Proof).[Proof of Theorem 2.1] Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$.
+Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$.
 Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$.
@@
 where $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$ with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, and $A_{1}=r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$ and $A_{2}=2r_{\delta}^{\prime}\mathcal{M}_{1}\epsilon$.
-Now, $|A_{1}|\leq\|r_{\delta}\|^{2}\|\mathcal{M}_{1}\|=o_{p}(1)$, since $E\|r_{\delta}\|^{2}=o(1)$, and $\|\mathcal{M}_{1}\|=O_{p}(1)$ by Assumptions 1 and 4.
+Now, $|A_{1}|\leq\|r_{\delta}\|^{2}\|\mathcal{M}_{1}\|=o_{p}(1)$, since $E\|r_{\delta}\|^{2}=o(1)$, and $\|\mathcal{M}_{1}\|=O_{p}(1)$ by Assumptions 2.1 and 2.4.
 A similar argument holds for $A_{2}$ because it has (conditionally) zero mean and variance $4r_{\delta}^{\prime}\mathcal{M}_{1}E(\epsilon\epsilon^{\prime})\mathcal{M}_{1}r_{\delta}=4r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$, by idempotence of $\mathcal{M}_{1}$.
 Thus $\frac{A_{1}}{\sqrt{d_{\alpha}}}$ and $\frac{A_{2}}{\sqrt{d_{\alpha}}}$ are both $o_{p}(1)$ and establishing the theorem.
-∎
-Proof of Theorem $2.$
-This is a special case of Theorem 3.2 and therefore omitted. ∎
-Proof of Lemma 3.1.
-This is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
-Proof of Theorem 3.1.
-$\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(L^{\prime}{P_{K_{1}}}L)^{-1}L^{\prime}{P_{K_{1}}}(r_{\delta}+Bv)$ and under $H_{0}:\alpha=0$,
+$\blacksquare$
+(Proof).[Proof of Theorem 2.2] This is a special case of Theorem 3.2 and therefore omitted. $\blacksquare$
+(Proof).[Proof of Lemma 3.1] This is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. $\blacksquare$
+(Proof).[Proof of Theorem 3.1] $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(L^{\prime}{P_{K_{1}}}L)^{-1}L^{\prime}{P_{K_{1}}}(r_{\delta}+Bv)$ and under $H_{0}:\alpha=0$,
 $\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{2}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{2}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.2)
@@
 The remainder of the proof is similar to the proof of Theorem 5.1.
-∎
-Proof of Theorem 3.2.
-The proof follows by setting $J=J_{1}$, $\mathscr{A}=\mathcal{M}_{2}$ and $\mathscr{S}=\frac{1}{n}K_{1}\mathcal{V}_{2}K_{1}^{\prime}$ in Theorem A.1.
+$\blacksquare$
+(Proof).[Proof of Theorem 3.2] The proof follows by setting $J=J_{1}$, $\mathscr{A}=\mathcal{M}_{2}$ and $\mathscr{S}=\frac{1}{n}K_{1}\mathcal{V}_{2}K_{1}^{\prime}$ in Theorem A.1.
 Checking Assumption A follows exactly as in the proof of Theorem 5.2 below and is omitted.
-∎
-Proof of Lemma 4.1.
-This proof is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
-Proof of Theorem 4.1.
-Define $r_{G}=(r_{1G},\cdots,r_{nG})^{\prime}$ and $r_{G\delta}=r_{G}+r_{\delta}$.
-Then, we have $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(F^{\prime}{P_{K_{2}}}F)^{-1}F^{\prime}{P_{K_{2}}}(r_{G\delta}+Bv)$ and $\big\|r_{G\delta}\big\|^{2}=o_{p}(1)$ by Assumption 12.
+$\blacksquare$
+(Proof).[Proof of Lemma 4.1] This proof is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. $\blacksquare$
+(Proof).[Proof of Theorem 4.1] Define $r_{G}=(r_{1G},\cdots,r_{nG})^{\prime}$ and $r_{G\delta}=r_{G}+r_{\delta}$.
+Then, we have $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(F^{\prime}{P_{K_{2}}}F)^{-1}F^{\prime}{P_{K_{2}}}(r_{G\delta}+Bv)$ and $\big\|r_{G\delta}\big\|^{2}=o_{p}(1)$ by Assumption 4.2.
 Under $H_{0}:\alpha=0$,
 $\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{3}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{3}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.3)
-where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, $\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$, $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and $\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$.
+where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, $\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$ with $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and $\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$.
 The remainder of the proof is similar to the proof of Theorem 5.1.
-∎
-Proof of Theorem 4.2.
-The proof is similar to that of Theorem 5.2 and omitted. ∎
-Proof of Lemma 5.1.
-Define $r_{\lambda}=(r_{1\lambda},...,r_{n\lambda})^{\prime}$. First, note that
+$\blacksquare$
+(Proof).[Proof of Theorem 4.2] The proof is similar to that of Theorem 5.2 and omitted. $\blacksquare$
+(Proof).[Proof of Lemma 5.1] Define $r_{\lambda}=(r_{1\lambda},...,r_{n\lambda})^{\prime}$. First, note that
 $\hat{\gamma}-\gamma=(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}u=(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta}+\epsilon),$ (B.4)
@@
 $\begin{split}\mathrm{E}\left\|n^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta})\right\|^{2}&\leq n^{-2}\left\|K_{3}\right\|^{2}\mathrm{E}\left\|r_{\lambda}+r_{\delta}\right\|^{2}\\ &\leq Cn^{-1}\overline{\alpha}(n^{-1}K_{3}^{\prime}K_{3})\left(\mathrm{E}\left\|r_{\lambda}\right\|^{2}+\mathrm{E}\left\|r_{\delta}\right\|^{2}\right)=o(1),\end{split}$ (B.5)
-because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17. Next,
+because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 5.2 and 5.3.
+Next,
 $\mathrm{E}\left\|n^{-1}K_{3}^{\prime}\epsilon\right\|^{2}$ $\leq$ $n^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}\mathrm{E}\epsilon_{i}^{2}+n^{-2}\sum_{i\neq j}\left\|k_{i}\right\|\left\|k_{j}\right\|\left|\mathrm{E}\epsilon_{i}\epsilon_{j}\right|$ (B.6)
@@
 $=$ $O(d_{\gamma}/n),$
-by Assumptions 7 and 15.
-Combining Assumption 17, (B.4), (B.5) and (B.6), we obtain
+by Assumptions 3.2 and 5.1.
+Combining Assumption 5.3, (B.4), (B.5) and (B.6), we obtain
 $\left\|\hat{\gamma}-\gamma\right\|=O_{p}(\sqrt{d_{\gamma}/n}).$ (B.7)
@@
 $\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j}=g_{i}^{\prime}(\hat{\gamma}-\gamma)g_{j}^{\prime}(\hat{\gamma}-\gamma)+g_{i}^{\prime}(\hat{\gamma}-\gamma)r_{j}+r_{i}r_{j}+g_{j}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{i}+g_{i}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{j}+\epsilon_{i}r_{j}+\epsilon_{j}r_{i}.$ (B.11)
-Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$.
+Now, by Assumptions 5.1 and 5.2, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$.
 Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, writing $\pi_{M}=1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})$ and also using (B.7),
@@
 Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$.
-By Assumption 9(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$.
+By Assumption 3.4(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$.
 We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$.
@@
 This concludes the proof.
-∎
-Proof of Theorem 5.1.
-Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$.
-Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big\|r_{\lambda\delta}\big\|^{2}=o_{p}(1)$ by Assumption 16.
+$\blacksquare$
+(Proof).[Proof of Theorem 5.1] Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$.
+Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big\|r_{\lambda\delta}\big\|^{2}=o(1)$ by Assumption 5.2.
 Under $H_{0}:\vartheta=0$,
@@
 where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, $\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$, $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and $\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$.
-Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big\|\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17.
-Thus, $\big\|\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big\|\leq\big\|\widehat{\mathcal{U}}_{3}^{-1}\big\|\big\|\mathcal{U}_{3}-\widehat{\mathcal{U}}_{3}\big\|\big\|\mathcal{U}_{3}^{-1}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$.
-Similarly, we have $\big\|\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big\|=o_{p}(1)$, and $\big\|\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big\|=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
+Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big\|\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 5.1 and 5.3.
+Thus, $\big\|\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big\|\leq\big\|\widehat{\mathcal{U}}_{3}^{-1}\big\|\big\|\mathcal{U}_{3}-\widehat{\mathcal{U}}_{3}\big\|\big\|\mathcal{U}_{3}^{-1}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 3.1 and 5.3, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$.
+Similarly, we have $\big\|\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big\|=o_{p}(1)$, and $\big\|\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big\|=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 3.2 and 5.1.
 Since $\mathrm{E}(v_{i}^{2})=1$, we have
 $\frac{1}{\sqrt{d_{\vartheta}}}\big|v^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v\big|\leq Cr_{n}\cdot O_{p}\left(d_{\vartheta}^{-1/2}\right)=o_{p}(1)$ (B.13)
-Similar to (B.13) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\vartheta}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\vartheta}}}|A_{22}|=o_{p}(1)$.
+Similar to (B.13) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\vartheta}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 3.2 and 5.1, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\vartheta}}}|A_{22}|=o_{p}(1)$.
 Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible.
-∎
-Proof of Theorem 5.2.
-(i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1.
-We now note that the entries of $\mathscr{S}$ are $s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so $\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above.
+$\blacksquare$
+(Proof).[Proof of Theorem 5.2] (i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1.
+We now note that the entries of $\mathscr{S}$ are $s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so $\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i,j$, by Assumptions 5.1 and 5.3, and arguments used in the proof of Theorem 5.1 above.
 Similarly, we also observe that $\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i$.
 Thus Assumption A(b) is satisfied.
-Assumption A is satisfied to due to Assumption 17.
+Assumption A is satisfied to due to Assumption 5.3.
 The conclusion now follows by Theorem A.1.
@@
 $\left\|\hat{\vartheta}-\vartheta\right\|=O_{p}\left(\sqrt{d_{\vartheta}/n}\right).$ (B.14)
-From Assumptions 15 and 17, we have $\left\|\widehat{\mathscr{D}}_{4}-\mathscr{D}_{4}\right\|=o_{p}(1)$ and
+From Assumptions 5.1 and 5.3, we have $\left\|\widehat{\mathscr{D}}_{4}-\mathscr{D}_{4}\right\|=o_{p}(1)$ and
 $\left\|\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1}\right\|\leq\left\|\widehat{\mathscr{D}}_{4}^{-1}\right\|\left\|\mathscr{D}_{4}-\widehat{\mathscr{D}}_{4}\right\|\left\|\mathscr{D}_{4}^{-1}\right\|=o_{p}(1).$ (B.15)
@@
 We choose $\Gamma_{4n}=\widehat{\mathscr{D}}_{4}^{-1}$ and then finish the proof.
-The proofs of $\mathbb{W}_{2}^{*}$ and $\mathbb{W}_{3}^{*}$ are similar.
-∎
-Appendix C Empirical Results
-Table 1: Estimation and testing results in Section 6.1.
-Panel A: polynomial series, $h=2$
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.189* 0.130** 0.224*** 0.256*** 0.451*** 0.457*** 0.499***
-(0.112) (0.062) (0.014) (0.081) (0.157) (0.042) (0.028)
-$\tau_{2}/\mu_{2}$ -0.095 0.405*** 0.484***
-(0.126) (0.047) (0.030)
-$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha_{3}=0$ 1.593 3.311 3.280 1.419 4.020 0.226 1.716 2.589
-$H_{0}:\mu=0$ 68.365*** 92.106***
-Panel B: trigonometric series, $h=2$
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.065 0.162* 0.065** 0.162*** 0.582*** 0.607*** 0.297***
-(0.052) (0.087) (0.033) (0.014) (0.098) (0.166) (0.088)
-$\tau_{2}/\mu_{2}$ -0.276*** 0.147** 0.494***
-(0.041) (0.074) (0.088)
-$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha_{3}=0$ 0.777 1.178 0.877 1.086 0.943 1.575 1.707 2.169
-$H_{0}:\mu=0$ 321.305*** 406.511***
-Panel C: polynomial series, $h=4$
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.330*** 0.265** 0.330*** 0.065*** 0.444*** 0.236*** 0.313***
-(0.095) (0.134) (0.019) (0.024) (0.135) (0.043) (0.0969)
-$\tau_{2}/\mu_{2}$ -0.013 0.219*** 0.330***
-(0.045) (0.052) (0.064)
-$\tau_{3}/\mu_{3}$ 0.025** 0.203** 0.258**
-(0.012) (0.082) (0.126)
-$\tau_{4}/\mu_{4}$ -0.088 0.189* 0.096*
-(0.068) (0.110) (0.051)
-$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha_{3}=0$ 0.075 1.707 1.464 2.401 1.808 0.030 0.098 0.385
-$H_{0}:\mu=0$ 101.231*** 122.398***
-Panel D: trigonometric series, $h=4$
-Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.224*** 0.256*** 0.224*** 0.256*** 0.196** 0.123* 0.136***
-(0.040) (0.083) (0.014) (0.081) (0.098) (0.067) (0.033)
-$\tau_{2}/\mu_{2}$ 0.261* 0.298** 0.269**
-(0.136) (0.143) (0.110)
-$\tau_{3}/\mu_{3}$ 0.029* 0.293** 0.299**
-(0.018) (0.145) (0.140)
-$\tau_{4}/\mu_{4}$ -0.011 0.107* 0.138*
-(0.124) 0.026** 0.035***
-$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha_{3}=0$ 0.290 1.848 1.497 0.038 1.443 0.034 0.375 1.022
-$H_{0}:\mu=0$ 86.301*** 90.233***
-Standard errors in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Table 2: Comparison of parametric method and semiparametric method in Section 6.1.
-Without $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
-parametric $\hat{\delta}_{1}$ 0.5536 0.5438 0.5330 0.5989 0.5996 0.5960
-$\hat{\delta}_{2}$ 0.4022 0.3794 0.3725 0.3145 0.3226 0.3099
-$\hat{\delta}_{1}$$\hat{\delta}_{2}$ 0.9558 0.9232 0.9055 0.9134 0.9222 0.9060
-polynomial, $h=2$ $\hat{\delta}_{1}$ 0.6228 0.6488 0.6179 0.6263 0.6051 0.6108
-$\hat{\delta}_{2}$ 0.3664 0.3212 0.3691 0.2930 0.3502 0.3220
-$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9892 0.9699 0.9869 0.9194 0.9553 0.9328
-trigonometric, $h=2$ $\hat{\delta}_{1}$ 0.6233 0.6257 0.6149 0.6284 0.6298 0.6806
-$\hat{\delta}_{2}$ 0.3517 0.3099 0.3381 0.3040 0.3377 0.2816
-$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9750 0.9356 0.9530 0.9324 0.9675 0.9621
-polynomial, $h=4$ $\hat{\delta}_{1}$ 0.6840 0.6390 0.5432 0.6077 0.6213 0.6342
-$\hat{\delta}_{2}$ 0.3089 0.3395 0.4345 0.3252 0.3471 0.3025
-$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9929 0.9785 0.9777 0.9330 0.9683 0.9368
-trigonometric, $h=4$ $\hat{\delta}_{1}$ 0.6792 0.6896 0.6814 0.6152 0.6289 0.6569
-$\hat{\delta}_{2}$ 0.3121 0.2702 0.3002 0.3043 0.3649 0.3199
-$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9913 0.9598 0.9816 0.9195 0.9938 0.9768
+The proofs of $\mathbb{W}_{2}^{*}$ and $\mathbb{W}_{3}^{*}$ $\blacksquare$
+Appendix C Supplementary materials
+This section presents additional simulation results.
+Additional results for the polynomial functions in the main text are shown in Tables S1-S5, and those for the trigonometric functions in the main text are shown in Tables S6-S14.
+We also considered tests of the CRS hypothesis to facilitate the empirical part in Section 6.1, with $d_{\delta}=2$.
+Our null of interest when testing CRS is $H_{0}^{true}:\delta_{1}(z)+\delta_{2}(z)=1,z\in\mathcal{Z}.$ We reparameterized with $\theta(z)=\delta_{1}(z)+\delta_{2}(z)-1$ and the model is $y_{i}-p_{i1}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i2}-p_{i1})\delta_{2}(z_{i})+\epsilon_{i}.$ For the alternatives, we set $\delta_{1}(\mathrm{x})=1-\mathrm{x}^{2}$ and $\delta_{2}(\mathrm{x})=1+\mathrm{x}^{2}$.
+We studied all four statistics proposed in the main text.
+These results are similar to our main results, and so the corresponding tables are omitted and available upon request.
+Table S1: Empirical sizes of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\mu=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Polynomial basis functions.
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.081 0.109 0.126 0.044 0.072 0.102 0.018 0.044 0.063
+chi-p 0.050 0.100 0.122 0.012 0.063 0.097 0.005 0.029 0.059
+$h=4$ asy-p 0.100 0.146 0.171 0.052 0.081 0.103 0.011 0.027 0.041
+chi-p 0.072 0.128 0.169 0.039 0.068 0.100 0.005 0.039 0.091
+$h=8$ asy-p 0.129 0.162 0.198 0.087 0.126 0.159 0.058 0.085 0.122
+chi-p 0.100 0.150 0.190 0.068 0.116 0.153 0.020 0.066 0.106
+7.1 $h=2$ asy-p 0.082 0.117 0.143 0.057 0.104 0.137 0.021 0.046 0.068
+chi-p 0.057 0.093 0.140 0.025 0.077 0.130 0.004 0.032 0.063
+$h=4$ asy-p 0.127 0.156 0.188 0.060 0.085 0.114 0.007 0.022 0.032
+chi-p 0.085 0.141 0.185 0.037 0.075 0.111 0.004 0.013 0.031
+$h=8$ asy-p 0.118 0.151 0.174 0.109 0.145 0.168 0.063 0.096 0.117
+chi-p 0.089 0.140 0.166 0.090 0.133 0.166 0.019 0.070 0.114
+7.1 $h=2$ asy-p 0.094 0.127 0.151 0.052 0.090 0.126 0.021 0.040 0.071
+chi-p 0.051 0.110 0.146 0.027 0.071 0.122 0.007 0.039 0.086
+$h=4$ asy-p 0.117 0.168 0.200 0.068 0.107 0.141 0.019 0.030 0.043
+chi-p 0.076 0.145 0.194 0.048 0.084 0.137 0.008 0.027 0.081
+$h=8$ asy-p 0.128 0.176 0.197 0.079 0.126 0.151 0.068 0.093 0.118
+chi-p 0.103 0.155 0.194 0.054 0.112 0.147 0.021 0.062 0.114
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.115 0.176 0.209 0.044 0.072 0.102 0.018 0.044 0.085
+chi-p 0.072 0.149 0.204 0.012 0.063 0.097 0.005 0.046 0.088
+$h=4$ asy-p 0.166 0.233 0.272 0.052 0.081 0.103 0.011 0.042 0.081
+chi-p 0.119 0.208 0.263 0.039 0.068 0.100 0.005 0.041 0.086
+$h=8$ asy-p 0.219 0.284 0.319 0.087 0.126 0.119 0.028 0.060 0.112
+chi-p 0.194 0.261 0.314 0.068 0.116 0.113 0.018 0.058 0.108
+7.1 $h=2$ asy-p 0.126 0.180 0.224 0.057 0.104 0.114 0.012 0.046 0.096
+chi-p 0.081 0.151 0.217 0.025 0.077 0.111 0.004 0.032 0.093
+$h=4$ asy-p 0.184 0.245 0.293 0.060 0.085 0.117 0.007 0.063 0.088
+chi-p 0.149 0.224 0.285 0.037 0.075 0.112 0.004 0.049 0.086
+$h=8$ asy-p 0.224 0.284 0.313 0.109 0.145 0.121 0.022 0.062 0.117
+chi-p 0.192 0.263 0.310 0.090 0.133 0.116 0.013 0.042 0.114
+7.1 $h=2$ asy-p 0.112 0.155 0.202 0.052 0.090 0.122 0.021 0.040 0.092
+chi-p 0.068 0.135 0.194 0.027 0.071 0.114 0.007 0.029 0.067
+$h=4$ asy-p 0.192 0.254 0.301 0.068 0.107 0.132 0.019 0.030 0.066
+chi-p 0.145 0.230 0.294 0.048 0.084 0.124 0.008 0.027 0.064
+$h=8$ asy-p 0.193 0.249 0.287 0.079 0.126 0.127 0.042 0.072 0.113
+chi-p 0.159 0.230 0.280 0.054 0.112 0.120 0.040 0.082 0.110
+Table S2: Empirical powers of $\mathbb{W}_{2}$ for the baseline SAR model with SHAC correction proposed in Section 3.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Polynomial basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.979 0.989 0.991 0.996 0.997 0.997 1 1 1
+chi-p 0.969 0.984 0.991 0.995 0.996 0.997 0.999 1 1
+$h=4$ asy-p 0.997 0.999 0.999 0.999 0.999 0.999 1 1 1
+chi-p 0.995 0.997 0.999 0.999 0.999 0.999 1 1 1
+$h=8$ asy-p 0.920 0.921 0.923 0.987 0.987 0.987 1 1 1
+chi-p 0.920 0.920 0.923 0.987 0.987 0.987 1 1 1
+7.1 $h=2$ asy-p 0.980 0.987 0.990 0.993 0.999 1 1 1 1
+chi-p 0.962 0.985 0.990 0.987 0.996 1 0.999 1 1
+$h=4$ asy-p 0.998 0.998 0.998 0.999 0.999 1 0.999 1 1
+chi-p 0.996 0.998 0.998 0.998 0.999 1 0.999 0.999 1
+$h=8$ asy-p 0.937 0.938 0.939 0.990 0.990 0.990 0.998 0.998 0.998
+chi-p 0.934 0.938 0.939 0.989 0.990 0.990 0.998 0.998 0.998
+7.1 $h=2$ asy-p 0.977 0.984 0.986 0.997 0.999 0.999 0.997 0.997 0.998
+chi-p 0.963 0.979 0.986 0.995 0.998 0.999 0.997 0.997 0.998
+$h=4$ asy-p 0.998 0.998 0.998 0.999 1 1 1 1 1
+chi-p 0.991 0.998 0.998 0.999 1 1 1 1 1
+$h=8$ asy-p 0.908 0.912 0.913 0.991 0.992 0.992 0.999 0.999 0.999
+chi-p 0.907 0.909 0.913 0.991 0.991 0.992 0.999 0.999 0.999
+$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.993 0.994 0.997 1 1 1 1 1 1
+chi-p 0.986 0.993 0.997 1 1 1 1 1 1
+$h=4$ asy-p 0.999 1 1 1 1 1 1 1 1
+chi-p 0.999 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.987 0.988 0.988 0.998 0.998 0.998 0.998 0.998 0.998
+chi-p 0.986 0.988 0.988 0.998 0.998 0.998 0.998 0.998 0.998
+7.1 $h=2$ asy-p 0.994 0.997 0.998 0.997 0.998 0.998 1 1 1
+chi-p 0.985 0.996 0.998 0.996 0.997 0.998 1 1 1
+$h=4$ asy-p 0.999 1 1 0.999 1 1 1 1 1
+chi-p 0.999 1 1 0.999 0.999 1 1 1 1
+$h=8$ asy-p 0.984 0.984 0.984 0.995 0.995 0.995 0.999 0.999 0.999
+chi-p 0.984 0.984 0.984 0.995 0.995 0.995 0.999 0.999 0.999
+7.1 $h=2$ asy-p 0.997 0.998 0.998 0.996 0.999 0.999 1 1 1
+chi-p 0.995 0.997 0.998 0.996 0.999 0.999 1 1 1
+$h=4$ asy-p 1 1 1 1 1 1 1 1 1
+chi-p 1 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.977 0.978 0.978 0.999 1 1 0.998 0.998 0.998
+chi-p 0.977 0.978 0.978 0.998 1 1 0.998 0.998 0.998
+Table S3: Empirical powers of $\mathbb{W}_{3}$ for the SAR model with nonparametric spatial weights proposed in Section 4.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Polynomial basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\tau}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.979 0.989 0.991 0.996 0.997 0.997 1 1 1
+chi-p 0.969 0.984 0.991 0.995 0.996 0.997 0.999 1 1
+$h=4$ asy-p 0.997 0.999 0.999 0.999 0.999 0.999 1 1 1
+chi-p 0.995 0.997 0.999 0.999 0.999 0.999 1 1 1
+$h=8$ asy-p 0.920 0.921 0.923 0.987 0.987 0.987 1 1 1
+chi-p 0.920 0.920 0.923 0.987 0.987 0.987 1 1 1
+7.1 $h=2$ asy-p 0.980 0.987 0.990 0.993 0.999 1 1 1 1
+chi-p 0.962 0.985 0.990 0.987 0.996 1 0.999 1 1
+$h=4$ asy-p 0.998 0.998 0.998 0.999 0.999 1 0.999 1 1
+chi-p 0.996 0.998 0.998 0.998 0.999 1 0.999 0.999 1
+$h=8$ asy-p 0.937 0.938 0.939 0.990 0.990 0.990 0.998 0.998 0.998
+chi-p 0.934 0.938 0.939 0.989 0.990 0.990 0.998 0.998 0.998
+7.1 $h=2$ asy-p 0.977 0.984 0.986 0.997 0.999 0.999 0.997 0.997 0.998
+chi-p 0.963 0.979 0.986 0.995 0.998 0.999 0.997 0.997 0.998
+$h=4$ asy-p 0.998 0.998 0.998 0.999 1 1 1 1 1
+chi-p 0.991 0.998 0.998 0.999 1 1 1 1 1
+$h=8$ asy-p 0.908 0.912 0.913 0.991 0.992 0.992 0.999 0.999 0.999
+chi-p 0.907 0.909 0.913 0.991 0.991 0.992 0.999 0.999 0.999
+$d_{\tau}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.993 0.994 0.997 1 1 1 1 1 1
+chi-p 0.986 0.993 0.997 1 1 1 1 1 1
+$h=4$ asy-p 0.999 1 1 1 1 1 1 1 1
+chi-p 0.999 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.987 0.988 0.988 0.998 0.998 0.998 0.998 0.998 0.998
+chi-p 0.986 0.988 0.988 0.998 0.998 0.998 0.998 0.998 0.998
+7.1 $h=2$ asy-p 0.994 0.997 0.998 0.997 0.998 0.998 1 1 1
+chi-p 0.985 0.996 0.998 0.996 0.997 0.998 1 1 1
+$h=4$ asy-p 0.999 1 1 0.999 1 1 1 1 1
+chi-p 0.999 1 1 0.999 0.999 1 1 1 1
+$h=8$ asy-p 0.984 0.984 0.984 0.995 0.995 0.995 0.999 0.999 0.999
+chi-p 0.984 0.984 0.984 0.995 0.995 0.995 0.999 0.999 0.999
+7.1 $h=2$ asy-p 0.997 0.998 0.998 0.996 0.999 0.999 1 1 1
+chi-p 0.995 0.997 0.998 0.996 0.999 0.999 1 1 1
+$h=4$ asy-p 1 1 1 1 1 1 1 1 1
+chi-p 1 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.977 0.978 0.978 0.999 1 1 0.998 0.998 0.998
+chi-p 0.977 0.978 0.978 0.998 1 1 0.998 0.998 0.998
+Table S4: Empirical powers of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Polynomial basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.972 0.977 0.984 0.990 0.991 0.996 0.996 0.997 0.998
+chi-p 0.977 0.980 0.984 0.991 0.993 0.996 0.997 0.997 0.998
+$h=4$ asy-p 0.952 0.969 0.981 0.973 0.983 0.989 0.991 0.997 0.999
+chi-p 0.963 0.976 0.982 0.980 0.986 0.990 0.995 0.998 0.999
+$h=8$ asy-p 0.901 0.924 0.945 0.948 0.962 0.968 0.972 0.980 0.986
+chi-p 0.915 0.931 0.947 0.955 0.966 0.971 0.978 0.982 0.986
+7.1 $h=2$ asy-p 0.975 0.986 0.992 0.990 0.991 0.992 0.990 0.995 0.997
+chi-p 0.983 0.990 0.992 0.991 0.992 0.992 0.995 0.997 0.997
+$h=4$ asy-p 0.951 0.968 0.973 0.977 0.985 0.987 0.989 0.991 0.995
+chi-p 0.966 0.973 0.974 0.983 0.985 0.988 0.991 0.993 0.996
+$h=8$ asy-p 0.912 0.935 0.947 0.946 0.961 0.970 0.973 0.983 0.986
+chi-p 0.924 0.940 0.947 0.957 0.964 0.971 0.979 0.984 0.986
+7.1 $h=2$ asy-p 0.967 0.981 0.984 0.994 0.997 0.998 0.994 0.993 0.996
+chi-p 0.978 0.982 0.984 0.996 0.997 0.998 0.996 0.997 1
+$h=4$ asy-p 0.948 0.968 0.976 0.975 0.985 0.987 0.990 0.997 0.997
+chi-p 0.960 0.972 0.976 0.985 0.985 0.987 0.992 0.993 1
+$h=8$ asy-p 0.896 0.922 0.941 0.953 0.973 0.975 0.978 0.987 1
+chi-p 0.912 0.932 0.944 0.965 0.974 0.976 0.982 0.990 1
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.986 0.993 0.996 0.992 0.996 0.997 0.995 0.998 0.998
+chi-p 0.991 0.993 0.997 0.995 0.997 0.997 0.996 0.998 0.998
+$h=4$ asy-p 0.919 0.954 0.972 0.981 0.989 0.995 0.990 0.993 0.995
+chi-p 0.942 0.963 0.973 0.988 0.992 0.995 0.991 0.993 0.995
+$h=8$ asy-p 0.810 0.875 0.898 0.939 0.960 0.969 0.963 0.969 0.974
+chi-p 0.843 0.886 0.901 0.952 0.964 0.969 0.966 0.971 0.974
+7.1 $h=2$ asy-p 0.970 0.983 0.986 0.996 0.998 0.999 0.996 0.998 0.999
+chi-p 0.980 0.985 0.988 0.996 0.998 0.999 0.998 0.999 0.999
+$h=4$ asy-p 0.916 0.947 0.960 0.976 0.985 0.989 0.991 0.991 0.993
+chi-p 0.934 0.953 0.961 0.981 0.986 0.990 0.991 0.992 0.993
+$h=8$ asy-p 0.812 0.873 0.908 0.918 0.948 0.957 0.967 0.981 0.986
+chi-p 0.850 0.887 0.912 0.935 0.954 0.958 0.975 0.981 0.986
+7.1 $h=2$ asy-p 0.980 0.992 0.994 0.993 0.995 0.998 0.998 0.998 0.998
+chi-p 0.989 0.992 0.994 0.994 0.996 0.998 0.998 0.998 0.998
+$h=4$ asy-p 0.940 0.960 0.971 0.982 0.986 0.990 0.990 0.996 0.996
+chi-p 0.957 0.967 0.972 0.986 0.988 0.991 0.995 0.996 0.996
+$h=8$ asy-p 0.847 0.889 0.915 0.920 0.942 0.955 0.962 0.972 0.977
+chi-p 0.873 0.896 0.916 0.936 0.947 0.955 0.970 0.975 0.977
+Table S5: Empirical powers of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\mu=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Polynomial basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\mu}}-d_{\mu})/\sqrt{2d_{\mu}}$).
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.974 0.980 0.980 0.980 0.989 0.989 0.988 0.992 0.992
+chi-p 0.998 0.998 0.998 0.998 0.999 0.999 0.999 1 1
+$h=4$ asy-p 0.923 0.958 0.966 0.973 0.990 0.990 0.988 0.992 0.992
+chi-p 0.946 0.976 0.984 0.991 1 1 0.999 1 1
+$h=8$ asy-p 0.785 0.847 0.876 0.791 0.849 0.878 0.810 0.869 0.898
+chi-p 0.805 0.863 0.893 0.806 0.858 0.887 0.820 0.877 0.906
+7.1 $h=2$ asy-p 0.765 0.828 0.863 0.881 0.909 0.922 0.945 0.964 0.967
+chi-p 0.784 0.844 0.879 0.898 0.919 0.932 0.956 0.972 0.975
+$h=4$ asy-p 0.749 0.817 0.828 0.845 0.853 0.859 0.907 0.927 0.963
+chi-p 0.768 0.832 0.844 0.861 0.862 0.868 0.918 0.935 0.971
+$h=8$ asy-p 0.740 0.785 0.808 0.814 0.851 0.855 0.886 0.907 0.923
+chi-p 0.759 0.800 0.823 0.829 0.860 0.864 0.896 0.915 0.931
+7.1 $h=2$ asy-p 0.968 0.974 0.975 0.979 0.988 0.989 0.988 0.991 0.991
+chi-p 0.992 0.992 0.993 0.997 0.998 0.999 0.999 0.999 0.999
+$h=4$ asy-p 0.953 0.969 0.976 0.977 0.988 0.990 0.989 0.992 0.992
+chi-p 0.977 0.987 0.994 0.995 0.998 1 1 1 1
+$h=8$ asy-p 0.709 0.726 0.781 0.794 0.841 0.866 0.874 0.899 0.913
+chi-p 0.727 0.740 0.796 0.809 0.850 0.875 0.884 0.907 0.921
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.764 0.795 0.846 0.862 0.895 0.917 0.943 0.971 0.979
+chi-p 0.783 0.810 0.862 0.878 0.905 0.927 0.954 0.979 0.987
+$h=4$ asy-p 0.758 0.820 0.826 0.828 0.849 0.861 0.925 0.939 0.969
+chi-p 0.777 0.836 0.842 0.844 0.858 0.870 0.936 0.947 0.977
+$h=8$ asy-p 0.809 0.821 0.831 0.855 0.869 0.879 0.878 0.888 0.900
+chi-p 0.829 0.837 0.847 0.871 0.878 0.888 0.888 0.896 0.908
+7.1 $h=2$ asy-p 0.787 0.822 0.844 0.870 0.893 0.914 0.923 0.951 0.969
+chi-p 0.807 0.838 0.860 0.886 0.903 0.924 0.934 0.959 0.977
+$h=4$ asy-p 0.807 0.824 0.832 0.834 0.864 0.867 0.874 0.881 0.886
+chi-p 0.827 0.840 0.848 0.850 0.873 0.876 0.884 0.889 0.894
+$h=8$ asy-p 0.751 0.794 0.834 0.837 0.866 0.877 0.885 0.899 0.913
+chi-p 0.770 0.809 0.850 0.853 0.875 0.886 0.895 0.907 0.921
+7.1 $h=2$ asy-p 0.734 0.756 0.792 0.857 0.893 0.914 0.923 0.951 0.981
+chi-p 0.753 0.770 0.807 0.873 0.903 0.924 0.934 0.959 0.989
+$h=4$ asy-p 0.741 0.771 0.780 0.822 0.841 0.877 0.884 0.940 0.969
+chi-p 0.760 0.786 0.795 0.838 0.850 0.886 0.894 0.948 0.977
+$h=8$ asy-p 0.798 0.814 0.817 0.852 0.860 0.872 0.883 0.885 0.892
+chi-p 0.818 0.829 0.832 0.868 0.869 0.881 0.893 0.893 0.900
+Table S6: Empirical sizes of $\mathbb{W}_{1}$ for the baseline SAR model proposed in Section 2.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.034 0.075 0.107 0.029 0.068 0.100 0.037 0.077 0.106
+chi-p 0.014 0.050 0.103 0.010 0.045 0.099 0.011 0.055 0.103
+$h=4$ asy-p 0.032 0.085 0.112 0.026 0.069 0.112 0.038 0.085 0.118
+chi-p 0.017 0.060 0.108 0.009 0.049 0.104 0.014 0.062 0.113
+$h=8$ asy-p 0.028 0.052 0.085 0.024 0.050 0.077 0.034 0.084 0.126
+chi-p 0.014 0.046 0.078 0.007 0.038 0.072 0.012 0.064 0.117
+7.1 $h=2$ asy-p 0.036 0.073 0.109 0.041 0.079 0.105 0.039 0.074 0.109
+chi-p 0.014 0.048 0.108 0.013 0.058 0.101 0.015 0.056 0.107
+$h=4$ asy-p 0.049 0.101 0.135 0.047 0.084 0.127 0.031 0.069 0.110
+chi-p 0.018 0.072 0.130 0.010 0.069 0.119 0.009 0.051 0.105
+$h=8$ asy-p 0.026 0.070 0.102 0.019 0.053 0.083 0.024 0.069 0.117
+chi-p 0.011 0.045 0.095 0.009 0.035 0.077 0.009 0.052 0.105
+7.1 $h=2$ asy-p 0.041 0.077 0.104 0.041 0.079 0.107 0.030 0.058 0.096
+chi-p 0.010 0.054 0.103 0.009 0.057 0.106 0.008 0.042 0.092
+$h=4$ asy-p 0.038 0.078 0.120 0.037 0.081 0.119 0.025 0.060 0.093
+chi-p 0.009 0.057 0.112 0.007 0.062 0.109 0.006 0.038 0.092
+$h=8$ asy-p 0.024 0.060 0.104 0.023 0.058 0.086 0.029 0.068 0.111
+chi-p 0.010 0.040 0.098 0.008 0.045 0.078 0.012 0.051 0.105
+$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.033 0.082 0.118 0.043 0.070 0.105 0.036 0.068 0.110
+chi-p 0.011 0.049 0.118 0.015 0.054 0.102 0.007 0.049 0.106
+$h=4$ asy-p 0.038 0.092 0.132 0.044 0.087 0.126 0.035 0.078 0.118
+chi-p 0.009 0.061 0.127 0.016 0.066 0.120 0.012 0.054 0.111
+$h=8$ asy-p 0.022 0.049 0.085 0.025 0.060 0.102 0.033 0.075 0.114
+chi-p 0.011 0.038 0.078 0.012 0.044 0.094 0.018 0.050 0.104
+7.1 $h=2$ asy-p 0.041 0.078 0.116 0.055 0.090 0.129 0.033 0.073 0.100
+chi-p 0.019 0.052 0.114 0.017 0.068 0.127 0.014 0.050 0.098
+$h=4$ asy-p 0.048 0.095 0.136 0.042 0.091 0.131 0.035 0.072 0.105
+chi-p 0.017 0.074 0.122 0.021 0.066 0.122 0.011 0.048 0.102
+$h=8$ asy-p 0.035 0.070 0.106 0.028 0.061 0.083 0.032 0.079 0.117
+chi-p 0.016 0.051 0.097 0.010 0.048 0.082 0.009 0.057 0.113
+7.1 $h=2$ asy-p 0.045 0.087 0.113 0.043 0.081 0.108 0.032 0.075 0.110
+chi-p 0.015 0.064 0.113 0.013 0.059 0.104 0.009 0.052 0.107
+$h=4$ asy-p 0.046 0.093 0.127 0.039 0.073 0.111 0.036 0.084 0.113
+chi-p 0.015 0.072 0.122 0.014 0.054 0.105 0.011 0.055 0.110
+$h=8$ asy-p 0.031 0.063 0.107 0.027 0.049 0.088 0.028 0.068 0.113
+chi-p 0.014 0.044 0.099 0.009 0.041 0.080 0.007 0.056 0.101
+Table S7: Empirical sizes of $\mathbb{W}_{2}$ for the baseline SAR model with SHAC correction proposed in Section 3.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.038 0.062 0.080 0.026 0.061 0.094 0.042 0.082 0.118
+chi-p 0.010 0.050 0.078 0.009 0.040 0.090 0.013 0.058 0.115
+$h=4$ asy-p 0.051 0.081 0.107 0.034 0.063 0.092 0.040 0.082 0.114
+chi-p 0.027 0.062 0.103 0.013 0.047 0.086 0.013 0.065 0.110
+$h=8$ asy-p 0.082 0.120 0.146 0.034 0.066 0.081 0.032 0.070 0.111
+chi-p 0.051 0.104 0.141 0.016 0.049 0.078 0.021 0.054 0.103
+7.1 $h=2$ asy-p 0.022 0.045 0.074 0.030 0.055 0.089 0.029 0.080 0.120
+chi-p 0.008 0.029 0.070 0.008 0.040 0.086 0.006 0.052 0.119
+$h=4$ asy-p 0.036 0.066 0.089 0.023 0.060 0.090 0.025 0.055 0.108
+chi-p 0.017 0.056 0.087 0.007 0.042 0.087 0.007 0.037 0.100
+$h=8$ asy-p 0.065 0.105 0.139 0.032 0.062 0.085 0.031 0.064 0.101
+chi-p 0.045 0.090 0.136 0.012 0.042 0.081 0.012 0.048 0.092
+7.1 $h=2$ asy-p 0.034 0.062 0.080 0.033 0.075 0.098 0.033 0.069 0.108
+chi-p 0.009 0.048 0.078 0.006 0.047 0.096 0.005 0.052 0.107
+$h=4$ asy-p 0.050 0.086 0.114 0.033 0.074 0.107 0.032 0.073 0.098
+chi-p 0.018 0.069 0.109 0.006 0.054 0.102 0.007 0.056 0.095
+$h=8$ asy-p 0.080 0.123 0.149 0.025 0.056 0.082 0.032 0.072 0.109
+chi-p 0.046 0.108 0.144 0.012 0.041 0.077 0.018 0.056 0.097
+$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.042 0.063 0.078 0.033 0.055 0.086 0.040 0.067 0.093
+chi-p 0.014 0.053 0.077 0.013 0.043 0.085 0.016 0.046 0.093
+$h=4$ asy-p 0.048 0.081 0.118 0.039 0.064 0.092 0.032 0.068 0.101
+chi-p 0.026 0.069 0.111 0.020 0.052 0.087 0.015 0.052 0.097
+$h=8$ asy-p 0.085 0.126 0.155 0.048 0.071 0.093 0.041 0.076 0.109
+chi-p 0.056 0.109 0.152 0.028 0.062 0.091 0.013 0.057 0.102
+7.1 $h=2$ asy-p 0.030 0.046 0.073 0.027 0.048 0.074 0.027 0.071 0.096
+chi-p 0.012 0.033 0.070 0.012 0.038 0.072 0.006 0.048 0.095
+$h=4$ asy-p 0.048 0.083 0.112 0.026 0.055 0.075 0.029 0.059 0.085
+chi-p 0.027 0.060 0.106 0.010 0.035 0.072 0.009 0.039 0.081
+$h=8$ asy-p 0.080 0.119 0.152 0.030 0.055 0.081 0.043 0.079 0.108
+chi-p 0.054 0.100 0.143 0.013 0.045 0.076 0.017 0.064 0.103
+7.1 $h=2$ asy-p 0.028 0.051 0.072 0.028 0.059 0.078 0.026 0.058 0.083
+chi-p 0.013 0.039 0.070 0.009 0.043 0.077 0.011 0.043 0.080
+$h=4$ asy-p 0.049 0.078 0.098 0.036 0.064 0.087 0.033 0.060 0.082
+chi-p 0.027 0.061 0.095 0.018 0.051 0.085 0.011 0.049 0.077
+$h=8$ asy-p 0.094 0.128 0.152 0.037 0.064 0.084 0.039 0.064 0.091
+chi-p 0.065 0.118 0.150 0.021 0.055 0.079 0.023 0.058 0.083
+Table S8: Empirical sizes of $\mathbb{W}_{3}$ for the SAR model with nonparametric spatial weights proposed in Section 4.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\tau}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.042 0.067 0.087 0.029 0.054 0.086 0.036 0.062 0.080
+chi-p 0.018 0.060 0.084 0.007 0.042 0.084 0.012 0.042 0.080
+$h=4$ asy-p 0.054 0.090 0.119 0.029 0.067 0.091 0.028 0.061 0.089
+chi-p 0.024 0.071 0.115 0.011 0.046 0.085 0.011 0.046 0.089
+$h=8$ asy-p 0.060 0.117 0.147 0.040 0.073 0.104 0.034 0.065 0.092
+chi-p 0.012 0.095 0.143 0.019 0.059 0.101 0.016 0.056 0.089
+7.1 $h=2$ asy-p 0.022 0.048 0.075 0.036 0.060 0.077 0.031 0.056 0.089
+chi-p 0.010 0.032 0.074 0.013 0.046 0.076 0.007 0.042 0.089
+$h=4$ asy-p 0.038 0.067 0.096 0.038 0.067 0.095 0.028 0.055 0.083
+chi-p 0.017 0.053 0.091 0.019 0.051 0.089 0.012 0.041 0.079
+$h=8$ asy-p 0.078 0.112 0.147 0.048 0.086 0.113 0.039 0.072 0.099
+chi-p 0.053 0.096 0.141 0.034 0.072 0.107 0.019 0.058 0.096
+7.1 $h=2$ asy-p 0.038 0.067 0.087 0.026 0.057 0.086 0.029 0.052 0.070
+chi-p 0.013 0.053 0.086 0.004 0.040 0.086 0.012 0.035 0.067
+$h=4$ asy-p 0.053 0.101 0.122 0.035 0.064 0.091 0.030 0.058 0.077
+chi-p 0.022 0.080 0.115 0.013 0.052 0.086 0.012 0.045 0.076
+$h=8$ asy-p 0.088 0.123 0.155 0.030 0.058 0.094 0.041 0.069 0.099
+chi-p 0.061 0.111 0.144 0.019 0.046 0.091 0.029 0.061 0.095
+$d_{\tau}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.036 0.061 0.081 0.039 0.069 0.098 0.029 0.059 0.086
+chi-p 0.016 0.040 0.079 0.013 0.052 0.097 0.013 0.039 0.083
+$h=4$ asy-p 0.051 0.079 0.106 0.037 0.079 0.124 0.031 0.062 0.083
+chi-p 0.027 0.064 0.102 0.015 0.056 0.120 0.017 0.052 0.079
+$h=8$ asy-p 0.089 0.128 0.154 0.056 0.090 0.127 0.045 0.073 0.094
+chi-p 0.062 0.109 0.152 0.026 0.077 0.125 0.030 0.057 0.091
+7.1 $h=2$ asy-p 0.031 0.049 0.083 0.041 0.070 0.097 0.030 0.050 0.071
+chi-p 0.012 0.039 0.080 0.014 0.050 0.095 0.011 0.035 0.069
+$h=4$ asy-p 0.054 0.082 0.124 0.033 0.075 0.101 0.025 0.055 0.075
+chi-p 0.027 0.064 0.117 0.014 0.052 0.099 0.013 0.041 0.070
+$h=8$ asy-p 0.073 0.133 0.168 0.056 0.103 0.151 0.040 0.067 0.089
+chi-p 0.014 0.114 0.162 0.031 0.086 0.143 0.021 0.056 0.085
+7.1 $h=2$ asy-p 0.031 0.056 0.086 0.034 0.069 0.095 0.030 0.058 0.078
+chi-p 0.016 0.041 0.083 0.012 0.050 0.092 0.009 0.044 0.076
+$h=4$ asy-p 0.050 0.088 0.116 0.039 0.073 0.098 0.038 0.060 0.078
+chi-p 0.028 0.068 0.113 0.009 0.059 0.096 0.015 0.052 0.076
+$h=8$ asy-p 0.082 0.121 0.155 0.044 0.091 0.129 0.043 0.067 0.089
+chi-p 0.032 0.102 0.148 0.022 0.072 0.126 0.029 0.058 0.084
+Table S9: Empirical sizes of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.057 0.076 0.097 0.042 0.056 0.078 0.044 0.068 0.085
+chi-p 0.033 0.063 0.096 0.024 0.050 0.075 0.019 0.056 0.084
+$h=4$ asy-p 0.088 0.126 0.147 0.054 0.086 0.100 0.041 0.072 0.092
+chi-p 0.059 0.106 0.143 0.034 0.068 0.097 0.027 0.062 0.088
+$h=8$ asy-p 0.085 0.126 0.143 0.029 0.045 0.059 0.023 0.030 0.074
+chi-p 0.067 0.109 0.139 0.018 0.038 0.058 0.015 0.038 0.083
+7.1 $h=2$ asy-p 0.074 0.103 0.121 0.044 0.077 0.096 0.037 0.054 0.071
+chi-p 0.052 0.090 0.119 0.021 0.057 0.093 0.015 0.046 0.070
+$h=4$ asy-p 0.105 0.143 0.163 0.053 0.080 0.107 0.041 0.066 0.075
+chi-p 0.078 0.125 0.162 0.023 0.064 0.105 0.021 0.057 0.073
+$h=8$ asy-p 0.098 0.135 0.154 0.032 0.046 0.055 0.010 0.027 0.071
+chi-p 0.076 0.117 0.150 0.028 0.041 0.054 0.007 0.036 0.086
+7.1 $h=2$ asy-p 0.077 0.102 0.124 0.052 0.070 0.080 0.041 0.069 0.083
+chi-p 0.050 0.087 0.122 0.029 0.062 0.078 0.020 0.054 0.078
+$h=4$ asy-p 0.084 0.126 0.156 0.041 0.073 0.098 0.046 0.073 0.096
+chi-p 0.059 0.115 0.147 0.021 0.060 0.095 0.029 0.058 0.092
+$h=8$ asy-p 0.108 0.142 0.167 0.032 0.045 0.055 0.013 0.033 0.062
+chi-p 0.082 0.134 0.162 0.025 0.041 0.053 0.008 0.038 0.082
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.089 0.122 0.140 0.042 0.068 0.082 0.029 0.048 0.063
+chi-p 0.055 0.103 0.137 0.025 0.056 0.081 0.016 0.038 0.083
+$h=4$ asy-p 0.089 0.128 0.156 0.041 0.068 0.084 0.024 0.041 0.072
+chi-p 0.063 0.115 0.152 0.024 0.053 0.088 0.013 0.034 0.085
+$h=8$ asy-p 0.101 0.136 0.162 0.034 0.052 0.062 0.019 0.035 0.077
+chi-p 0.085 0.129 0.159 0.021 0.045 0.060 0.012 0.033 0.086
+7.1 $h=2$ asy-p 0.084 0.111 0.129 0.045 0.071 0.091 0.032 0.053 0.065
+chi-p 0.053 0.097 0.128 0.024 0.059 0.090 0.013 0.043 0.092
+$h=4$ asy-p 0.087 0.123 0.153 0.035 0.057 0.075 0.020 0.038 0.071
+chi-p 0.063 0.104 0.144 0.023 0.049 0.088 0.007 0.025 0.089
+$h=8$ asy-p 0.092 0.120 0.140 0.037 0.056 0.072 0.009 0.017 0.073
+chi-p 0.068 0.112 0.138 0.024 0.047 0.083 0.005 0.014 0.089
+7.1 $h=2$ asy-p 0.088 0.121 0.143 0.052 0.079 0.094 0.030 0.058 0.071
+chi-p 0.059 0.108 0.138 0.028 0.066 0.092 0.015 0.045 0.088
+$h=4$ asy-p 0.089 0.125 0.150 0.038 0.062 0.080 0.023 0.041 0.073
+chi-p 0.063 0.113 0.143 0.025 0.049 0.077 0.011 0.032 0.088
+$h=8$ asy-p 0.103 0.136 0.159 0.033 0.057 0.064 0.013 0.021 0.070
+chi-p 0.082 0.123 0.155 0.031 0.048 0.061 0.007 0.019 0.088
+Table S10: Empirical sizes of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\mu=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\mu}}-d_{\mu})/\sqrt{2d_{\mu}}$).
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.059 0.097 0.131 0.020 0.049 0.076 0.014 0.028 0.037
+chi-p 0.029 0.082 0.122 0.006 0.031 0.072 0.002 0.039 0.080
+$h=4$ asy-p 0.097 0.136 0.165 0.019 0.027 0.040 0.001 0.004 0.011
+chi-p 0.061 0.121 0.158 0.010 0.033 0.078 0.001 0.036 0.080
+$h=8$ asy-p 0.074 0.120 0.141 0.054 0.061 0.169 0.017 0.037 0.076
+chi-p 0.048 0.103 0.139 0.023 0.025 0.070 0.014 0.034 0.073
+7.1 $h=2$ asy-p 0.061 0.100 0.132 0.026 0.050 0.075 0.023 0.033 0.088
+chi-p 0.029 0.077 0.124 0.009 0.041 0.069 0.007 0.027 0.084
+$h=4$ asy-p 0.093 0.138 0.164 0.014 0.033 0.052 0.011 0.012 0.085
+chi-p 0.060 0.120 0.162 0.005 0.023 0.048 0.002 0.011 0.085
+$h=8$ asy-p 0.104 0.127 0.151 0.038 0.049 0.119 0.042 0.058 0.083
+chi-p 0.071 0.121 0.146 0.029 0.046 0.102 0.031 0.053 0.084
+7.1 $h=2$ asy-p 0.072 0.119 0.157 0.022 0.048 0.069 0.015 0.025 0.054
+chi-p 0.042 0.100 0.151 0.005 0.030 0.076 0.005 0.033 0.080
+$h=4$ asy-p 0.084 0.127 0.166 0.014 0.034 0.045 0.005 0.009 0.016
+chi-p 0.064 0.110 0.158 0.006 0.037 0.072 0.002 0.037 0.074
+$h=8$ asy-p 0.107 0.148 0.180 0.033 0.046 0.090 0.053 0.069 0.079
+chi-p 0.086 0.131 0.175 0.019 0.041 0.077 0.035 0.065 0.077
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.060 0.098 0.133 0.020 0.050 0.077 0.014 0.038 0.077
+chi-p 0.029 0.083 0.123 0.006 0.031 0.073 0.002 0.039 0.081
+$h=4$ asy-p 0.098 0.138 0.167 0.019 0.027 0.040 0.001 0.024 0.071
+chi-p 0.062 0.122 0.160 0.010 0.033 0.079 0.000 0.036 0.081
+$h=8$ asy-p 0.075 0.121 0.143 0.055 0.062 0.091 0.017 0.037 0.077
+chi-p 0.049 0.104 0.141 0.023 0.026 0.071 0.014 0.035 0.074
+7.1 $h=2$ asy-p 0.062 0.101 0.134 0.026 0.051 0.076 0.023 0.033 0.089
+chi-p 0.029 0.078 0.125 0.009 0.041 0.070 0.007 0.027 0.085
+$h=4$ asy-p 0.094 0.140 0.166 0.014 0.033 0.053 0.011 0.012 0.086
+chi-p 0.061 0.121 0.164 0.005 0.023 0.049 0.002 0.011 0.086
+$h=8$ asy-p 0.105 0.129 0.153 0.039 0.049 0.120 0.042 0.059 0.084
+chi-p 0.072 0.122 0.148 0.029 0.046 0.103 0.031 0.054 0.085
+7.1 $h=2$ asy-p 0.073 0.120 0.159 0.022 0.049 0.070 0.015 0.025 0.055
+chi-p 0.043 0.101 0.153 0.005 0.030 0.077 0.005 0.033 0.081
+$h=4$ asy-p 0.085 0.129 0.168 0.014 0.034 0.046 0.005 0.049 0.076
+chi-p 0.065 0.111 0.160 0.006 0.037 0.073 0.002 0.037 0.075
+$h=8$ asy-p 0.108 0.150 0.182 0.034 0.047 0.091 0.054 0.070 0.080
+chi-p 0.087 0.133 0.177 0.019 0.042 0.078 0.035 0.062 0.078
+Table S11: Empirical powers of $\mathbb{W}_{2}$ for the baseline SAR model with SHAC correction proposed in Section 3.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.995 0.997 0.999 0.997 0.998 0.998 1 1 1
+chi-p 0.980 0.996 0.999 0.997 0.998 0.998 1 1 1
+$h=4$ asy-p 0.996 0.999 0.999 1 1 1 1 1 1
+chi-p 0.996 0.997 0.999 1 1 1 1 1 1
+$h=8$ asy-p 0.978 0.979 0.979 0.999 0.999 0.999 1 1 1
+chi-p 0.976 0.979 0.979 0.999 0.999 0.999 1 1 1
+7.1 $h=2$ asy-p 0.987 0.992 0.992 0.996 0.998 0.998 0.999 0.999 1
+chi-p 0.977 0.990 0.992 0.993 0.997 0.998 0.998 0.999 1
+$h=4$ asy-p 0.996 0.998 0.998 0.999 1 1 1 1 1
+chi-p 0.996 0.998 0.998 0.999 1 1 1 1 1
+$h=8$ asy-p 0.980 0.980 0.981 0.996 0.996 0.996 1 1 1
+chi-p 0.977 0.980 0.981 0.996 0.996 0.996 1 1 1
+7.1 $h=2$ asy-p 0.986 0.994 0.996 0.999 1 1 0.998 0.999 1
+chi-p 0.978 0.989 0.996 0.998 0.999 1 0.998 0.999 1
+$h=4$ asy-p 0.998 0.999 0.999 0.999 1 1 1 1 1
+chi-p 0.995 0.998 0.999 0.999 1 1 0.999 1 1
+$h=8$ asy-p 0.981 0.982 0.982 0.998 0.998 0.998 1 1 1
+chi-p 0.980 0.982 0.982 0.998 0.998 0.998 1 1 1
+$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.995 0.997 0.998 0.998 0.999 0.999 0.999 0.999 0.999
+chi-p 0.991 0.996 0.998 0.997 0.998 0.999 0.999 0.999 0.999
+$h=4$ asy-p 1 1 1 1 1 1 1 1 1
+chi-p 1 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.943 0.943 0.945 0.999 0.999 0.999 1 1 1
+chi-p 0.943 0.943 0.945 0.998 0.999 0.999 1 1 1
+7.1 $h=2$ asy-p 0.998 0.998 1 0.997 0.997 0.999 1 1 1
+chi-p 0.996 0.998 1 0.997 0.997 0.999 1 1 1
+$h=4$ asy-p 0.998 0.999 1 0.999 0.999 0.999 1 1 1
+chi-p 0.998 0.999 1 0.999 0.999 0.999 1 1 1
+$h=8$ asy-p 0.947 0.949 0.949 0.998 0.998 0.998 1 1 1
+chi-p 0.946 0.948 0.949 0.997 0.998 0.998 1 1 1
+7.1 $h=2$ asy-p 0.997 0.999 0.999 0.998 0.999 0.999 0.999 1 1
+chi-p 0.995 0.999 0.999 0.998 0.998 0.999 0.999 1 1
+$h=4$ asy-p 1 1 1 0.999 0.999 1 1 1 1
+chi-p 1 1 1 0.999 0.999 0.999 1 1 1
+$h=8$ asy-p 0.959 0.959 0.960 0.996 0.996 0.996 1 1 1
+chi-p 0.959 0.959 0.960 0.996 0.996 0.996 1 1 1
+Table S12: Empirical powers of $\mathbb{W}_{3}$ for the SAR model with nonparametric spatial weights proposed in Section 4.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\tau}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.974 0.983 0.988 0.973 0.983 0.989 0.982 0.992 0.994
+chi-p 0.942 0.980 0.988 0.954 0.977 0.988 0.966 0.988 0.994
+$h=4$ asy-p 0.986 0.996 0.998 0.986 0.995 0.996 0.994 0.997 0.998
+chi-p 0.976 0.994 0.998 0.973 0.991 0.996 0.985 0.997 0.998
+$h=8$ asy-p 0.853 0.857 0.859 0.894 0.901 0.903 0.996 0.997 0.998
+chi-p 0.849 0.855 0.858 0.689 0.901 0.903 0.992 0.997 0.998
+7.1 $h=2$ asy-p 0.966 0.978 0.983 0.973 0.982 0.987 0.982 0.990 0.992
+chi-p 0.942 0.976 0.983 0.946 0.978 0.987 0.962 0.987 0.992
+$h=4$ asy-p 0.992 0.994 0.997 0.990 0.994 0.997 0.991 0.996 0.998
+chi-p 0.980 0.994 0.995 0.972 0.992 0.997 0.980 0.995 0.998
+$h=8$ asy-p 0.858 0.858 0.858 0.899 0.905 0.906 0.991 0.995 0.999
+chi-p 0.855 0.858 0.858 0.893 0.904 0.906 0.983 0.995 0.998
+7.1 $h=2$ asy-p 0.973 0.985 0.987 0.975 0.991 0.992 0.986 0.995 0.997
+chi-p 0.946 0.980 0.987 0.955 0.984 0.992 0.973 0.989 0.997
+$h=4$ asy-p 0.988 0.994 0.998 0.993 0.996 0.998 0.992 0.998 0.999
+chi-p 0.971 0.990 0.997 0.982 0.995 0.998 0.986 0.993 0.999
+$h=8$ asy-p 0.865 0.866 0.866 0.904 0.910 0.911 0.997 0.999 0.999
+chi-p 0.864 0.865 0.866 0.901 0.909 0.910 0.991 0.999 0.999
+$d_{\tau}=4$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.996 0.998 1 0.999 0.999 1 1 1 1
+chi-p 0.992 0.996 1 0.998 0.999 1 1 1 1
+$h=4$ asy-p 1 1 1 1 1 1 1 1 1
+chi-p 1 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.815 0.816 0.819 0.944 0.944 0.944 0.982 0.982 0.982
+chi-p 0.813 0.815 0.817 0.944 0.944 0.944 0.982 0.982 0.982
+7.1 $h=2$ asy-p 0.998 0.999 0.999 0.999 0.999 0.999 1 1 1
+chi-p 0.997 0.998 0.999 0.999 0.999 0.999 1 1 1
+$h=4$ asy-p 0.999 1 1 1 1 1 1 1 1
+chi-p 0.999 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.810 0.810 0.812 0.957 0.957 0.957 0.975 0.975 0.975
+chi-p 0.810 0.810 0.812 0.957 0.957 0.957 0.975 0.975 0.975
+7.1 $h=2$ asy-p 0.996 0.999 0.999 0.998 0.998 0.998 1 1 1
+chi-p 0.994 0.997 0.999 0.998 0.998 0.998 1 1 1
+$h=4$ asy-p 1 1 1 1 1 1 1 1 1
+chi-p 1 1 1 1 1 1 1 1 1
+$h=8$ asy-p 0.833 0.833 0.834 0.949 0.949 0.949 0.979 0.979 0.979
+chi-p 0.833 0.833 0.834 0.949 0.949 0.949 0.979 0.979 0.979
+Table S13: Empirical powers of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\alpha=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$).
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.973 0.982 0.990 0.988 0.992 0.995 0.996 0.998 0.998
+chi-p 0.977 0.986 0.990 0.992 0.995 0.995 0.997 0.998 0.998
+$h=4$ asy-p 0.947 0.968 0.975 0.975 0.982 0.985 0.993 0.996 0.997
+chi-p 0.959 0.972 0.975 0.976 0.984 0.985 0.996 0.996 0.997
+$h=8$ asy-p 0.913 0.931 0.943 0.945 0.961 0.969 0.976 0.984 0.988
+chi-p 0.922 0.935 0.943 0.956 0.963 0.969 0.980 0.985 0.988
+7.1 $h=2$ asy-p 0.976 0.989 0.992 0.989 0.993 0.996 0.995 0.999 1
+chi-p 0.988 0.990 0.992 0.992 0.996 0.996 0.998 1 1
+$h=4$ asy-p 0.948 0.968 0.975 0.978 0.986 0.991 0.989 0.994 0.996
+chi-p 0.962 0.971 0.976 0.985 0.989 0.991 0.992 0.994 0.996
+$h=8$ asy-p 0.892 0.927 0.938 0.948 0.964 0.973 0.976 0.985 0.988
+chi-p 0.915 0.931 0.942 0.957 0.969 0.973 0.983 0.988 0.989
+7.1 $h=2$ asy-p 0.962 0.973 0.985 0.986 0.991 0.993 0.997 0.998 0.999
+chi-p 0.969 0.981 0.985 0.991 0.991 0.993 0.998 0.998 0.999
+$h=4$ asy-p 0.952 0.973 0.978 0.975 0.984 0.988 0.998 0.999 1
+chi-p 0.963 0.974 0.979 0.980 0.985 0.988 0.999 0.999 1
+$h=8$ asy-p 0.884 0.918 0.934 0.953 0.965 0.977 0.971 0.981 0.984
+chi-p 0.906 0.927 0.934 0.960 0.971 0.978 0.979 0.983 0.985
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.977 0.990 0.992 0.993 0.996 0.997 0.997 1 1
+chi-p 0.986 0.990 0.992 0.996 0.997 0.998 1 1 1
+$h=4$ asy-p 0.917 0.946 0.964 0.986 0.991 0.994 0.992 0.994 0.994
+chi-p 0.935 0.957 0.966 0.991 0.992 0.994 0.993 0.994 0.995
+$h=8$ asy-p 0.817 0.868 0.889 0.920 0.951 0.960 0.966 0.975 0.984
+chi-p 0.845 0.878 0.894 0.942 0.955 0.960 0.971 0.979 0.984
+7.1 $h=2$ asy-p 0.983 0.991 0.994 0.991 0.996 0.996 0.997 0.998 0.999
+chi-p 0.986 0.992 0.994 0.995 0.996 0.996 0.997 0.998 0.999
+$h=4$ asy-p 0.937 0.955 0.968 0.968 0.982 0.986 0.993 0.996 0.996
+chi-p 0.945 0.961 0.970 0.979 0.982 0.988 0.995 0.996 0.996
+$h=8$ asy-p 0.813 0.884 0.910 0.919 0.951 0.959 0.967 0.980 0.985
+chi-p 0.854 0.897 0.913 0.939 0.954 0.960 0.975 0.982 0.986
+7.1 $h=2$ asy-p 0.974 0.985 0.988 0.992 0.995 0.999 0.999 0.999 0.999
+chi-p 0.982 0.988 0.988 0.995 0.997 0.999 0.999 0.999 0.999
+$h=4$ asy-p 0.944 0.961 0.972 0.972 0.980 0.985 0.990 0.994 0.994
+chi-p 0.955 0.968 0.973 0.975 0.982 0.986 0.993 0.994 0.995
+$h=8$ asy-p 0.823 0.868 0.895 0.925 0.944 0.959 0.966 0.980 0.983
+chi-p 0.850 0.879 0.902 0.937 0.950 0.959 0.975 0.982 0.983
+Table S14: Empirical powers of $\mathbb{W}_{4}$ for the varying spatial coefficient SAR model proposed in Section 5.
+Null approximating $H_{0}^{true}$, $H_{0}:\mu=0$.
+Rejection probabilities at 1%, 5% and 10% levels.
+Trigonometric basis functions.
+P-values: asy-p (standard normal) and chi-p (chi-square calibration $(\chi^{2}_{d_{\mu}}-d_{\mu})/\sqrt{2d_{\mu}}$).
+$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.973 0.979 0.979 0.979 0.989 0.989 0.988 0.991 0.992
+chi-p 0.997 0.997 0.997 0.997 0.999 0.999 0.999 0.999 1
+$h=4$ asy-p 0.969 0.978 0.979 0.982 0.990 0.990 0.989 0.992 0.992
+chi-p 0.993 0.996 0.997 1 1 1 1 1 1
+$h=8$ asy-p 0.923 0.941 0.947 0.975 0.985 0.986 0.985 0.991 0.992
+chi-p 0.946 0.959 0.965 0.993 0.995 0.996 0.996 0.999 1
+7.1 $h=2$ asy-p 0.765 0.817 0.844 0.881 0.909 0.922 0.945 0.964 0.967
+chi-p 0.784 0.832 0.860 0.898 0.919 0.932 0.956 0.972 0.975
+$h=4$ asy-p 0.808 0.833 0.866 0.868 0.906 0.930 0.936 0.962 0.968
+chi-p 0.828 0.849 0.882 0.884 0.916 0.940 0.947 0.970 0.976
+$h=8$ asy-p 0.714 0.720 0.739 0.766 0.798 0.822 0.826 0.866 0.888
+chi-p 0.732 0.734 0.753 0.781 0.807 0.831 0.836 0.873 0.896
+7.1 $h=2$ asy-p 0.975 0.981 0.981 0.981 0.990 0.990 0.989 0.992 0.992
+chi-p 0.999 0.999 0.999 0.999 1 1 1 1 1
+$h=4$ asy-p 0.970 0.978 0.979 0.982 0.990 0.990 0.989 0.992 0.992
+chi-p 0.994 0.996 0.997 1 1 1 1 1 1
+$h=8$ asy-p 0.919 0.945 0.951 0.971 0.984 0.985 0.985 0.989 0.991
+chi-p 0.942 0.963 0.969 0.989 0.994 0.995 0.996 0.997 0.999
+$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
+1% 5% 10% 1% 5% 10% 1% 5% 10%
+7.1 $h=2$ asy-p 0.740 0.816 0.837 0.879 0.916 0.927 0.931 0.951 0.965
+chi-p 0.759 0.831 0.853 0.896 0.926 0.937 0.942 0.959 0.973
+$h=4$ asy-p 0.723 0.766 0.793 0.832 0.874 0.891 0.901 0.925 0.971
+chi-p 0.741 0.781 0.808 0.848 0.883 0.900 0.912 0.933 0.979
+$h=8$ asy-p 0.728 0.802 0.804 0.806 0.840 0.842 0.869 0.883 0.908
+chi-p 0.746 0.817 0.819 0.821 0.849 0.851 0.879 0.891 0.916
+7.1 $h=2$ asy-p 0.786 0.816 0.830 0.858 0.867 0.884 0.898 0.926 0.940
+chi-p 0.806 0.831 0.846 0.874 0.876 0.893 0.908 0.934 0.948
+$h=4$ asy-p 0.708 0.769 0.796 0.823 0.858 0.875 0.908 0.927 0.944
+chi-p 0.726 0.784 0.811 0.839 0.867 0.884 0.919 0.935 0.952
+$h=8$ asy-p 0.784 0.789 0.801 0.803 0.824 0.830 0.832 0.872 0.882
+chi-p 0.804 0.804 0.816 0.818 0.833 0.839 0.842 0.880 0.890
+7.1 $h=2$ asy-p 0.763 0.813 0.833 0.868 0.906 0.930 0.936 0.962 0.968
+chi-p 0.782 0.828 0.849 0.884 0.916 0.940 0.947 0.970 0.976
+$h=4$ asy-p 0.713 0.762 0.791 0.830 0.865 0.884 0.898 0.926 0.940
+chi-p 0.731 0.776 0.806 0.846 0.874 0.893 0.908 0.934 0.948
+$h=8$ asy-p 0.700 0.775 0.789 0.818 0.830 0.833 0.870 0.896 0.908
+chi-p 0.718 0.790 0.804 0.833 0.839 0.842 0.880 0.904 0.916
 (a) Capital
@@
 (c) CRS
-Figure 1: Cobb-Douglas versus semiparametric production.
-Table 3: Estimation and testing results in Section 6.2
-Panel A: polynomial series $h=2$
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.238*** 0.219*** 0.238*** 0.219*** 0.214** 0.512*** 0.468***
-(0.049) (0.044) (0.028) (0.022) (0.084) (0.073) (0.089)
-$\tau_{2}/\mu_{2}$ 0.330** -0.356*** -0.311***
-(0.129) (0.049) (0.057)
-Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 40.041*** 33.913*** 290.165*** 33.913*** 183.122*** 91.459*** 123.439***
-$H_{0}:\mu=0$ 32.336*** 35.141***
-Panel B: polynomial series $h=4$
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.239*** 0.269*** 0.239** 0.269*** 0.162** 0.169** 0.169***
-(0.050) (0.043) (0.095) (0.027) (0.069) (0.080) (0.019)
-$\tau_{2}/\mu_{2}$ 0.175*** -0.363** -0.363***
-(0.042) (0.177) (0.080)
-$\tau_{3}/\mu_{3}$ 0.241*** 0.345** 0.345***
-(0.062) (0.147) (0.113)
-$\tau_{4}/\mu_{4}$ 0.082** -0.122** -0.123**
-(0.035) (0.060) (0.053)
-$\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 33.876*** 20.486*** 392.020*** 928.892*** 601.967*** 394.776*** 352.386***
-$H_{0}:\mu=0$ 56.124*** 66.981***
-Panel C: trigonometric series $h=2$
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.230*** 0.275*** 0.230*** 0.275*** 0.211*** 0.039 0.043***
-(0.050) (0.043) (0.026) (0.031) (0.065) (0.284) (0.005)
-$\tau_{2}/\mu_{2}$ 0.258*** -0.397** -0.437***
-(0.039) (0.186) (0.048)
-$\tau_{3}/\mu_{3}$ 0.255*** 0.229** 0.253***
-(0.055) (0.107) (0.028)
-$\tau_{4}/\mu_{4}$ 0.092*** 0.333* 0.367***
-(0.031) (0.190) (0.041)
-Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 40.041*** 33.913*** 290.165*** 33.913*** 258.704*** 91.459*** 123.439***
-$H_{0}:\mu=0$ 43.155*** 46.397***
-Panel D: trigonometric series $h=4$
-Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{G}$ $W_{n}^{q}$ $W_{n}^{t}$
-$\lambda/\tau_{1}/\mu_{1}$ 0.254*** 0.225*** 0.254*** 0.225*** 0.259*** 0.177*** 0.121**
-(0.049) (0.044) (0.028) (0.023) (0.069) (0.027) (0.057)
-$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
-(0.149) (0.040) (0.083)
-Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
-$H_{0}:\alpha=0$ 38.897*** 34.423*** 229.115*** 325.327*** 266.301*** 175.332*** 144.982***
-$H_{0}:\mu=0$ 28.883*** 21.326***
-Panel E: Comparison of parametric method and semiparametric method in Section 6.2
-polynomial, $h=2$ polynomial, $h=4$ trigonometric, $h=2$ trigonometric, $h=4$
-$W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$
-0.3136 0.3371 0.3252 0.3406 0.3074 0.3221 0.2996 0.3125
-Standard errors in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
-Appendix D Simulation Results
-Table 4: Empirical sizes for $\mathbb{W}_{1}$, rejection probabilities at 1%, 5% and 10% levels.
-$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.035 0.075 0.115 0.033 0.063 0.098 0.042 0.074 0.107
-chi-p 0.013 0.051 0.112 0.006 0.049 0.096 0.010 0.056 0.106
-$h=4$ asy-p 0.033 0.074 0.112 0.031 0.074 0.107 0.038 0.077 0.114
-chi-p 0.013 0.056 0.106 0.017 0.050 0.105 0.010 0.055 0.106
-$h=8$ asy-p 0.033 0.073 0.117 0.025 0.060 0.091 0.011 0.060 0.096
-chi-p 0.014 0.057 0.108 0.011 0.046 0.088 0.008 0.040 0.091
-7.1 $h=2$ asy-p 0.036 0.077 0.111 0.049 0.086 0.109 0.034 0.071 0.089
-chi-p 0.015 0.048 0.109 0.014 0.062 0.107 0.007 0.044 0.087
-$h=4$ asy-p 0.038 0.090 0.126 0.052 0.092 0.125 0.027 0.070 0.094
-chi-p 0.015 0.059 0.118 0.015 0.070 0.117 0.005 0.045 0.094
-$h=8$ asy-p 0.029 0.069 0.101 0.023 0.072 0.105 0.030 0.075 0.106
-chi-p 0.012 0.055 0.097 0.009 0.053 0.097 0.013 0.054 0.102
-7.1 $h=2$ asy-p 0.042 0.072 0.111 0.037 0.085 0.113 0.025 0.059 0.091
-chi-p 0.012 0.057 0.106 0.011 0.059 0.112 0.004 0.039 0.088
-$h=4$ asy-p 0.038 0.082 0.125 0.031 0.082 0.116 0.025 0.069 0.102
-chi-p 0.013 0.056 0.114 0.012 0.058 0.111 0.007 0.046 0.096
-$h=8$ asy-p 0.023 0.066 0.108 0.031 0.067 0.098 0.027 0.069 0.104
-chi-p 0.011 0.043 0.100 0.012 0.048 0.095 0.011 0.052 0.103
-$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.037 0.084 0.123 0.037 0.071 0.101 0.033 0.081 0.111
-chi-p 0.012 0.055 0.120 0.011 0.049 0.098 0.008 0.060 0.106
-$h=4$ asy-p 0.040 0.086 0.141 0.034 0.072 0.101 0.030 0.073 0.112
-chi-p 0.011 0.058 0.131 0.011 0.052 0.099 0.014 0.046 0.104
-$h=8$ asy-p 0.020 0.053 0.086 0.021 0.056 0.089 0.024 0.045 0.090
-chi-p 0.009 0.042 0.083 0.008 0.038 0.081 0.008 0.035 0.088
-7.1 $h=2$ asy-p 0.048 0.083 0.110 0.045 0.083 0.115 0.036 0.069 0.100
-chi-p 0.016 0.061 0.108 0.015 0.062 0.113 0.014 0.055 0.099
-$h=4$ asy-p 0.049 0.093 0.129 0.046 0.095 0.125 0.035 0.084 0.117
-chi-p 0.015 0.070 0.121 0.017 0.070 0.119 0.012 0.051 0.115
-$h=8$ asy-p 0.031 0.066 0.106 0.020 0.058 0.086 0.016 0.041 0.086
-chi-p 0.013 0.049 0.103 0.005 0.042 0.085 0.003 0.033 0.086
-7.1 $h=2$ asy-p 0.037 0.074 0.107 0.048 0.078 0.111 0.038 0.080 0.103
-chi-p 0.016 0.049 0.106 0.014 0.060 0.107 0.010 0.055 0.101
-$h=4$ asy-p 0.039 0.089 0.113 0.035 0.071 0.107 0.032 0.071 0.110
-chi-p 0.014 0.069 0.110 0.015 0.057 0.099 0.011 0.047 0.103
-$h=8$ asy-p 0.033 0.063 0.096 0.019 0.037 0.069 0.016 0.052 0.088
-chi-p 0.015 0.050 0.090 0.007 0.032 0.062 0.006 0.032 0.088
-Table 5: Empirical sizes for $\mathbb{W}_{2}$, rejection probabilities at 1%, 5% and 10% levels.
-$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.037 0.071 0.099 0.033 0.061 0.095 0.048 0.085 0.124
-chi-p 0.008 0.055 0.096 0.007 0.040 0.092 0.016 0.061 0.124
-$h=4$ asy-p 0.046 0.083 0.109 0.030 0.064 0.097 0.046 0.078 0.118
-chi-p 0.024 0.057 0.108 0.010 0.046 0.093 0.016 0.061 0.115
-$h=8$ asy-p 0.079 0.111 0.138 0.031 0.059 0.079 0.026 0.047 0.082
-chi-p 0.053 0.105 0.132 0.012 0.046 0.078 0.011 0.036 0.085
-7.1 $h=2$ asy-p 0.022 0.050 0.079 0.024 0.056 0.084 0.034 0.075 0.102
-chi-p 0.004 0.034 0.079 0.010 0.035 0.080 0.003 0.047 0.102
-$h=4$ asy-p 0.040 0.071 0.104 0.027 0.057 0.086 0.026 0.066 0.104
-chi-p 0.016 0.052 0.099 0.009 0.040 0.082 0.007 0.043 0.098
-$h=8$ asy-p 0.069 0.102 0.130 0.025 0.049 0.072 0.017 0.042 0.089
-chi-p 0.047 0.086 0.124 0.011 0.041 0.072 0.009 0.029 0.091
-7.1 $h=2$ asy-p 0.029 0.069 0.087 0.035 0.070 0.104 0.028 0.072 0.106
-chi-p 0.010 0.041 0.085 0.005 0.051 0.104 0.010 0.048 0.104
-$h=4$ asy-p 0.039 0.073 0.103 0.035 0.066 0.102 0.034 0.067 0.107
-chi-p 0.018 0.062 0.098 0.009 0.048 0.095 0.008 0.051 0.101
-$h=8$ asy-p 0.069 0.113 0.146 0.022 0.053 0.083 0.019 0.046 0.092
-chi-p 0.049 0.094 0.143 0.012 0.039 0.077 0.008 0.036 0.092
-$d_{\lambda}=4$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.036 0.064 0.085 0.036 0.062 0.088 0.040 0.074 0.097
-chi-p 0.015 0.046 0.083 0.011 0.045 0.088 0.018 0.052 0.097
-$h=4$ asy-p 0.046 0.089 0.116 0.037 0.060 0.092 0.040 0.074 0.104
-chi-p 0.028 0.063 0.113 0.019 0.050 0.086 0.018 0.052 0.097
-$h=8$ asy-p 0.097 0.139 0.170 0.042 0.061 0.083 0.024 0.054 0.086
-chi-p 0.065 0.122 0.166 0.028 0.053 0.077 0.006 0.043 0.088
-7.1 $h=2$ asy-p 0.028 0.053 0.082 0.026 0.055 0.075 0.032 0.071 0.099
-chi-p 0.013 0.032 0.081 0.006 0.031 0.074 0.010 0.048 0.096
-$h=4$ asy-p 0.044 0.079 0.109 0.027 0.054 0.073 0.032 0.063 0.095
-chi-p 0.023 0.058 0.106 0.011 0.043 0.072 0.010 0.043 0.087
-$h=8$ asy-p 0.070 0.110 0.140 0.028 0.054 0.073 0.026 0.053 0.088
-chi-p 0.050 0.094 0.137 0.018 0.044 0.070 0.006 0.042 0.090
-7.1 $h=2$ asy-p 0.029 0.052 0.070 0.027 0.061 0.086 0.027 0.054 0.077
-chi-p 0.010 0.039 0.067 0.009 0.039 0.085 0.009 0.037 0.074
-$h=4$ asy-p 0.055 0.093 0.128 0.027 0.058 0.089 0.031 0.060 0.089
-chi-p 0.029 0.071 0.122 0.012 0.036 0.086 0.014 0.043 0.082
-$h=8$ asy-p 0.090 0.128 0.157 0.029 0.053 0.075 0.022 0.044 0.085
-chi-p 0.061 0.112 0.150 0.015 0.040 0.068 0.012 0.035 0.086
-Table 6: Empirical sizes for $\mathbb{W}_{3}$, rejection probabilities at 1%, 5% and 10% levels.
-$d_{\tau}=2$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.037 0.071 0.099 0.033 0.061 0.095 0.048 0.085 0.124
-chi-p 0.008 0.055 0.096 0.007 0.040 0.092 0.016 0.061 0.124
-$h=4$ asy-p 0.046 0.083 0.109 0.030 0.064 0.097 0.046 0.078 0.118
-chi-p 0.024 0.057 0.108 0.010 0.046 0.093 0.016 0.061 0.115
-$h=8$ asy-p 0.079 0.111 0.138 0.031 0.059 0.079 0.026 0.047 0.082
-chi-p 0.053 0.105 0.132 0.012 0.046 0.078 0.011 0.036 0.085
-7.1 $h=2$ asy-p 0.022 0.050 0.079 0.024 0.056 0.084 0.034 0.075 0.102
-chi-p 0.004 0.034 0.079 0.010 0.035 0.080 0.003 0.047 0.102
-$h=4$ asy-p 0.040 0.071 0.104 0.027 0.057 0.086 0.026 0.066 0.104
-chi-p 0.016 0.052 0.099 0.009 0.040 0.082 0.007 0.043 0.098
-$h=8$ asy-p 0.069 0.102 0.130 0.025 0.049 0.072 0.017 0.042 0.089
-chi-p 0.047 0.086 0.124 0.011 0.041 0.072 0.009 0.029 0.091
-7.1 $h=2$ asy-p 0.029 0.069 0.087 0.035 0.070 0.104 0.028 0.072 0.106
-chi-p 0.010 0.041 0.085 0.005 0.051 0.104 0.010 0.048 0.104
-$h=4$ asy-p 0.039 0.073 0.103 0.035 0.066 0.102 0.034 0.067 0.107
-chi-p 0.018 0.062 0.098 0.009 0.048 0.095 0.008 0.051 0.101
-$h=8$ asy-p 0.069 0.113 0.146 0.022 0.053 0.083 0.019 0.046 0.092
-chi-p 0.049 0.094 0.143 0.012 0.039 0.077 0.008 0.036 0.092
-$d_{\tau}=4$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.036 0.064 0.085 0.036 0.062 0.088 0.040 0.074 0.097
-chi-p 0.015 0.046 0.083 0.011 0.045 0.088 0.018 0.052 0.097
-$h=4$ asy-p 0.046 0.089 0.116 0.037 0.060 0.092 0.040 0.074 0.104
-chi-p 0.028 0.063 0.113 0.019 0.050 0.086 0.018 0.052 0.097
-$h=8$ asy-p 0.097 0.139 0.170 0.042 0.061 0.083 0.024 0.054 0.086
-chi-p 0.065 0.122 0.166 0.028 0.053 0.077 0.006 0.043 0.088
-7.1 $h=2$ asy-p 0.028 0.053 0.082 0.026 0.055 0.075 0.032 0.071 0.099
-chi-p 0.013 0.032 0.081 0.006 0.031 0.074 0.010 0.048 0.096
-$h=4$ asy-p 0.044 0.079 0.109 0.027 0.054 0.073 0.032 0.063 0.095
-chi-p 0.023 0.058 0.106 0.011 0.043 0.072 0.010 0.043 0.087
-$h=8$ asy-p 0.070 0.110 0.140 0.028 0.054 0.073 0.026 0.053 0.088
-chi-p 0.050 0.094 0.137 0.018 0.044 0.070 0.006 0.042 0.090
-7.1 $h=2$ asy-p 0.029 0.052 0.070 0.027 0.061 0.086 0.027 0.054 0.077
-chi-p 0.010 0.039 0.067 0.009 0.039 0.085 0.009 0.037 0.074
-$h=4$ asy-p 0.055 0.093 0.128 0.027 0.058 0.089 0.031 0.060 0.089
-chi-p 0.029 0.071 0.122 0.012 0.036 0.086 0.014 0.043 0.082
-$h=8$ asy-p 0.090 0.128 0.157 0.029 0.053 0.075 0.022 0.044 0.085
-chi-p 0.061 0.112 0.150 0.015 0.040 0.068 0.012 0.035 0.086
-Table 7: Empirical sizes for $\mathbb{W}_{4}$, rejection prob. at 1%, 5% and 10% levels. Testing $\delta(z)$.
-$d_{\lambda}=1$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.065 0.080 0.129 0.053 0.086 0.110 0.050 0.072 0.090
-chi-p 0.022 0.053 0.124 0.027 0.069 0.109 0.021 0.059 0.089
-$h=4$ asy-p 0.042 0.085 0.112 0.071 0.106 0.130 0.047 0.072 0.098
-chi-p 0.039 0.071 0.103 0.026 0.090 0.126 0.029 0.061 0.092
-$h=8$ asy-p 0.169 0.124 0.159 0.019 0.060 0.074 0.022 0.039 0.071
-chi-p 0.144 0.101 0.154 0.025 0.054 0.071 0.016 0.036 0.080
-7.1 $h=2$ asy-p 0.076 0.084 0.144 0.056 0.078 0.090 0.041 0.060 0.070
-chi-p 0.031 0.055 0.137 0.028 0.071 0.087 0.017 0.053 0.069
-$h=4$ asy-p 0.134 0.085 0.133 0.075 0.097 0.115 0.054 0.077 0.093
-chi-p 0.099 0.064 0.125 0.015 0.065 0.111 0.014 0.060 0.091
-$h=8$ asy-p 0.174 0.124 0.153 0.053 0.081 0.094 0.025 0.045 0.080
-chi-p 0.142 0.103 0.150 0.022 0.072 0.092 0.015 0.036 0.083
-7.1 $h=2$ asy-p 0.032 0.059 0.126 0.058 0.080 0.098 0.051 0.077 0.095
-chi-p 0.018 0.039 0.122 0.033 0.067 0.094 0.017 0.062 0.093
-$h=4$ asy-p 0.072 0.094 0.141 0.060 0.100 0.129 0.055 0.084 0.104
-chi-p 0.095 0.070 0.137 0.041 0.077 0.122 0.016 0.070 0.101
-$h=8$ asy-p 0.063 0.089 0.151 0.047 0.067 0.083 0.025 0.049 0.076
-chi-p 0.109 0.070 0.147 0.020 0.063 0.078 0.016 0.036 0.088
-$d_{\lambda}=2$ $n=200$ $n=500$ $n=900$
-1% 5% 10% 1% 5% 10% 1% 5% 10%
-7.1 $h=2$ asy-p 0.086 0.117 0.135 0.061 0.086 0.099 0.028 0.053 0.076
-chi-p 0.061 0.101 0.132 0.036 0.075 0.098 0.016 0.041 0.082
-$h=4$ asy-p 0.122 0.163 0.187 0.063 0.084 0.111 0.041 0.059 0.068
-chi-p 0.091 0.144 0.184 0.044 0.072 0.107 0.030 0.048 0.082
-$h=8$ asy-p 0.145 0.195 0.227 0.043 0.065 0.089 0.028 0.047 0.058
-chi-p 0.109 0.177 0.225 0.030 0.057 0.083 0.018 0.038 0.081
-7.1 $h=2$ asy-p 0.083 0.113 0.141 0.057 0.081 0.098 0.034 0.051 0.067
-chi-p 0.052 0.101 0.136 0.036 0.067 0.096 0.018 0.042 0.085
-$h=4$ asy-p 0.117 0.154 0.183 0.060 0.090 0.113 0.028 0.046 0.063
-chi-p 0.087 0.138 0.176 0.045 0.078 0.109 0.019 0.040 0.081
-$h=8$ asy-p 0.134 0.188 0.218 0.041 0.067 0.086 0.027 0.046 0.068
-chi-p 0.105 0.167 0.214 0.032 0.060 0.085 0.016 0.036 0.081
-7.1 $h=2$ asy-p 0.093 0.131 0.148 0.045 0.066 0.091 0.036 0.058 0.069
-chi-p 0.057 0.114 0.145 0.030 0.056 0.087 0.020 0.048 0.087
-$h=4$ asy-p 0.135 0.177 0.203 0.064 0.088 0.107 0.030 0.047 0.076
-chi-p 0.090 0.163 0.200 0.043 0.074 0.104 0.018 0.038 0.084
-$h=8$ asy-p 0.148 0.202 0.228 0.060 0.080 0.092 0.022 0.036 0.072
-chi-p 0.115 0.179 0.227 0.039 0.077 0.089 0.016 0.030 0.082
+Figure S1: Cobb-Douglas versus semiparametric production.